A Complete Bibliography of Publications in Annals of Statistics (2010–2019)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

19 May 2021 Version 1.20

Title word cross-reference

1 ≤ r ≤ ∞ [GN11]. _{L} [ZFY11]. A [SFSL18]. β [MMS18, RPF13]. D [SH15]. ℓ_{0} [FG18a, vdGB13]. ℓ_{1} [JKNP12, BC11a, RWL10]. F [HPZ16]. g [MG11]. k [BSY19, BCR10, GKM16]. L^{r} [GN11]. L_{1} [HPŠ10a, WXX18]. L_{2} [EI12]. L_{∞} [DKL12]. l_{p} [LZ11]. L_{q} [FY10]. M [CH10, DN11, EKS12, Loh17, SO18, WZ18, ZBFL18]. n [CW12, Por12]. P [EJ10, CW12, HBZO19, RBWJ19]. Q [DRZ12, LR14]. s [CLSY18, HW16a]. T [DMS12, DMS13a, DMG15]. T^{2} [LS13]. U [Che18, LLBM^{+}11]. V [Zho14b]. ∥ · ∥_{2}^{2} [JN18]. z [BCCW18].

-aggregation [DRZ12, LR14]. -concave [HW16a]. -distance [WXX18]. -error [DKL12]. -estimation [BCCW18, CH10, SO18, WZ18, ZBFL18]. -estimator [EKS12]. -estimators [Loh17]. -filter [RBWJ19]. -learning [SFSL18]. -likelihood [FY10]. -loss [JN18]. -metrics [GN11]. -model [MMS18, RPF13]. -nearest [BSY19, BCR10, GKM16]. -optimal [DMS12, DMS13a, DMG15, SH15]. -penalized [BC11a, FG18a, vdGB13]. -prior [MG11]. -processes [LLBM^{+}11]. -regularized [RWL10]. -sample [CLSY18]. -statistic [LS13]. -statistics [Che18, Zho14b]. -theorem [DN11]. -value [HBZO19]. -values [EJ10].

1/64 [Pho12]. 1D [SGA11].

2010 [Ste11]. 2850-2876 [BGvZ21].

34 [Fuh10, Jen10].

abundant [CFR12]. accelerated [LJ18]. accounting [PP14]. Accuracy [CG18, JKNP12]. Accurate [HQ11]. achieve [Dic11]. acting [DKZ13]. Active [HSRW19, NJ13, Han11]. activity [ASJ11, Bul16, Tod15, Tod17]. acyclic [CMKR12, ER14, HJY13, KMH14, vdGB13]. Adaptation [KGS18, PR16, GZ16a, Lep13]. Adaptive [Arm15, BW19b, CY12, CLX13, CGW18, CV19a, CL13b, CJ12, FFB14, FLRB13, GMZ17, JMR14, LS14, Lep15, LMP18, Ray17, RBS10, Str18, TZ19a, Tib14, Ver10, dJvZ10, AGR18, AZ12, Atc11, BCDD14, Bul13, CMW13, CLZ16, Cas15, CC18a, CCK14a, CCTV18, DGL19, EL10, FMP11, Gho15, GRSP15, GL11, HWH13a, HWH13b, HMV13, HN11, HRSH15, HH12, KX14, LW19, LM15, MP17, Nic15, PR19, Rou10, Rou15, Ste15, SC16, SvdVvZ15a, SvdVvZ15b, TBK12, TLT16, WCT11, ZPH15, ZH10]. Adaptive-to-model [TZ19a]. adaptivity [CG17]. Addendum [KMH14]. Additive [ST19, BPE14, FJR15, HP18, HHW10, JFF10, LMP10, LMP12a, LLLH18, Ma12, MCLX15, RBM19, SW16a, TZ19b, WLLC11, WXQL14, YZ16, ZL11a]. Adjusted [LC10, AGR18, DHZ16, JW10, ZBFL18]. adjustment [SB10, WZHO17]. affine [Bel18a, DS12]. after [TRTW18]. against [CEL16, CPV17, MMY18]. aggregate [DGP18]. Aggregation [GRSP15, Bel18a, DRZ12, DS12, LR14, Rig12]. agnostic [KP19, LV15]. AIC [BCF18a]. Albert [QH19]. Algorithm [ZLZ18, BWY17, CMZ19, GRD13, JWJW18, JG12, JG13, KRS15, QH19, Yu10]. algorithmic [FLSZ18, Lop19]. algorithms [BCDD14, DR11, DMJ13, FMP11, KH11, Lia10, STRR15]. aliasing [CCC19]. aligned [CW15]. allocation [CGM^{+}13, Kau18, MP17]. almost [LPK17]. almost- [LPK17]. alternative [ZCX13]. alternatives [ACCP11, CEL16, CPV17, EH19b, MMY18]. Amplitude [PZ16]. analyses [Røy12]. analysis [Bac11, BFT13, BL12, BCF18a, BWY17, BCR10, BJV17, BFR19, BMBM15, BD11, CEL16, CGY18, CLLW15, DM18, DP19, DP18, FLW16, GMRZ15, JW10, JLP^{+}16, JLX17, Kat13, KLZ16, KS11, KMS12, LH10b, LJ10, LMY18, LY19, Ma13, Mül16, NC10, Nye11, PT13, Pen18, PM16a, QP16, QH19, RBS10, RH12, SY12, SWDW11, SC12, TS12, Wan11, WMZ18, WW18, Zhu17]. analytic [KH11]. angular [ZYL18]. Anisotropic [BPD14, Lep15, Str18].

Ann [BGvZ21, Fuh10, Jen10]. anomalous [ACCD11]. ANOVA [ACCP11, CZ10, CLZ19a]. Answer [Jia13]. Anti [CCK14a]. Anti-concentration [CCK14a]. application [BLPZ15, BBFGD12, DN11, EH19a, HS17, HLLZ19, HKKM10, Koi19, LSST16, LMY18, LS10, MDO14, NC10, PR16, Pho12, RBS10, RV11, WY17, ZPMX11]. applications [BB18a, BQS18, CJ11, CZ18, CTP19, CV19b, CI11, CL13a, CQ10, Che18, CD17, DE17, DSS11, FSZ18, God18, GZ16b, HJY16, KR11b, LS13, TPN19, ZHCC11, ZBY15, Zho14b, ZBFL18]. applied [DH12a]. approach [AcL17, BBTW11, BG10, CCG16, CC18b, CHZ11, DKZ17, DW19a, DP18, FLRB13, HW16b, HL13, JP13, KPS17, KX14, Lep13, Lep15, MZZ^{+}19, Pho12, Sch13, VS19, Yen11, YC10, ZPMX11, ZYFW14]. approaches [BD11]. Approximate [BO17, DGH^{+}19, FG18a, CC11, FLM11, MMB17, MMB18]. approximately [KP19]. Approximating [WRM19]. Approximation [DSS11, HW16a, Nor10, BCCZ13, CCK14b, DB16, FH14a, HPWW11, HBZO19, Koi19, KL17, LLV16, Lia10, Por12, ZW17]. approximations [Bob19, Che18, CCK13, FR13, KR11b]. arbitrary [CL15a, EKS12]. argument [Jia13]. arithmetic [Pho12]. ARMA [LXQ16, VL18, ZL11c]. ARMAX [GH19]. armed [PR13]. array [HQ14, LKA10]. array-valued [LKA10]. arrays [GX14, HT14, HCT18, LPK17, OR10, OB13, STY14]. arrival [CLN19]. article [Ano10, Fuh10, Jen10]. Assessing [ZYL18]. assessment [CG18]. assisted [KJF14]. associated [LX12, RS17]. association [GCW19, JLP^{+}16]. assumption [RWG19, URBY13]. assumption-lean [RWG19]. assumptions [HSRW19]. asymmetric [FZ13, GZ16b]. Asymptotic [Akr16, BM17, BCCZ13, BCFG11, CHI14, CGW15, CC19, DM12, EPFV18, GNZ10, HPWW11, HA13, HH12, KF11, KV10, LLBM^{+}11, MH14, MMB11, Mei11, OMH13, OPP17, PJS10, QZ10, Rei11, RSZZ15, RS17, SH14, TA17, Wan13, WW18, XH12, Atc11, BF18, BGN18, LX12, Loh17, MDO14, NS11, PS18, RS12, SC13, SF19, TRTW18, YFG13]. Asymptotically [CEKL15, Fel12, CR13, SC16]. Asymptotics [LMNP15, SW10a, SBK16, WF17, YLZ16, BFFP19, CS15, CW19, DW18, DMJ13, HE18, Li14, ZS13]. attack [CHL17]. augmentation [KH11]. auto [LWY17]. auto-covariance [LWY17]. autocorrelation [HA13]. autocovariance [BB16b]. autoregression [ZPL^{+}17]. autoregressive [GRSP15, KPP11, WLCY14, Zhu19]. auxiliary [Efr13]. average [YL11]. averages [DS11, JPV10]. averaging [AcL17, Lia10, ZL11a].

back [MC15]. back-door [MC15]. Backfitting [LMP10, LMP12a, HP18]. Backward [HE18]. balance [MR12]. Balanced [RW10b]. balancing [Zha19b]. Ball [PTWZ18, LM18]. balls [CCGB17]. band [WLCY14]. bandedness [QC12]. bandit [PR13, PRCS16]. bands [CCK14a, GN10, HH13, HN11, Jir12, PR19]. Bandwidth [CL15b, GL11, BPD14, CDH13, QC12, SW10a]. barrier [Men17]. barriers [GMZ17, MW15]. Barycentric [Pen18]. based [AGR18, BB16a, BWY17, CY11, CGT16, CC17, CZ15b, Che19a, CAGPFBGM19, FH14a, HP10, HPV10, HQ14, IP11, LLL19, LPZZ12, MZZ^{+}19, NHM18, NC10, OPP17, Røy12, SvdAW14, SSGF13, Ste15, TBK12, TA17, WB17, WXX18, WZ18, XZ18, XZ12, YFG13, YP15, YL11, ZDMD18]. Batch [FJ10]. Batched [PRCS16]. Bayes [BCFG11, BMW18, BD11, CN14, DP18, LM12d, MG11, MH11, MGC10, RS17, SB10, YL14]. Bayes-optimality [BCFG11]. Bayesian [Cas15, Gho15, LM15, Nic15, Rou15, SvdVvZ15b, Bac11, BFT13, BBFGD12, BPY19, BFR19, BD11, BD15, CKG19, Cas14, CSHvdV15, CC18b, DMG15, DP18, Eva18, GZ16a, HTX17, Jam17, Joh13, Kat13, Kau18, KPS17, KvdVvZ11, LG17, LJ10, MNP19, NH14, NS17, PBPD14, QH19, Roč18, RH12, RCL12, SvdVvZ15a, WM10, YH12, YD16, YWJ16, dJvZ10]. be [EW16]. behavior [BCF18b, PP14]. behaviour [BB16b, RS17]. benefit [HZ10]. Berkson [Sch13]. Bernstein [BK12, BG14, Bon11, CN13, CN14, CR15, HS17, Ray17, RR12]. Bernstein-type [HS17]. Best [BKM16, MM18]. Betas [Rou10]. between [DMS12, LZ17, SS13]. beyond [DS11, Jia13, ZW16]. bi [YLZ16]. bi-degree [YLZ16]. Bias [FdHM15]. biased [AZ12, CLSY18, CK15, MR10]. BIC [BCF18a]. big [BFR19]. binary [Fry14, MPL15, Wan11, Xu17]. bivariate [WZ12]. Blackwellization [DR11]. blind [BHM18]. block [AL18, CY12, CLX18, FdH15, GZ16a, GMZZ18, GLN18, JKNP12, LR15, Lei16, SR15, TPN19, WB17, ZZ16, ZLZ12]. block-sparse [JKNP12]. blockmodel [BRS19, RCY11]. blockmodels [BCCZ13, SB15]. blocks [God18]. BLUE [DPZ19]. Blumenthal [ASJ12]. Board [Ano16a, Ano16b, Ano17a, Ano16c]. boosting [FGM17, EI12]. Bootstrap [ABFRB15, CH10, SZ15, SW19, BB15, CL13b, Che16, Che18, CCK13, DGL19, FR13, FLRB13, GLN18, HH13, JP15, KP11, KPP11, Lop19, Pap18, PS18, RS12, SS11a, SBW10, TPN19, TRTW18]. Bootstrapping [RWG19]. both [XMW10]. bottom [Fry18]. bottom-up [Fry18]. bouncy [DBCD19]. bound [NS11]. boundaries [CLR17, KRS15, LG17]. boundary [TBK12]. bounds [ACL19, AGL14, Bel18a, BLT18, BJNP13, CZ18, CGM^{+}13, CI11, CHI13, CGS15, JvdG18, KP19, Lev15, LN11, PS18, TS12]. breakdown [ZYL18]. breaking [Men17]. Bridging [ELL15]. Brownian [ASJ10]. Buffet [Jam17]. bundled [DN11].

calculations [RV11]. calculus [Hua10b, Jam17]. Calibrating [WKL13]. calibration [TW15]. CAM [BPE14, PV17]. can [EW16]. Canonical [BHPZ19, GMRZ15, YP15]. Carlo [Atc11, BDD10, CDO11, FJ10, FMP11, GRD13, MDO14]. case [BHPZ19, FH14b, KF11, KMS12, LW17, OMH14]. case-control [FH14b, KF11]. categories [WZY18]. Cauchy [PM16b]. Causal [BPE14, REB18, RBM19, ST16a, ZJS11, CD17, JBGWS13, Lok17, MZZ^{+}19, TS12, URBY13, ZDMD18]. cause [VR12]. CCA [GMZ17]. cell [HE18]. censored [DN11, GK12, Hua10b, SW12, SW16b, ZPH18]. censoring [ACF12, Gro14]. Central [JY13, BLRG17, EH19a, HQ14]. centrality [ELL15]. centroid [HP10]. centroid-based [HP10]. cepstral [MH14]. chain [Atc11, BB16a, BDD10, CDO11, FJ10, FMP11, MDO14, Sad16, WR13]. chains [Bac11, BFT13, HW16b, LTG16, MDO14, Tru19]. Change [SS11a, WZY18, BJV17, CZ15b, Che19a, CC19, EH19b, Fry14, Fry18, HTX17, Jir15, KRD10, LSR^{+}17, PT10, SBK16, WZ18, XS13, ZYFW14]. Change-point [SS11a, WZY18, BJV17, CZ15b, Che19a, CC19, EH19b, Fry14, Fry18, HTX17, SBK16, XS13, ZYFW14]. changed [Bel11]. changes [Bul14, DW19a, VD15]. characteristic [Bel11]. characterization [HT14]. Characterizing [EI12]. Chebyshev [WY19]. checking [EJ10, TZ19a]. Cheng [Fuh10, Jen10]. Chernoff [Li16a, LLY18]. Chib [QH19]. CHIME [CMZ19]. choice [BBFGD12, CDH13]. Cholesky [LLL19]. circulant [CK15, LPK17]. circular [ZAL17]. class [KLN12, KH11, LPS08, LP10, Xu17]. classes [CCT17, DS13, HJY13, LM12c, Lep15, Str18]. classical [JY13]. Classification [BCDD14, GKM16, DW18, DKR18, FJY13, FKLZ15, HJY16, NR12, Sam16, SEG19, TSP13, ZYL18]. classified [BMW18]. classifier [CDH13]. classifiers [HP10, Sam12]. clinical [AZ12, ZH10]. closed [Li13]. closed-form [Li13]. CLT [ZPG18, ZBY15]. cluster [ACCD11, Che19c, DR10, Ma12]. clustered [CHL16, SW16b, Wan11]. Clustering [CMZ19, ACV16, CZ16, CW14, Cho17, JW16a, JW16b, JKW17, JY16, LR15, Nad16, NC10, RW10a, RCY11, SB15, SWY15, SC12, SEC14, Ste15, ST16b]. clusters [Wal10]. Co [CW14, Cho17]. Co-clustering [CW14, Cho17]. coalescent [CMW14]. CoCoLasso [DZ17]. code [MT13, Pho12, ZPMX11]. coefficient [BW19b, CHL16, FR13, KLZ11, KP15, LMP12b, LXQ16, LKZ15, MCLX15, ZLK12]. coefficients [BHPZ19, CGZ18, XMW10, YP15, Zhu17]. coherence [CJ11]. coin [AZ12, MR10]. collaborative [BCR10]. column [ZCX13]. column-wise [ZCX13]. Combinatorial [NLL19, ABBDL10]. comments [LK10]. common [BG10, FSD^{+}17]. communication [Fel12]. Community [ACV14, GMZZ18, ACBL13, BCZ^{+}15, CL15a, Jin15, LLV16, ZZ16, ZLZ12]. comparative [AZ12]. comparing [DS16]. Comparison [MDO14, KH11]. comparisons [ACCP11, HSRW19]. competing [BD13]. complementary [MT13]. Complete [DS13, KMS12, SR15]. completion [DP19, KLT11, Wan13, Zha19a]. complexity [FLSZ18, Men17, QH19, YWJ16]. component [BCF18a, CGY18, DM18, FLW16, IP11, LZW10, LH10b, Ma13, SY12]. components [BCF18a, BBTW11, BR13, CTT17, DE17, FJ19, HPV10, JW10, Nye11, WBS16]. composite [Bel11, CL11, KX14, XZC12]. compressed [CJ11]. compromise [MP17]. Computation [DFW19, XH12]. Computational [BD11, CLR17, MW15, GMZ17, WLZ14, WBS16, YWJ16]. computationally [CL15a]. compute [GRD13]. computer [HQ11, LBST10, TW15]. Computing [SH15, Yu10]. concave [DW19b, DSS11, FXZ14, FZ19, HW16a, KS16, KGS18, KM10b, LYL16, Zha10]. concentration [CvdV12, CCK14a, HRSH15, KL17, PV18]. Conditional [Jac10, ZYS18, CD17, Han19, Hua10a, Lee13, LJ11, Nor10, PR12, SSL19, XZC12]. conditioning [Sad16]. conditions [CD12, Ler11]. cone [CCF14, SH15]. cones [WWG19]. Confidence [ABR10a, CG17, FLR^{+}14, GN10, NvdG13, SHMD13, BLP19, BK18, BCCW18, CLX13, CGM^{+}13, CCGB17, CCK14a, GJ15, HH13, HN11, Jir12, LK10, NL17, PR19, PS18, SZ15, Wan10, WLCY14, YL14]. configurations [RV11]. Confounder [WZHO17, VS13]. conical [PJS10]. conical-hull [PJS10]. connection [EW16]. conquer [BDS19]. consequences [YS12]. considering [XMW10]. Consistency [BCF18a, CDO11, DL17, DHSM18, DMOvH11, GD17, LR15, MMNP15, MGC10, SBV15, SR13, ZLZ12, BCZ^{+}15, CKG19, CH10, CD12, Jia13, LLL19, LJ11, LZL18, LSR^{+}17, Loh17, NHM18]. Consistent [KJ16, MMB17, MMB18, And10, TSP13]. constant [FG18a, XMW10]. constants [ULR18]. constraint [Cha14, CLZ19b, LS14]. constraints [CLX13, El 10a, Fel12, SHMD13]. constructing [CDH13, HH13, LBST10]. construction [CJ11, LK10, SLL11, SLQ14, Wan10]. contamination [CGR18]. Contents [Ano16d, Ano16e, Ano16f, Ano17b]. context [BO17, CD12]. continuous [ASFJ10, CDO11, DPZ19, Roč18, WCT11, ZJS11]. continuous-time [ASFJ10, DPZ19, ZJS11]. contraction [Atc17, Cas14, GZ15, GN11, NS17, PBPD14, YG16]. contrastive [JWJW18]. contributions [Che16, CF16, Del16]. control [FLSZ18, FH14b, GS10, JM18b, KF11, Liu13, RW10b, SF19]. Controlling [BC15, DR15, GHS14, PHW11]. Convergence [FMP11, HW19, HN16, Jan14, JWJW18, LZW10, Mas13, Ngu13, QH19, RH12, WR13, XQB13, ANW12a, BB18b, BSV14, CJ10, CZZ10, CZ12, CLZ16, CLM16, CL13b, DHSM18, GCW19, Gun12, Han11, KS16, LLL19, LH10b, PS10, Rou10, WLZ14, Yu10]. Convex [RYC19, ANW12b, CGW18, CS12, CPW12a, CPW12b, Cha14, GT12, Gun12, Han19, LM12b, RZ12, SS11b, SW10b, Wai12, WWG19, XCL16, Yua12]. convex-transformed [SW10b]. Convexified [CLX18]. convexity [EKDN18]. convolution [LW19]. Convolved [TPN19]. Cook [ZIC12]. Coordinate [CZC10, ZLZ18]. Coordinate-independent [CZC10]. copula [LHY^{+}12, SvdAW14]. copulas [BV13, CEKL15]. correct [KP19]. corrected [CLX18, GMZZ18, ZLZ12]. Correction [BGvZ21, BCZ^{+}15, CL10, DR16, DG12, FXZ15, FZ21, HWH13a, JW15, JG13, LP10, LTTT14a, OB13, FdHM15]. correlated [CL18, DPZ13, DPZ16, DKZ17, DPZ19, HJ10, JFF10, WXQL14]. correlation [BLPZ15, BHPZ19, ESW17, FR13, GMRZ15, Hua10a, HJY16, KLFL17, LPZZ12, PY12, SR14, YP15, ZCGZ19]. correlations [ACBL12, Dob17, FSZ18, LD18]. count [BB15]. countably [SSK14]. Counterfactual [Røy12, Lok17]. counting [MH13]. Coupling [Cha15b].

Covariance [CXW13, JP15, XW12, APT17, BPZ15, BFFP19, CZZ10, CJ11, CY12, CZ12, CGR18, CW12, DGJ18, FLM11, FRW15, FLW18, HLLZ19, JPSZ10, KR11a, KL17, Lam16, LW12a, LC12, LWY17, LW13, PBPD14, QC12, SN19, WF17, XQB13, XZ18, ZPW13, ZL11b, ZBY15, ZCCL19, ZLST17]. Covariate [JW10, KJF14, Zha19b, AGR18, DHZ16, HH12, MP17, MR12]. covariate-adaptive [HH12, MP17]. covariate-adjusted [AGR18, DHZ16]. covariates [AZ12, JFF10, MCLX15, MRS12, PR13, Sch13, SW12, Wan11, WXQL14, ZJS11]. covariation [BHMR14]. coverage [Bel17, Cas15, Gao13, Gho15, LM15, Nic15, Rou15, SvdVvZ15a, SvdVvZ15b]. Cox [BFJ11, DML10, HSY^{+}13, LLY18, QWW16]. Cramér [CSWX16, LS10, LS13]. Cramér-type [LS10, CSWX16]. credible [Bel17, Cas15, Gho15, LM15, Nic15, Rou15, SvdVvZ15a, SvdVvZ15b, YG16]. Criteria [BBFGD12, AL11, HYS15, KJ16]. criterion [ATHW12, CHI14, FDD12, MC15, SEG19, SO18, ZL11a]. critical [Roh19]. criticism [ACCP11, HJ10, ZCX13]. Cross [RD19, BD14, BMW18, CGZ18, XH12, Zha19a]. cross-classified [BMW18]. cross-over [BD14]. cross-validation [CGZ18, XH12]. crossover [NK18, Zhe13a, Zhe13b]. Current [GH18, GJW10, TBK12, WZ12]. curse [BK19]. Curvature [Efr18, AL19]. curve [BYZ10]. curves [BG10, BG13, DS16, FSD^{+}17, KF11, Zho10]. cyclic [CLN19, DFW19].

DAG [CKG19, LLL19, MW16]. Dantzig [RBM19]. data [ASJ10, BL14, BLN15, BS14, BFR19, BJNP13, CY11, CDH13, CC17, Cha17, CZ10, CQ10, CM12, CLLW15, CHLP14, CHL16, CW14, CC19, CGC11, DM18, DH12a, DH12b, DP19, DN11, ELL15, EPFV18, ER14, FFN10, FH14a, FH14b, Fry18, GK12, HP10, Han19, HWH13a, HWH13b, HK10, JW10, JW11, JYW14, JW15, JKL12, KLZ16, KH11, Koi19, KLJ15, Kon18, KV10, Lee13, Ler11, LH10a, LH10b, LS17, LZL19, LMY18, LY19, LW12b, Loh15, MB15, MY10, Mül16, NMR17, OMH13, PM16a, Sae19, Sam16, SWDW11, SL19, SW12, SW16b, TZ19b, TBK12, WZ10, Wan11, WXQL14, WZY18, WZ12, YP15, ZFY11, ZW16, ZCL16, ZPH15]. Debiasing [JM18a]. Deciding [LH10a]. decision [PHW11]. decomposable [KR11a]. Decomposition [BB10, ANW12b]. decompositions [Fry18, JBD17]. deconvolution [BG10, DGJ11, Del16, GPPVW12, HKKM10, KNP11, LN11, PS10, Pen17, SHMD13]. deep [BK19]. defined [Rao18]. Defining [DH10]. definition [VS13]. definitive [SEG19]. degree [BCL11, CLX18, GMZZ18, YLZ16, ZLZ12]. degree-corrected [CLX18, GMZZ18, ZLZ12]. Degrees [TT12a, SBK16]. delivers [JJ12]. Delta [GZ11]. demographic [BS14]. denoising [DG14b]. dense [ACV14, CHL17, ZW16]. densities [BW19a, HW16a, LPS08, LP10, Nor10, QT19, SW10b]. Density [ASW13, AS11, ACF12, BGK10, CJ10, CLSY18, CL13c, CGW15, DH10, DW19b, GN10, GL11, GNZ10, HKKM10, JPSZ10, KS16, KGS18, KM10b, LMNP15, LMNP17, Lep13, LW19, Ler11, Li13, MJ15, PR16, Pen17, PM16a, RW10a, RR12, Rou10, RCL12, SW10a, Ste15, Str18]. density-based [Ste15]. Density-sensitive [ASW13]. dependence [BKRS19, BDV11, BSV14, CSWX16, Chi11, EKS12, EBGG18, Lah10, LLBM^{+}11, LZ17, MMB11, ZYS18, ZCX13, ZBFL18, ZZX18]. dependence-adjusted [ZBFL18]. dependent [EPFV18, HK10, SO18, SN19]. depending [Efr13]. depth [CGHH17, ELL15, HPŠ10a]. depths [PV18]. derivative [ABT11]. Descartes [BS14]. descriptors [HE18]. design [AGR18, AZ12, CZvdL17, DPZ13, LPK17, PP14, Roh19, SS11a, SD12]. Designs [God18, AZ12, BD14, BLRG17, BD13, CC18b, CK15, DGH^{+}19, DM11, DMS12, DS13, DMS13a, DG14a, DMG15, DS16, DPZ16, DKZ17, FSD^{+}17, HQ14, HYS15, HW17, JM18a, LZA15, LBST10, LMY19, MT13, NK18, Pho12, SH15, SR15, SEG19, SLL11, SLQ14, SWX19, TXL12, WXX18, YS12, Yu10, ZPMX11, ZDMD18, Zhe13a, Zhe13b, Zhe15, ZAL17]. Detecting [DW19a, QCN18, VD15, HJ10, PT10, STY14]. Detection [ACCD11, ACBL12, MMS18, VAC17, ACBL13, ACV14, BR13, BCZ^{+}15, Bul17, CL15a, CW15, Cha17, CZ15b, Che19a, CS17, CC19, Dob17, EH19b, Fry14, Fry18, GMZZ18, HTX17, ISS12, Jin15, LLV16, LG17, MW15, OMH14, SGA11, Wal10, WZY18, WZ18, XS13, ZZ16, ZLZ12]. deterministic [SD12]. Deviation [DRZ12, LS10, LS13]. deviations [CSWX16, GZ11, Gao13, GXZ18]. diagram [DFKP19, JJ12]. diagrams [FLR^{+}14]. dichotomous [VR12]. dictionaries [Pen16, WCT11]. difference [Liu17, LK10, Wan10]. differences [LZ17]. differentiable [HYS15]. differential [PL11, QZ10, XMW10]. differentiation [HE18]. diffusing [NH14]. diffusion [BGK10, CC11, Li13, ZL11b]. diffusions [MS10b, NS17, Str18]. digital [Dic11]. dimension [ABR10a, ABR10b, BL12, BR13, CZC10, FKLX15, JP15, Jir15, KX14, LLC13, LH10a, LKA10, LAL11, LS17, LLY14, MZ13, MZZ^{+}19, MGC10, OMH14, SW16a, WKL13, YL11, ZPW13, ZYS18]. dimensional [ANW12a, ATHW12, AcL17, ACV16, Atc17, BCF18a, BCF18b, BW19a, BLPZ15, BHPZ19, BC19, BM15, BFL^{+}18, BC11a, BB16b, BW19b, BJNP13, BDP19, BPE14, BSW11, BSW12, CZ16, CLR16, CG17, CG18, CZ18, CMZ19, CKG19, CTP19, CC17, CTW18, CQ10, CXW13, Che18, CL18, CLZ19a, CHLP14, CCK13, CMKR12, CFR12, CF19, CGZ18, CPV17, DZ17, DW18, EH19a, EH19b, FLM11, FKLZ15, FJ19, GKM16, Gol10, GZ16b, HP10, Han19, HWH13a, HWH13b, HLLZ19, HMLZ11, JvdG18, JJ12, JY13, JLP^{+}16, JW16a, JW16b, JKW17, KP15, KS11, KLFL17, KV10, LY12, LW12a, LZW10, LLL19, Lee13, LH10a, LC12, LKZ15, LZ17, LLYY19, LZL19, LHY^{+}12, LW12b, Loh17, MCLX15, MPL15, Nad16, NMR17, NHM18, NW11, NL19, NL17, NZL17, OWJ11, OMH13, QC12, RYC19, RWL10, RWG19, RLM^{+}17, RT11b, RCY11, Sam16, SWDW11, SD12, SFSL18]. dimensional [SSCL19, SL19, ST16b, TZ19b, TWZ13, VV10, WF17, WY17, WH19, XQB13, XZ18, XCL16, XZ12, YP15, YWJ16, YDS16, YZ16, Zha13, ZW17, ZPG18, ZLLW13, ZL11b, ZPH15, ZBY15, ZPH18, ZCCL19, ZCGZ19, ZCX13, ZLST17, Zhu17]. dimensionality [BYZ10, BK19, BFJ11, CD12, El 10a, FS10, Lv13]. dimensions [ANW12b, DGK19, EKS12, FL14, HWCS19, LM12a, LD18, LZL18, VL13, YT15]. direct [LSR^{+}17]. directed [CMKR12, DP13, ER14, HJY13, KMH14, YLZ16, vdGB13]. direction [Akr16]. directional [PV17, PJS10]. directions [MW16]. Dirichlet [KGC10]. discontinuous [SO18]. discoveries [FSZ18, SBC17]. discovery [BC15, DR15, GHS14, JM18b, Liu13, NR12, PHW11, RV11, WR13]. discrete [CDH13, DLP12, ER14, Eva18, Khm13, LM12d, LW13, MW16, WRM19, ZJS11]. discretely [ASJ12, BG13, CY11, NS17]. discretized [JPV10]. discriminant [SWDW11]. Discriminating [Li16a, BD13, DMS13a, DMG15]. discrimination [DMS12, HW17, Huc11]. Discussion [ACV16, BMvdG14, BB14, CY14, CZ16, CS12, Cas15, FK14, Gho15, GT12, KM10c, LM12b, LM15, LZ14, Nad16, Nic15, RZ12, Rou15, SZ10, ST16b, Wai12, Was14, Wei10, Yua12]. discussions [SvdVvZ15b]. dissimilarities [SR14]. Distance [BKRS19, KLFL17, Li16a, SSGF13, SR14, VL18, WXX18, ZIC12]. distance-based [SSGF13]. distances [BSY19]. Distributed [BFL^{+}18, FWWZ19, VCC19]. Distribution [CLOP19, CEKL15, CL13b, CC19, DGJ11, DH10, DKL12, EPFV18, Khm13, KSZ18, PJS10, PT10, WLCY14, WZ12, XQB13, XZ18]. distribution-free [CEKL15, CC19, EPFV18]. distributions [Atc17, BCL11, DP13, Des15, DSS11, FJ19, GN11, HKKR11, HLLZ19, JY13, KR11a, Khm13, Lee13, Rou10, ULR18, XKB16, FSZ18]. Divergence [PTWZ18, DKR18, JWJW18]. divergences [HW16a]. divergent [ZPW13]. diverging [TZ19a, Wan11, WXQL14]. Divide [BDS19]. Do [KNV15, HSRW19]. does [Ste11]. domain [BLN15, BFFP19, KP11, Rao18, VL18]. Dominating [SSK14]. door [MC15]. dose [FSD^{+}17]. Doubly [TZ19b, AZ12, SW16b]. doubly-censored [SW16b]. drift [GLM18]. driven [GLM18, Mas13]. dropout [Zhe13b]. Dynamic [Bow10, Pen19, SFSL18, SYZL14]. dynamical [HS17, MMNP15]. dynamics [MY10, PPB16].

E-optimal [DG14a]. early [SBC17]. Edge [PY12]. edges [PJS10]. Edgeworth [Bob19, Lah10]. Editor [LK10]. Editorial [Ano16a, Ano16b, Ano17a, Li16b, Ano16c]. effect [NMR17, Ste11]. effective [LH10a]. Effects [Chi11, BL14, CJ10, El 10a, FL12, FJ19, HL18, KGC10, Lok17, MB15, Roh19, SSL19, Zha18]. Efficiency [TS12, BDS19, BHMR14, HL13, JvdG18, QZ10, STRR15]. Efficient [AV13, Ano10, BSY19, CC18a, CHL16, FH14b, Fuh06, Fuh10, JR13, JT14, Jen10, LLX19, MZ13, MNP19, TW15, Vim10, Zha19a, BFR19, DN11, FL16, GLM18, IP11, KLZ11, LLY14, MZZ^{+}19, PO10, SvdAW14, SW16b, WLCY14, XH12, YL14]. EFM [CHZ11]. Efron [MR10]. eigenspaces [FWWZ19]. eigenstructure [WF17]. Eigenvalue [FJ19, BPZ15, HPZ16, Lam16]. eigenvalue-regularized [Lam16]. eigenvalues [BCF18b, Dob17, DGJ18, JLX17, WY17, ZPG18]. eigenvector [XQB13]. eigenvectors [TP18]. eighth [ZPMX11]. elicitability [FZ16, FZ21]. elimination [DGK19]. elliptical [FLW18, HLLZ19]. embedding [WW18]. Empirical [BMW18, BJL15, BSV14, KM10a, MM18, MY10, TW13, BLN15, BD11, CL08, CL10, CTW13, CTW16, CTW18, CCK14b, CL15b, CAGPFBGM19, DP18, DR10, ELL15, KF11, LM12a, LC10, MBM18, NBL13, RS17, SB10, WF17, XQB13, YH12, YL14]. empirical-Bayes [SB10]. emulation [GWB18]. emulators [HQ11]. encounter [PM16b]. Endogeneity [FL14]. enhanced [PHW11]. ensemble [YL11]. ensembles [BGN18, Lop19]. entries [Min18]. entropy [BSY19, Kol11]. envelope [KPS17, LM12c]. epi [BSV14]. epi- [BSV14]. equation [DFS11, DFW19, FDD12, REB18, RBM19, XMW10]. equations [CTW18, Ma12, OPP17, PL11, QZ10]. Equivalence [SSGF13, BGN18, GNZ10, HJY13, Mei11, Rei11, SH14, Wan13]. equivalency [PS10]. ergodic [Mas13, Str18]. ergodicity [DBCD19, JG12, JG13]. Erich [Roj11, vZ11]. error [DZ17, Dic11, FLSZ18, GS10, KSZ18, MM18, NS11, PHW11, RW10b, SF19, WLCY14, XMW10, DKL12]. error-in-variables [DZ17]. errors [CHI13, DPZ16, DPZ19, FR13, HP18, HW19, JMR14, LXQ16, Sch13, TWZ13, TGC^{+}19]. errors-in-variables [HP18]. estimate [JPSZ10, Lok17]. estimates [BMW18, DFW19, Efr18, MS10b]. Estimating [AGZ17, BHMR14, BJR16, Bul14, CLZ16, CD13, CFR12, HL18, Loh15, Lop19, NMR17, BCF18a, BGvZ12, BGvZ21, CJ10, CTW18, HJY16, Ma12, OPP17, TBM11]. Estimation [ACL19, Ano10, ABT11, BFFP19, CEdH11, CLSY18, CTT17, CGC11, DGT16, DPZ12, FRW15, Fuh10, GL12, Jen10, JFF10, Kat12, KGC10, LXQ16, LYW^{+}11, LTG16, LLLT10, MCLX15, Muk16, NW11, RT11b, SD12, SN19, WXZC10, WLLC11, WXQL14, AL19, ATHW12, AS11, BM15, BFL^{+}18, BCW14, BCCW18, Bel10, Bel11, BF12, BSY19, BPD14, BG10, BW19b, BGK10, Bul16, BSW12, CZZ10, CY11, CL11, CY12, CZ12, CMW13, CLZ16, CGW18, CKG19, CCGB17, CV19a, CC11, Cha15a, CXW13, CGR18, CLN19, CH10, CHL16, CF16, CW12, CZW11, CCT17, CCTV18, CGZ18, DL17, DE17, DN11, DM12, EL10, EvdG18, EBGG18, FFN10, FLM11, FXZ14, FL16, FG18a, FLW18, FWWZ19, Fel12, FY10, FR12, Fuh06, GMRZ15, GLZ15, GMZ17, GPPVW12, GPPVW14, GN10, GLM18, GL11, GNZ10, GJW10, Gun12, HKKR11, HW16a, HW16b, HK18]. estimation [HN16, HR11, HKKM10, Jac10, JR13, JT14, JP15, JY11, JMR14, KLZ11, KJF14, KS16, KGS18, KTV17, KM10b, Kol11, KR13, KV17, LUZ19, LW12a, LLC13, Lep13, LS14, Lep15, LW19, Ler11, Li13, LLX19, LZL19, Liu13, LW13, LMP18, MZ13, MMNP15, MMB17, MMB18, MJ15, PL11, PPB16, Pen17, Pen19, QZ10, QP16, QC12, QWW16, RSZZ15, RBS10, RT11a, RPF13, RLM^{+}17, Roč18, Rou10, RCL12, SW13, SW10a, SY12, SvdAW14, SS11b, SO18, SW10b, Spo12, Str18, SW16b, TZ19b, TWZ13, TPN19, Tib14, Tod15, TGC^{+}19, Uhl12, VR19, Ver10, Vim10, VL13, WZ10, WLCY14, WBS16, WZ12, WZ18, WY19, XW12, XKB16, XMW10, XZC12, XZ12, YZ16, ZW15, ZBC19, ZL11b, Zho14a, ZBFL18]. estimations [CL13c]. estimator [BB18b, DMOvH11, EKS12, Gao13, HMLZ11, Jan14, Lam16, PO10, RS17, SBW10, Zha13]. estimators [ACF12, Atc11, BB18a, Bel18a, Bel18b, BDV11, BSW11, CL13b, DS12, DLLO16, DKL12, FJ19, FdH15, FJ10, GZ11, Gro14, HW19, JT15, Jir12, LPS08, LP10, Loh17, LN11, LM19, Min18, PJS10, QCN18, TA17, YL11, ZL11c]. Euclidean [CC19, PM19]. Evaluating [LGS11]. event [JLX17]. Exact [BRS19, CR13, KP19, LSST16, MR10, MH13, MJ15, RV11, ULR18, GZ16a, SH15]. exceedance [JM18b]. excess [DW19a]. exchangeable [CW14, VR19]. expansions [Lah10, Li13, WCT11]. experimental [CK15]. experiments [AGL14, BB16a, BB10, CCC19, God18, HQ11, LBST10, LPK17, MR12]. Exponential [DBCD19, RT11a, BCZ10, CD13, DPZ12, HF19, JWJW18, Muk16, SR13, YLZ16, ZL11c]. exponentially [DGP18]. exposures [VR12]. expression [LS10, NC10]. Extended [CD17]. extension [YS12]. exterior [ABT11]. Extremal [Zha18, BB18b, JLX17]. Extreme [WY17, BDV11, CEdH11, ELL15, FdH15]. extreme-value [BDV11]. extremes [FdHM15]. extremity [ELL15].

faces [WRM19]. Factor [KS11, LY12, BL12, FLM11, FLW16, FLW18, PBPD14]. factorial [CC18b, TXL12]. factorization [BQS18]. factors [CCC19, LY12, LLL19, LM12d, LWY17, MG11, MGC10]. fails [BSV14]. failure [LJ18]. faint [QCN18]. Faithful [XCL16]. faithfulness [URBY13]. False [SBC17, BC15, DR15, GHS14, JM18b, Liu13, NR12, PHW11, RV11]. families [BCZ10, DPZ12, HF19, LLY18, Muk16]. family [HIT19, JWJW18, PHW11, XKB16, FLRB16]. family-wise [PHW11, FLRB16]. familywise [GS10]. Fast [ANW12a, Jin15, RBM19, SS13, YPW17, Fry18, Wal10]. favorable [RV11]. FDR [ZFY11]. feasible [CL15a]. Feature [DGK19, CTW13, CTW16, Jam17, VAC17, YDS16]. features [ACV16, BB15, CZ16, JW16a, JW16b, Nad16, SHMD13, ST16b]. Fiducial [TL13, CH12, WH19]. field [KLN12, Loh15]. fields [AST12, And10, CCC19, CS17, LX12, Mas13, MH14, Ver10]. filament [QP16]. filling [HQ14, SLQ14]. filter [BC19, MZ15, RBWJ19]. filtering [GLM18, IBAK11, ST19, Tib14]. filters [CL13a, WL14]. financial [JKL12, WZ10]. Finding [GL18]. Finite [BHPZ19, ZBFL18, ASJ11, BYZ10, GKM16, HK18, HN16, HIT19, Lv13, Ngu13, QZ10, TA17, Spo12]. finite-sample [QZ10, TA17]. first [Akr16, DS11, JPSZ10]. Fisher [CI11, WY17]. fit [Bul17, CEKL15, CAGPFBGM19, KLN12, KSZ18, Lei16, TWT17, VV10, WW18]. fitting [SL19]. Fixed [KRS15, ZS13, BFFP19, MH13, ZPW13]. fixed-margin [MH13]. Fixed-smoothing [ZS13]. Flexible [DE17, LMP12b, LBST10]. flow [CLM16]. fMRI [CK15, LPK17]. Focused [ZL11a]. folded [FXZ14, LYL16]. folding [LKA10]. forecasting [GRSP15, LMNP15, LMNP17]. forecasts [LGS11, SSK14]. forests [ATW19, SBV15]. form [Li13, Zhu19]. forms [AW12, DE17]. formula [AST12, TLT16]. formulas [ULR18]. Formulation [LLC13]. Fourier [PT13, Rao18]. Fractals [MS10a]. fractional [Bel10, BPY19, BF18, HR11, SH14, TXL12]. fractions [Pho12, ZPMX11]. frailty [SW16b]. framework [BCCW18, NZL17, ZCL16]. Fréchet [BG13, PM19]. free [CEKL15, CC19, EPFV18, HWH13a, HWH13b, Khm13, MZ15, YDS16]. freedom [TT12a]. Frequency [VL18, ASJ10, BLN15, BF18, CEL16, CGC11, JKL12, KP11, Koi19, KLJ15, Kon18, Mas13, WZ10, XZ18]. Frequentist [Cas15, Gho15, LM15, Nic15, Rou15, SvdVvZ15a, SvdVvZ15b]. Fuh [Jen10]. full [BD14, TW13]. Fully [MG11, Ste15]. function [AGZ17, Arm15, BGvZ12, BGvZ21, Bel11, BPD14, BDV11, CY11, CL13c, CF16, EBGG18, HL13, JMR14, LZL19, LMP18, PT13, SS11b, TBM11, WZ12]. Functional [BLRG17, DP19, FJR15, JW11, KPS16, PM16a, AGR18, AW12, BBTW11, BCCW18, CY11, CL11, CDH13, CHM11, CM12, CJ12, CAGPFBGM19, DM18, DH12a, DPZ12, HMV13, HK10, HKN18, JW10, JMW15, Kat12, Lep15, LH10a, LH10b, LS17, LMY18, LY19, LLY14, MS10a, Mei11, Mül16, Pap18, PS10, QWW16, RLM^{+}17, SC15, YC10, ZW16, ZLK12]. functional/longitudinal [LH10b]. functionals [CR15, CCT17, CCTV18, DR10, FRW15, JR13, Jan14, Koi19, LLX19, Pen17, RR12]. functions [ACL19, BFFP19, BSV14, CLX13, wCM19, DGJ11, DH10, GXZ18, GJ15, Gun12, HH13, JPSZ10, LX12, PHW11, PM16a, Zha19b]. Further [GHS14]. Fused [JMW15, MZ15].

Gamma [NC10]. Gamma-based [NC10]. GARCH [FZ13, ZL11c]. GARCH/IGARCH [ZL11c]. Garza [DM11, Yan10]. Gaussian [AST12, And10, BHPZ19, BPD14, Bon11, BF18, BGN18, CMZ19, CV19a, CN13, CCC19, Che18, CCK13, CCK14b, DLLO16, DKZ13, FJY13, FL16, GL18, GL12, GNZ10, GWB18, HPWW11, HR11, JM18a, JN18, KvdVvZ11, Koi19, KMH14, LUZ19, LHY^{+}12, LX12, Liu13, Liu17, Loh15, LM19, Mas13, Men17, Min18, MJ15, PO10, Ray17, RSZZ15, RCL12, SvdAW14, SW19, STD10, Uhl12, Ver10, VV10, ZW17]. GEE [CLLW15, Wan11]. Gemini [Zho14a]. gene [CQ10, LS10, NC10]. gene-expression [NC10]. gene-set [CQ10]. General [BB18a, LM12c, VR12, ZBC19, BPZ15, CLR16, CL13a, CH10, DPZ12, DMOvH11, GKM16, HWCS19, HQ14, JL14, LLC13, NL17, RH12, Yu10]. generalised [LKZ15]. generalizations [GZ16b]. Generalized [ATW19, Che19c, CH12, GX14, LW13, RW10a, WWG19, AcL17, BFFP19, CHI14, FS10, HBZO19, Jam17, JMW15, LMP12b, Ma12, MCLX15, MC15, MG11, OPP17, QZ10, Rig12, RW10b, SC15, SSCL19, SF19, TT11, WLLC11, WXQL14, WH19, ZL11a]. generated [MRS12]. generic [FDD12]. genewise [FFN10]. genome [JLP^{+}16, RBS10]. genome-wide [JLP^{+}16]. genomic [BS14, WR13]. geodesic [Huc11]. Geometric [CLR16, BM17, CCGB17, JG12, JG13, LZ11, SC12]. Geometry [SvdAW14, Uhl12, URBY13, CTP19, HW16b, LM12d, SWY15, WWG19]. geostatistical [CHI14]. geostatistics [CZW11]. Getoor [ASJ12]. Gibbs [DL17, RH12]. glimpse [Roj11]. GLMM [Jia13]. Global [ACCP11, CEL16, DFS11, KS16, LYL16, LN11, MMY18, ZL11c, ANW12a, LS13, Men17, PP14, SC13]. Globally [MS10b, ZPH15, WW18]. good [WXX18]. Goodness [CAGPFBGM19, KLN12, KSZ18, VV10, Bul17, CEKL15, Lei16, TWT17]. Goodness-of-fit [CAGPFBGM19, KSZ18, VV10, Bul17, CEKL15, Lei16, TWT17]. Gradient [WZ18, ANW12a, CL15b]. Gradient-based [WZ18]. gradients [TA17]. gradual [VD15]. Graph [CZ15b, EW16, Zho14a, BD15, CKG19, CD13, KR11a, SR13, YLZ16]. Graph-based [CZ15b]. graphic [FJY13]. graphical [AV13, BK18, CS12, CPW12a, CPW12b, DKZ13, FL16, GL12, GT12, KMH14, LM12b, LS18, LHY^{+}12, Liu13, Liu17, LW13, NLL19, RZ12, RSZZ15, Røy12, ST16a, STD10, Uhl12, ULR18, Wai12, XZ12, Yua12]. graphon [GLZ15, KTV17, Pen19]. graphons [Cho17]. graphs [Bow10, CMKR12, ER14, FG18a, HJY13, KR13, MMS18, Sad16, TSP13, TP18, VR19, vdGB13]. greater [CW12]. greedy [DRZ12, Fry18]. Grenander [DKL12, Jan14, SBW10]. Grenander-type [DKL12]. grid [TBK12]. group [BO17, DH12b, HZ10, LPvdGT11]. Groups [DKZ13]. grows [MGC10]. growth [Huc11]. guaranteed [GRD13]. guarantees [AV13, BWY17, JKNP12, LW12b, QM11].

Half [FDD12]. Half-trek [FDD12]. Halfspace [PV18, HPŠ10a]. Hall [Che16, CF16, Del16, Mül16, Sam16]. Hamming [BNST18]. Harris [LTG16]. Hastings [DR11]. Hausdorff [GPPVW12]. Hawkes [RBS10]. Haystack [CvdV12]. hazards [BFJ11]. heavy [BJL15, HW19, Min18]. heavy-tailed [BJL15, HW19, Min18]. Hellinger [HW17]. help [HSRW19]. Hermite [CL11]. heterogeneous [HWH13a, HWH13b, ZCL16]. heteroscedasticity [Zhu19]. heuristic [DR15]. hidden [CL13a, Chi11, DMOvH11, DM12, RBM19]. hierarchical [BTT13, CLOP19, HA13, LM12d, WRM19]. hierarchy [ST16a]. High [ATHW12, DW18, El 10a, EH19b, FLM11, GZ16b, HLLZ19, LHY^{+}12, LW12b, NHM18, RWL10, SFSL18, ZLLW13, ZPH18, ANW12a, ANW12b, ASJ10, AcL17, ACV16, ABR10a, ABR10b, Atc17, BL12, BCF18a, BCF18b, BW19a, BHPZ19, BC19, BM15, BFL^{+}18, BC11a, BR13, BB16b, BW19b, BJNP13, BF18, BPE14, BSW11, BSW12, CZ16, CLR16, CG17, CG18, CZ18, CMZ19, CKG19, CTP19, CC17, CTW18, CQ10, CXW13, Che18, CL18, CLZ19a, CHLP14, CCK13, CMKR12, CD12, CGC11, CFR12, CF19, CGZ18, CPV17, DGK19, DZ17, EH19a, FL14, FKLZ15, FJ19, Gol10, HP10, HWH13a, HWH13b, HMLZ11, JvdG18, JJ12, JY13, JLP^{+}16, JW16a, JW16b, JKW17, JKL12, Jir15, KP15, KS11, Koi19, KLJ15, KLFL17, Kon18, KV10, LM12a, LY12, LZW10, LLL19, Lee13, LD18, LH10a, LC12, LKZ15, LZ17, LLYY19, LZL19, LZL18, LC10, Loh17, Lv13]. high [MCLX15, Mas13, MPL15, Nad16, NMR17, NW11, NL19, NL17, NZL17, OWJ11, OMH13, OMH14, QC12, RYC19, RWG19, RLM^{+}17, RT11b, RCY11, Sam16, SWDW11, SD12, SW16a, SSCL19, SL19, ST16b, TZ19b, TWZ13, VV10, VL13, WZ10, WKL13, WF17, WH19, XZ18, XCL16, XZ12, YP15, YT15, YWJ16, YZ16, Zha13, ZW17, ZPG18, ZYS18, ZL11b, ZPH15, ZBY15, ZCCL19, ZCGZ19, ZCX13, ZLST17, Zhu17]. High-dimensional [ATHW12, DW18, EH19b, FLM11, GZ16b, HLLZ19, LHY^{+}12, LW12b, NHM18, RWL10, SFSL18, ZLLW13, ANW12a, AcL17, Atc17, BCF18a, BCF18b, BW19a, BHPZ19, BC19, BM15, BC11a, BW19b, BJNP13, BPE14, BSW11, BSW12, CLR16, CG17, CG18, CZ18, CMZ19, CKG19, CTP19, CC17, CTW18, CQ10, CXW13, Che18, CL18, CCK13, CMKR12, CFR12, CF19, CGZ18, CPV17, DZ17, EH19a, FKLZ15, FJ19, Gol10, HP10, HWH13a, HWH13b, HMLZ11, JvdG18, JJ12, JY13, JLP^{+}16, KP15, KS11, KLFL17, KV10, LY12, LZW10, LLL19, Lee13, LH10a, LC12, LZ17, LLYY19, LZL19, Loh17, MCLX15, MPL15, NMR17, NW11, NL19, NZL17, OWJ11, OMH13, QC12, RYC19, RWG19, RLM^{+}17, RT11b, RCY11, Sam16, SD12, SL19, TZ19b, TWZ13, VV10, WH19, XZ18, XCL16, XZ12, YWJ16, Zha13, ZPG18, ZBY15, ZCCL19, ZCGZ19, ZCX13, ZLST17, Zhu17]. High-dimensionality [El 10a]. high-frequency [ASJ10, BF18, JKL12, Koi19, KLJ15, Kon18, WZ10, XZ18]. high-order [LC10]. Higher [Dic11, FZ16, ACCP11, CL13b, HJ10, Loh15, WMZ18, ZCX13, FZ21]. higher-order [Loh15]. highest [SW10a]. highly [JFF10]. Hilbert [AW12, CCGB17, Lev15, LLLH18, PM16a, YC10]. hold [Ste11]. homodyne [LMP18]. homogeneous [DH12b]. honest [CCK14a]. Hotelling [LS13]. Huber [CGR18]. hull [PJS10]. hybrid [NHM18, TBM11, ZvdAW19]. hyperbolic [HKKM10]. hypercube [WXX18]. hypergraph [GD17]. hypographs [BSV14]. hypotheses [CL11, NS11]. Hypothesis [BW19a, MPL15, ZCCL19, ASFJ10, Li14, NJ13, NL17, PV17, SSGF13, SSCL19, WZHO17, WWG19, ZBY15].

I-LAMM [FLSZ18]. Identifiability [Xu17, BS14, DFS11, FDD12, HK18]. identifiable [HN16]. Identification [ZAL17, CHLP14, KLZ16, SQ17]. Identifying [ASJ12, BYZ10, LWY17, YS12]. idiosyncratic [Kon18]. IGARCH [ZL11c]. II [ABR10b]. II. [BB10]. ill [ISS12, Pen16]. ill-posed [ISS12, Pen16]. image [AL11, LG17]. imaging [ZFY11]. imbalanced [FH14b]. Immigrated [ZHCC11]. Impact [JY16, KPS16, MS10a]. Impacts [Lv13]. impedance [HKKM10]. imperfect [TW15]. implied [QT19]. implies [EKDN18]. importance [EL10]. improper [MH11, RH12]. improve [MR12]. Improved [BLT18, CW12]. in-sample [LMNP15, LMNP17]. incoherence [LW17]. Inconsistency [SBW10]. increasing [Bon11, JP15, YLZ16]. Independence [YP15, ABFRB15, BDP19, Bow10, CTW13, CTW16, CL18, CHLP14, CD17, FS10, Hua10a, Lep13, LD18, Røy12]. Independent [SY12, CZC10, IP11]. index [BB18b, CHM11, CLLW15, CHZ11, DGT16, DZ14, JW11, JL14, JMW15, Kau18, LS14, LYW^{+}11, LLY18, LLLT10, LY11, MH16, WXZC10]. Indian [Jam17]. indices [ASJ12]. individualized [QM11]. inequalities [ACL19, Bel18b, DS12, GL11, Gol10, HSY^{+}13, KTV17, LM12c, LW19, LPvdGT11, LLLH18]. inequality [HS17, LJ10]. Inference [DW19b, FZ13, KRD10, LY12, MH16, PV17, TLT16, ZW12, Zho14b, ASW13, BC19, Bel11, BBB^{+}13, BV13, CLR16, CGT16, CZvdL17, CH12, CD17, DP13, DHSM18, DP18, Efr18, FH14a, GXZ18, HPWW11, HN11, HBT16, Huc11, HE18, IP11, KV10, LSST16, LV15, LW10, LPvdGT11, LvdL16, MCLX15, MZZ^{+}19, MNP19, NLL19, Rao18, Rei11, RWG19, REB18, RBM19, SC13, SC15, SQ17, ST16a, TBK12, TL13, TT18, TRTW18, URBY13, VL18, VCC19, WH19, WM10, XZ18, ZJS11, ZBC19, ZDMD18, Zho10, Zhu19, dJvZ10]. infinite [ASJ11, CGT16, JT14, Ngu13]. infinite-variance [CGT16]. infinity [CTP19]. influence [ZLLW13]. influences [JBGWS13]. Influential [ACV16, CZ16, JW16a, JW16b, Nad16, ST16b]. Information [HW16b, MP17, AL11, CI11, CHI14, DKR18, El 10b, KNV15, LPS08, LP10, ZL11a]. Information-regret [MP17]. inhomogeneous [MDO14, MB15, Pen13, Tru19]. inner [LMY18]. Innovated [FKLZ15, FL16, HJ10]. instances [Zho14a]. instrumental [Kat13]. integer [SH15]. integrals [JY11]. integrands [Dic11]. integrated [DGT16, JT14, LLX19, ZL11b]. interacting [BDD10, FMP11]. Interaction [KLFL17, BD14, FKLZ15]. interactions [BTT13, MCLX15, VR12]. interactive [BL14]. interest [LLY14]. interference [LZA15, Zhe15, ZAL17]. intermediate [RCL12]. Intermittent [BMBM15]. interval [CGC11, Gro14, LK10, Wan10, YL14, ZZX18]. intervals [BLP19, BK18, CLX13, CG17, GJ15]. interventions [NMR17, RBM19, ST16a]. Intrinsic [Huc11, LY19, HBT16]. Introduction [BC11b]. invariance [HBZO19]. Invariant [EJ10, Str18, Vim10]. Inverse [JYW14, JW15, CLR16, ISS12, JL14, KM10a, KvdVvZ11, LZL18, Mei11, Pen13, PWM18, TBM11, YDS16]. inverses [LW13]. irregularly [BLN15, Loh15]. Ising [ATHW12, BRS19, MMY18, NL19, RWL10, XZC12]. Isotonic [HWCS19, CGS15]. Itô [TWZ13]. Iterated [IBAK11]. iterative [Ma13].

jack [OPP17]. jack-knife [OPP17]. jackknife [LLX19]. Joint [BSW12, CS15, NMR17, WZ12]. Jump [GLM18, Tod15, Bul16, CMW14, JKL12, KLJ15, TT12b, Tod17]. jumps [ASJ11, JT14].

Kac [TLT16]. Kantorovich [CGHH17]. Kendall [Bao19, BK18]. Kernel [Atc11, BGK10, AW12, ACF12, CL15b, DGK19, El 10b, GL11, JMW15, El 10c, KY10, LLLH18, MN10, SS13, YC10]. kernel-spline [JMW15]. kernelized [SWY15]. kernels [KR13, WCT11, YPW17]. Khinchine [Bob19]. kinematic [AST12]. kinetic [Bow10]. knife [OPP17]. knockoff [BC19]. knockoffs [BC15]. knowledge [RBWJ19]. Kolmogorov [FG18b, MZ15]. Kullback [CGM^{+}13, Rig12].

labels [AGZ17]. LAMM [FLSZ18]. landscape [MBM18]. Laplace [DGP18, TT12b]. Laplacian [EW16, HMLZ11, TP18]. Large [ACF12, BB16b, FLW18, Sae19, ACBL13, BLPZ15, BDP19, CJ10, CKW^{+}16, CLR17, DB16, FL16, GL18, GZ11, HQ11, HJY16, Kon18, LW12a, LWY17, Ma12, MB15, RSZZ15, SN19, WY17, WZY18, XQB13, ZFY11]. large-dimensional [LW12a, WY17]. Large-sample [ACF12]. large-scale [CJ10, CKW^{+}16, HQ11, MB15, ZFY11]. largest [BPZ15, HPZ16, ZPG18]. Lasso [DGL19, LTTT14a, BTT13, BMvdG14, BB14, CY14, FK14, HSY^{+}13, JM18a, LSST16, LTTT14b, LTTT14c, LZ14, TT11, TT12a, Was14, CL13b, MMB17, MMB18, BLT18, BCW14, SBC17]. Latent [CS12, CPW12a, CPW12b, GT12, LM12b, RZ12, Wai12, Yua12, AV13, BJR16, CMKR12, Jam17, Ngu13, TSP13, Xu17]. latent-structure [BJR16]. latin [WXX18]. lattice [WXX18]. lava [CHL17]. law [HPZ16, LAP15]. laws [BD15, CJ11]. leaf [Huc11]. lean [RWG19]. Learning [AV13, CMKR12, BK19, CZvdL17, DRZ12, FLSZ18, GK12, HS17, Han11, KY10, KP19, LR14, LYL16, LSR^{+}17, MN10, NHM18, SFSL18, SS13, WLZ14, ZLZ18]. Least [Bel18b, BM14, AC11, Cha14, CF19, DH12a, GZ16b, HW19, Jac10, RV11, SS11b]. leave [XH12]. leave-subject-out [XH12]. Lecture [Ste11]. left [SW12]. left-truncated [SW12]. Lehmann [BC11b, Roj11, vZ11]. Leibler [CGM^{+}13, Rig12]. lens [BKM16]. Leo [Roj11]. Letter [LK10]. level [CC18b, STY14, SLL11]. levels [FG18b]. Lévy [Bel10, Bel11, BV13, Bul14, CGC11, GLM18, Mas13, PM16b, QT19, WCT11]. Lévy-driven [GLM18]. life [Roj11]. like [KR11b]. Likelihood [Ano10, Fuh10, Jen10, TBK12, WB17, ACBL13, BLN15, BB18b, BCCZ13, CC11, CTW13, CTW16, CTW18, CZW11, DL17, DMOvH11, DM12, DFW19, DFKP19, Efr18, FY10, FR12, Fuh06, GJW10, Gro14, HR11, JY13, KR11b, LM12a, LUZ19, Li13, LC10, Mas13, MMNP15, NZL17, NBL13, QCN18, RS17, SW13, SY12, TGC^{+}19, TW13, Uhl12, WWG19, XZC12, YFG13, YH12, ZL11c, ZYFW14, vdGB13]. Likelihood-based [WB17]. lilkelihood [RPF13]. Limit [DR10, JPV10, TP18, Bao19, BLRG17, DKL12, EH19a, HQ14, JY13, KNV15]. limitations [WMZ18]. limited [SO18]. Limiting [BCF18b, CJ11]. Linear [SSCL19, AcL17, AC11, CLR16, CG17, CG18, CSHvdV15, CGT16, CC18a, CLLW15, CS15, CH12, CCT17, CCTV18, CJ12, CAGPFBGM19, DP13, DPZ13, DN11, DGT16, DFS11, El 10a, FS10, FL12, FR12, FDD12, FGM17, GRSP15, Gol10, GH18, HMV13, Jan14, JP15, JMW15, JBD17, JN18, KLZ11, Kat12, KPS16, KSZ18, LYW^{+}11, LAL11, LZL19, LLLT10, LAP15, Ma12, MS10a, Mei11, Pen13, Pen16, Pen17, Rig12, RR12, RH12, REB18, RBM19, SC15, SWDW11, SD12, SW16a, TGC^{+}19, VV10, WXZC10, WLLC11, WXQL14, WH19, WW18, YC10, ZL11a, Zha13, ZCL16, ZBY15, ZCCL19]. link [CHM11]. Lipschitz [ACL19]. Local [ATHW12, BHMR14, BF18, BGN18, CTW16, EG12, EBGG18, FH14b, HBT16, SO18, SC13, Tru19, BW19a, Bel17, CS17, DLP12, FG18b, PDL12, Røy12, SGA11, YFG13, ZL11c, Men17]. localization [CLR17]. Localized [KNP11]. Locally [PR19, DM11, HYS15, PO10, RD19, Vog12, VD15, WW18, YS12]. location [BGvZ12, BGvZ21, Des15, dJvZ10]. location-scale [Des15, dJvZ10]. log [Des15, DW19b, DSS11, FR12, JBD17, KS16, KGS18, MM18]. log-concave [DW19b, DSS11, KS16, KGS18]. log-linear [FR12, JBD17]. log-regularly [Des15]. logistic [RWL10]. loglinear [LM12d]. long [BT17, CL08, CL10, CHI13, CGC11, LLBM^{+}11, PO10, RCL12]. long-memory [CHI13, PO10]. long-range [LLBM^{+}11]. longitudinal [CZ10, CLLW15, CHLP14, CHL16, JW10, JW11, JYW14, JW15, LH10b, MY10, SW12]. longitudinal/clustered [CHL16]. loopy [AV13]. loss [ACL19, BNST18, GPPVW12, Han19, HL13, JN18, Lep13, Zha19b]. losses [BJL15, MBM18]. Low [KR13, EvdG18, FH14a, Kol11, KLT11, LLV16, Lee13, NW11, RT11b, Zha19a]. low-dimensional [Lee13]. low-rank [EvdG18, FH14a, Kol11, KLT11, LLV16, NW11, RT11b, Zha19a]. lower [KP19]. lowest [PR16]. LWF [Sad16].

machine [KP19]. machines [DGK19, LAL11]. made [EW16]. main [Del16]. majorization [Yen11]. Manifold [GPPVW12, AL19, CV19b, CM12, WW18, YD16]. manifolds [ABT11, DM18, EH19a, Pen18]. MANOVA [BCF18b]. many [BCCW18, SSK14]. Marcenko [LAP15]. margin [MH13]. Marginal [CTW13, CTW16, RS17, STRR15, WRM19, YDS16]. Marginalization [Sad16]. Margins [Eva18, wCM19]. Markov [ASFJ10, Atc11, Bac11, BFT13, BDD10, BD15, CL13a, CDO11, CMW14, Chi11, DMOvH11, DM12, FLS^{+}17, FJ10, FMP11, HW16b, HJY13, KLN12, LS18, LTG16, LSR^{+}17, MDO14, Tru19, WR13]. Markovian [ER14]. Markowitz [El 10a]. mass [AGZ17, DW19a]. massive [LZL19, PBPD14, ZCL16]. matching [WY19]. matrices [BPZ15, BB16b, BSW11, BSW12, CJ11, CW19, DMO08, DG12, El 10b, FRW15, HPZ16, HLLZ19, HJY16, JLX17, JPSZ10, El 10c, LW12a, LC12, LW13, MH13, NW11, PV18, PBPD14, PY12, QC12, RT11b, SD12, SN19, WY17, ZL11b, ZBY15, ZCGZ19]. Matrix [CZ15a, Cha15a, ANW12b, BLPZ15, BHMR14, BW19b, CZZ10, CY12, CZ12, CLZ16, CLR17, CXW13, CGR18, CTT17, DP19, DG14b, EvdG18, FLM11, FH14a, GL18, JP15, Kol11, KLT11, Lam16, LKA10, LWY17, Min18, PWBM18, RT10, TWZ13, WZ10, Wan13, XQB13, XZ18, XW12, ZPW13, ZCCL19, Zho14a]. matrix- [LKA10]. matter [Dob17]. maxima [CS17, CCK13, FdH15, Koi19, SGA11]. maximal [Hua10a]. Maximin [MB15, WXX18]. maximization [CLX18]. Maximum [FY10, FR12, GJW10, Gro14, LUZ19, Li13, RPF13, TGC^{+}19, BGvZ12, BGvZ21, BB18b, BCCZ13, CC11, CZW11, DMOvH11, DM12, DFW19, DFKP19, Efr18, FSZ18, HR11, LS10, MMNP15, QCN18, RS17, SY12, Uhl12, ZL11c, ZYFW14, vdGB13]. Maximum-likelihood [Li13]. MCMC [DP18, HJY13, Lia10]. mean [CY11, CZvdL17, CHI13, CW12, DW19a, DLLO16, Dic11, Huc11, LX12, LvdL16, LM19, Min18, XKB16, ZYS18]. means [BG13, CC17, CLZ19a, FJ10, GL12, LZ17, NC10, ZBC19, ZCX13]. measure [ZLLW13]. measured [JMW15]. measurement [KSZ18, TWZ13, XMW10]. measurements [CKW^{+}16, ZLST17]. measures [BKRS19, BV13, Che19c, Ngu13]. Measuring [ZZX18]. median [CCGB17]. mediation [TS12]. meets [BLT18]. memory [BT17, CL08, CL10, CHI13, CLZ19b, PO10, RCL12]. merged [Sae19]. Merging [PANS14]. message [MMB17, MMB18]. method [BLN15, BHMR14, BCL11, FKLX15, FdH15, GZ11, HH13, LM12a, LM18, LBST10, MZ15]. Methodology [DH12a]. methods [ANW12a, ACBL13, AV13, BL14, BT17, BDD10, BD11, Cha15b, EW16, GCW19, KLZ11, MDO14, SHMD13, SR14, WR13]. metrics [GN11]. Metropolis [JG13, DR11, JG12, STRR15]. microarray [FFN10, God18]. middle [CW19]. middle-scale [CW19]. Mimicking [Lok17]. Minimax [BG13, BJNP13, CG17, CCT17, DG14b, GMRZ15, HKKR11, HMV13, ISS12, LLL19, Pen17, RLM^{+}17, VL13, YT15, YZ16, ZZ16, BW19a, GL11, HK18, KP15, MW15, MJ15, SC16, WWG19, Zha10, Zha13]. Minimax-optimal [YT15]. minimization [BJL15, CL15b, KM10a, LZ11, Yen11]. minimum [VL18, YL11]. minors [DMO08, DG12]. mis [SBK16]. mis-specification [SBK16]. Mises [BK12, BG14, Bon11, CN13, CN14, CR15, Ray17, RR12]. missing [CZ10, LW12b]. misspecification [Jan14, SZ15]. misspecified [DM12, HIT19, JLP^{+}16, Rig12]. mixed [CH12, ER14, FL12, JLP^{+}16, RH12, SH15, STY14, SLL11]. mixed-level [SLL11]. mixing [HS17, Ler11, Ngu13]. Mixture [LMY18, BTWB10, HK18, Ngu13, VS19, VAC17, dJvZ10]. mixtures [CMZ19, HKKR11, HN16, MH11, Nor10, Rou10]. MLE [Jia13]. Möbius [HKKM10]. modal [CGTW16]. mode [DW19b, PV17]. Model [LKZ15, ASJ10, AGR18, AL18, AcL17, BLP19, BBFGD12, BG10, BFJ11, BMW18, BD15, CS12, CV19a, CPW12a, CPW12b, CHI14, CL13c, CGR18, Chi11, CAGPFBGM19, Des15, DN11, DGJ18, EJ10, FJY13, GD17, GT12, GJW10, Gro14, HWH13a, HWH13b, HMV13, HL13, HIT19, HSY^{+}13, JMW15, JLP^{+}16, KJ16, KP15, KP19, LM12b, LLL19, LS14, LW19, Ler11, LYW^{+}11, LZA15, LJ18, LY11, Liu13, MZ15, MM18, MGC10, MMS18, PP14, QWW16, RWL10, RZ12, RPF13, RLM^{+}17, SQ17, SZ15, SW19, SYZL14, TZ19a, TRTW18, Wai12, WXZC10, WXQL14, WB17, WH19, YDS16, Yua12, ZL11a, ZW15, Zhe13a, ZAL17, ZLK12]. model-free [HWH13a, HWH13b, MZ15, YDS16]. Modeling [HM10, JKL12, SW12, CMW14, CF16, Jam17, LY12]. Models [Ano10, Fuh10, Jen10, ASFJ10, ATHW12, AV13, AcL17, BL12, BL14, BD14, BB16a, BK18, BM15, BFL^{+}18, BC11a, BF12, BS14, BCL11, BW19b, BG14, BJR16, Bow10, BD13, BPE14, Bul14, BTWB10, CKG19, CR15, CL13a, CLSY18, CTW16, CD13, CHM11, CLLW15, CLX18, CS15, CHL16, CH12, CHZ11, DMS12, DS13, DPZ13, DG14a, DPZ19, DN11, DGT16, DMOvH11, DM12, DKZ13, DFS11, DFW19, DML10, FS10, FLM11, FL12, FL16, FLW16, FLW18, FJ19, FR12, FDD12, FZ13, Fuh06, GMZZ18, GL12, Gol10, HP18, HA13, HIT19, HR11, HHW10, IP11, Jac10, JvdG18, JFF10, JW11, JL14, JBD17, KLZ11, KLN12, KRD10, Kat13, KR11a, KPS17, KTV17, KS11, KSZ18, KMH14, KGC10, LS18, LUZ19, LMP10, LMP12a, LMP12b, LLL19, LR15, LXQ16, Lei16, LM12d, LKZ15, LZL19, LLLT10, LJ10, LJ11, LMY19]. models [LHY^{+}12, Liu17, LW13, Ma12, MCLX15, MH16, MW16, MJ15, MMY18, NLL19, NL19, Ngu13, NL17, PBPD14, PS10, Pen19, PWBM18, RSZZ15, Rig12, RH12, REB18, RBM19, Røy12, ST19, SvdAW14, SR13, SC13, SC15, SD12, SSCL19, SBK16, SL19, SW16b, STD10, TW15, Uhl12, ULR18, VR12, VS19, VL18, VV10, VAC17, Vim10, WLLC11, WP12, WXQL14, WB17, WRM19, Xu17, XMW10, XZC12, XZ12, YLZ16, YS12, YZ16, ZL11a, ZHCC11, ZW12, ZW15, ZZ16, ZLZ12, Zhe15, ZL11c, Zhu17, Zhu19]. Moderate [Gao13, GXZ18, CSWX16, GZ11, LS10, LS13]. moderately [DGK19]. modern [BM14, BKM16]. modified [GZ16a]. modularity [Bow10, CLX18]. modulated [DZ14]. Moment [CHI13, CI11, DS13, LJ10, LJ11, WY19]. Moments [DMO08, DG12, BHMR14, BCL11, BMBM15]. Monge [CGHH17]. monitoring [PR12, ZH10]. Monotone [AS11, CLSY18, CPV17, GXZ18, GJ15, PPB16, WZ12]. Monotonic [Yu10]. Monro [BF12]. Monte [Atc11, BDD10, CDO11, FJ10, FMP11, GRD13, MDO14]. most [Joh13]. motif [WR13]. motion [ASJ10]. moving [DS11, JPV10]. MSE [NK18]. MSE-optimal [NK18]. Multi [AZ12, LJ18, BPD14, CW15, Cha17, CC18b, KLFL17, PR13, XS13]. multi-armed [PR13]. multi-bandwidth [BPD14]. Multi-objective [AZ12]. multi-response [KLFL17]. multi-sample [CW15]. multi-sensor [XS13]. multi-stratum [CC18b]. multi-stream [Cha17]. Multi-threshold [LJ18]. Multiclass [DKR18]. multidimensional [LPS08, LP10]. Multilayer [BQS18]. multinomial [WZY18]. multinomials [BW19a]. Multiple [ABR10b, CS17, SGA11, ZFY11, ACCP11, BCFG11, CJ10, wCM19, CHM11, Chi11, DZ14, FLRB16, Fry14, Fry18, HPŠ10a, HL18, KY10, LZ11, Li16a, Liu17, MMB11, NS11, PHW11, RBWJ19, Sae19, SF19, SS13, TS12, WZHO17, ZBFL18, ZYFW14]. multiple-output [HPŠ10a]. multiplicative [ACF12]. multiplicity [SB10]. multiplier [CCK13]. multiresponse [RYC19]. Multiscale [BHM18, PWM18, SHMD13, LLX19]. multispiked [OMH14]. multistage [Cha15b]. multitiered [BB16a]. multivariance [BKRS19]. Multivariate [HPŠ10a, Lep13, ZLK12, BGvZ12, BGvZ21, BW11, BSY19, BW19b, BJR16, CEdH11, CW12, CC19, DP13, DGH^{+}19, FdHM15, JP15, OWJ11, SS11b, SW10b, YG16].

Natural [DMS13b]. Near [Bul16, JN18, NW11, PV17]. Near-optimal [Bul16]. Near-optimality [JN18]. nearest [BSY19, BCR10, Che19a, GKM16, Sam12]. Nearly [Por12, Zha10, Zha13]. necessary [ASJ10]. Needles [CvdV12]. Negative [GCW19]. neighbor [BCR10, GKM16]. neighbors [Che19a]. neighbour [BSY19, Sam12]. neighbourhoods [HW17]. nested [HE18, MM18, SLQ14]. net [BM17]. nets [Dic11]. Network [ZPL^{+}17, ACCD11, BCL11, CW14, KTV17, Pen19, Roh19]. networks [ACBL13, ACV14, BB15, Eva18, LLV16, LSR^{+}17, ZLZ12]. Neumann [Kol11]. nodes [CL15a]. noise [BF18, BGN18, CN13, El 10b, EW16, GNZ10, HJ10, JPV10, LLYY19, Mei11, NW11, Ray17, SH14]. Noisy [ANW12b, BHMR14, BJNP13, Bul14, CLM16, CLR17, CTT17, KLT11, LW12b, LMP18, Rei11, Wan13, XZ18]. non [CC19, LvdL16]. non-Euclidean [CC19]. non-unique [LvdL16]. Nonasymptotic [AL19, CCGB17, KP15, Lev15, Zhu17, ABR10a, ABR10b, FLRB13]. noncausal [VL18]. Nonconcave [XZC12]. nonconvex [LYL16, LW17, MBM18, WKL13, WLZ14]. nonconvexity [LW12b]. nonexact [LM12c]. nonhighly [JFF10]. noninferiority [wCM19]. noninvertible [VL18]. Nonlinear [CM12, LW12a, LS17, DS13, FKLZ15, Hua10a, Jac10, LLC13, LAL11, LTG16, MCLX15, MS10b, PP14, WP12, XMW10, YS12, ZW15]. nonnegative [JT15, KRS15]. nonnormal [TGC^{+}19]. nonnull [CJ10]. Nonparametric [ASFJ10, AGR18, BJV17, BCZ10, BV13, CN13, wCM19, CGTW16, CHLP14, DH12b, DHZ16, DB16, Efr13, FFN10, GPPVW14, GJ15, Han19, JY11, JPSZ10, Lam16, MRS12, NS17, PTWZ18, PPB16, QT19, SS11b, SW10b, SC15, TWT17, Vog12, ZW15, Zho10, ZYFW14, APT17, BFT13, BK19, BCW14, Bul13, CN14, Cas15, CTW16, CHM11, CS15, CF16, Gao13, GXZ18, Gho15, HH13, HHW10, JMR14, Kat13, LH10b, LZL19, LJ11, LW10, LY11, LM15, MZ15, MNP19, Nic15, QP16, Rou10, RCL12, Rou15, SY12, Sch13, SvdVvZ15a, SvdVvZ15b, VS19, YT15, YPW17, YG16, dJvZ10]. nonparametrically [MRS12]. nonparanormal [XZ12]. Nonpenalized [WH19]. nonregular [BG14, LMY19]. nonsense [ESW17]. nonsmooth [CL11, Cho17]. nonstandard [BDS19, NBL13]. nonstationary [DW19a, FZ13, GRSP15, Jac10, WP12, WZ18, ZPG18, Zho10, Zho14b]. norm [CTP19, Cas14, KLT11, Lep13, LMP18, YG16]. Normal [KL17, CW12, CH12, JY13, Zho14a]. normality [BM17, BCCZ13, BF18, HPWW11, Loh17, RSZZ15, YFG13]. normalization [OR10, OB13, SB15]. normalized [CSWX16, TP18, Tod15]. normalizing [ULR18]. Note [Khm13, BGvZ21, DM11, FZ21]. NP [BFJ11, FS10]. NP-dimensionality [BFJ11, FS10]. NPV [KF11]. Nuclear [KLT11]. Nuclear-norm [KLT11]. null [CJ10]. number [BCF18a, Bon11, CTT17, LY12, LWY17, TZ19a, Wan11, WXQL14, WZY18]. numerical [XMW10]. objective [AZ12, MGC10]. objects [LKA10, Mül16, PM19, SW18]. observation [JN18]. observational [NMR17]. observations [BHMR14, BF18, DPZ13, DKZ17, OPP17, Rei11, SO18, SN19, XZ18]. observed [ASJ12, Mas13, NS17, ZJS11]. obtain [JG12, JG13, KP11]. occupation [LTT13]. occur [SBC17]. off [SS13]. offs [WBS16]. Old [FG18b]. one [BB10, CZ15a, JMR14, LK10, Wan10, ZPMX11]. one-eighth [ZPMX11]. one-sided [JMR14, LK10, Wan10]. one-sixteenth [ZPMX11]. Online [CCGB17, JM18b, GRSP15]. only [ZJS11]. open [Jia13]. Operational [LMNP17]. operator [VS19]. operators [APT17]. opinions [PANS14]. Optimal [ANW12b, BD14, BGvZ12, BGvZ21, Bel18a, BR13, BD13, BSW11, CJ10, CZZ10, CY11, CZ12, CMW13, CLZ16, CKW^{+}16, CLM16, CW15, Cha17, CK15, CCTV18, DPZ13, DS16, DPZ16, DGJ18, FJY13, FSD^{+}17, Gun12, HP10, HPV10, JM18a, LR14, Ler11, LZA15, LPK17, QWW16, Sam12, TWZ13, Wal10, WLZ14, WXX18, XKB16, Zhe13a, AZ12, Bul16, CL11, CZ18, CGM^{+}13, CCG16, CZvdL17, DRZ12, DGH^{+}19, DM11, DMS12, DMS13a, DG14a, DMG15, DKZ17, Dic11, Fel12, GLZ15, GZ15, HK18, HYS15, JJ12, KLT11, LZ17, LMY19, LPvdGT11, LvdL16, MS10b, NK18, RT11a, SH15, SW10a, SFSL18, SLL11, TL13, WY19, YS12, YT15, YPW17, Yu10, YZ16, Zha13, Zhe13b, Zhe15, ZAL17, ZvdAW19]. optimalities [RSZZ15]. Optimality [PWBM18, BLT18, BCFG11, CMZ19, FXZ14, GL11, HTX17, HYS15, JN18, MMB11, PT10, SF19, XH12]. Optimization [LLV16, BM14, BKM16, CS12, CPW12a, CPW12b, FKLX15, FGM17, GT12, HPŠ10a, LM12b, MH16, RZ12, Wai12, Yua12, ZLZ18]. Optimum [PP14, STY14]. Optional [WM10]. Oracally [WLCY14]. Oracle [CL13b, GL11, HSY^{+}13, KTV17, Lep13, LW19, LPvdGT11, LLLH18, ACL19, BLT18, Bel18b, DS12, FXZ14, Gol10, LM12c, Lep15]. Order [KA10, Bac11, Bel10, BPE14, CGY18, CL13b, DP13, DS11, DG14a, Dic11, EG12, FZ16, FZ21, Jir12, Li14, LC10, Loh15, SH15, WMZ18, YL14]. ordered [DP13, NC10, SW19]. ordering [GCW19, TWT17]. orders [BW11]. ordinary [QZ10, XMW10]. Ornstein [HA13]. orthogonal [GX14, HQ14, HT14, HCT18, LPK17]. Osband [FZ21, FZ16]. other [CGS15, El 10a, JR13]. outcome [LvdL16]. outcomes [JMW15, Lok17]. outlier [CL15a]. outliers [Des15, SC12]. output [HPŠ10a]. Overcoming [WMZ18]. overcomplete [Pen16].

PAC [KP19]. pairwise [HSRW19]. panel [BL14, KLZ16, Kon18]. Parameter [PL11, Des15, Fel12, HW16b, HN16, MS10b, MMB17, MMB18, TW13]. parameters [BCCW18, DN11, FSD^{+}17, Men17, QZ10, XKB16]. Parametric [LY11, Spo12, HSRW19, LLY18, Zhu17]. parametricness [LY11]. Pareto [SW18]. parsimonious [BSW12]. Partial [CF19, HJY16, SR14, BW11, DH12a, GH19, LYW^{+}11, WXZC10, WLLC11, WXQL14, ZL11a]. partial-linear [LYW^{+}11, WXZC10]. partial-sum [GH19]. Partially [SW16a, WZ12, CLLW15, CS15, DGT16, JMW15, KLZ11, LZL19, LLLT10, Ma12, REB18, ZCL16]. particle [CL13a, DBCD19, WL14]. partition [GD17]. partitioning [BCDD14, GD17]. passing [MMB17, MMB18]. Pastur [LAP15]. path [DFKP19, SBC17, TT11]. Pathwise [ZLZ18]. Pauli [CKW^{+}16]. PCA [ACV16, CZ16, JW16b, Nad16, ST16b, AW12, BJNP13, CMW13, Dob17, GZ15, JW16a, KNV15, LV15, PWBM18]. peaks [CS17, SGA11]. penalization [Kol11, KLT11]. Penalized [CZW11, DML10, BC11a, BPE14, CTW18, FXZ14, FG18a, FZ19, LM12a, LYL16, SQ17, TZ19b, WKL13, XZC12, vdGB13]. penalty [Zha10]. Performance [QM11]. periodicity [HKN18]. periodogram [LS10]. Permutation [HBZO19, ABFRB15, CR13, KR11b, MMB11, SR15]. permutations [Muk16]. persistence [FLR^{+}14]. perspective [Cha14, FGM17, PV17, ZBFL18]. Perturbation [DGL19, ZIC12, CZ18, JP13]. Peter [Che16, CF16, Del16, Mül16, Sam16]. Phase [CY11, Che19b, JKW17, CLM16, JJ12, PZ16, SW13, WMZ18]. phenomenon [BDS19, CN14, DM11, TBK12, Yan10]. phylogenetic [DHSM18, Nye11]. Pickands [BDV11, EBGG18]. piecewise [FG18a, Tib14]. piecewise-constant [FG18a]. Pivotal [BCW14]. planar [CGW18, Huc11]. plane [HKKM10]. planted [GD17]. plot [ZDMD18]. plus [El 10b, HCT18, JPV10]. Poincaré [CCF14]. point [ABFRB15, Arm15, BJV17, CZ15b, Che19a, CC19, DL17, EH19b, Fry14, Fry18, HTX17, Jir15, MS10a, PZ16, SS11a, SBK16, WZY18, WXX18, XS13, ZYL18, ZvdAW19, ZYFW14]. point-optimal [ZvdAW19]. points [KPS16, KRS15, MJ18]. Poisson [FLRB13, Jam17]. policies [Kau18]. Pólya [WM10]. polynomial [DGH^{+}19, DMS12, Tib14]. polynomials [CL11, WY19]. polytopes [WRM19]. population [BWY17, BPZ15, BS14]. posed [ISS12, Pen16]. positions [TSP13]. positivity [FLS^{+}17, LUZ19]. possible [Pen13]. possibly [CvdV12, HIT19, JP15, LvdL16, VL18]. post [BLP19, BCCW18, BBB^{+}13, LSST16]. post-model-selection [BLP19]. post-regularization [BCCW18]. post-selection [BBB^{+}13, LSST16]. Posterior [CKG19, CvdV12, LJ11, PBPD14, Atc17, GZ15, GN11, HRSH15, LLL19, NS17, Rou10, YG16]. posteriors [BPY19, RS17]. Power [PHW11, GRD13, OMH13]. Power-enhanced [PHW11]. powerful [Joh13]. PPV [KF11]. precision [CLZ16, CXW13, GRD13, HJY16, Lam16, LC10, SN19]. Prediction [SL19, BFFP19, CHI13, CF19, DW18, LZW10, MM18]. predictive [MJ15]. predictors [BLP19, CFR12, GRSP15, PM19, TZ19a]. presence [CL15a, JT14]. Principal [CGY18, DM18, LAL11, Nye11, BCF18a, BBTW11, BR13, CTT17, DS13, FLW16, FWWZ19, HPV10, JW10, LZW10, LH10b, Ma13, VL13, WBS16]. principle [FZ16, GS10, KMS12, ZBY15, FZ21]. prior [Che19b, DGP18, MG11, MW16, RBWJ19, Roč18]. priors [CSHvdV15, GZ16a, Jam17, KvdVvZ11, NH14, RH12, Yen11, dJvZ10]. probability [AGZ17, DH10, LGS11]. probably [KP19]. probably-approximately-correct [KP19]. probit [QH19]. problem [El 10a, FLRB13, Jia13, Mei11, PR13, SB10]. problems [ABBDL10, Ano10, BDS19, CLR16, CGS15, Fuh10, HL18, ISS12, Jen10, JKW17, KM10a, KvdVvZ11, Pen13, Pen16, PRCS16, PWM18, TT12a, WLZ14, ZYFW14]. procedure [BF12, DML10, LZ17, MR10, MMB11, PT10, TBM11, TBK12]. procedures [ACL19, BGvZ12, BGvZ21, BCFG11, CN14, DR15]. process [ASJ12, Bel10, BFR19, BMBM15, GWB18, JP13, JP15, RCL12, Jam17]. processes [ABFRB15, Bel11, BPD14, BSV14, CLOP19, CGT16, CL08, CL10, CC11, CMW14, CCK14b, CGC11, CAGPFBGM19, DL17, DW19a, DP18, DR10, FLRB13, GRSP15, GH19, HS17, JPV10, JKL12, KLJ15, LLBM^{+}11, Li13, LX12, PO10, PZ16, Por12, RBS10, RD19, TWZ13, VD15, VCC19, ZJS11, ZL11b]. product [LMY18]. profile [MH16]. profiling [QZ10]. profound [YS12]. programming [SH15]. programs [El 10a]. Projected [FLW16, LZL19, CAGPFBGM19]. projection [BBTW11, SWX19]. projection-pursuit [BBTW11]. projections [CZ15a, Lee13]. projective [Akr16]. projectors [KL17]. proof [BCZ^{+}15]. propensity [Zha19b]. Proper [DLP12, PDL12, EG12]. properties [Atc17, BG13, CDH13, DM12, HP10, HH12, KF11, LS18, LLBM^{+}11, MR10, QZ10, TA17, ZHCC11]. Property [NL19, BF18, CCF14]. proportion [CJ10, DR15, GHS14, RV11]. proportional [BFJ11, LZA15, Zhe13a]. proportions [LK10, Wan10]. Provable [LW12b]. Pseudo [ACBL13, BB18b, HR11, OPP17, STRR15]. Pseudo-likelihood [ACBL13]. pseudo-marginal [STRR15]. pseudo-maximum [HR11]. pseudo-observations [OPP17]. pure [BGN18, JKL12, KLJ15, TT12b, Tod15, Tod17]. pure-jump [KLJ15, TT12b, Tod15]. pursuit [BBTW11, KLFL17]. PWM [FdH15].

Q [GK12]. Q-learning [GK12]. quadratic [BHMR14, CCT17, DE17, El 10a, Loh15, XKB16]. quadratically [KJ16]. quadrature [BM17]. QUADRO [FKLX15]. qualitative [CCC19, SHMD13, SSL19]. Quantifying [JBGWS13]. Quantile [CL13c, CLZ19b, EL10, HWH13a, HWH13b, Hua10b, BC11a, BM14, CCG16, GLN18, Kat12, KX14, LMP10, LMP12a, LLLH18, MH16, Por12, SW16a, VCC19, YH12, ZYS18, Zha18, ZPH15, ZPH18, Zho10, ZZX18]. Quantile-adaptive [HWH13a, HWH13b]. quantile-function [CL13c]. quantiles [BW11, CGHH17, HPŠ10a, SW18]. quantization [Lev15]. Quantum [YFG13, BGN18, CKW^{+}16, Li14, Li16a, LMP18, NS11, Wan13]. Quarticity [JR13]. Quasi [Kat13, KM10b, Atc17, CDO11, Mas13, ZL11c]. Quasi-Bayesian [Kat13]. Quasi-concave [KM10b]. quasi-likelihood [Mas13]. quasi-maximum [ZL11c]. quasi-Monte [CDO11]. quasi-posterior [Atc17]. quaternary [MT13, Pho12, ZPMX11]. quotient [FKLX15].

Rademacher [Che19b]. radial [Bel17]. radii [WWG19]. random [AST12, And10, ACV14, APT17, ATW19, BKRS19, CJ11, CCC19, CD13, Che19b, CS17, CCK13, DH10, El 10b, FJ19, GL18, HL18, JP13, JG12, JG13, KLN12, El 10c, KGC10, LX12, Loh15, LM19, Mas13, MH14, Min18, Mül16, PWBM18, PM19, SBV15, SR13, STRR15, TP18, YLZ16]. random-walk [JG12, JG13]. Randomization [BB16a, ZDMD18, HH12, MR10, PR12]. Randomization-based [BB16a, ZDMD18]. randomizations [BB16a, BB10]. Randomized [YPW17, AGL14, Lop19, TT18, ZH10]. randomly [CAGPFBGM19]. range [KPP11, LLBM^{+}11]. Rank [HF19, BLPZ15, BHPZ19, BW19b, BSW11, BSW12, CZ15a, CTT17, EvdG18, FH14a, HPV10, JP13, Kol11, KLT11, KR13, KRS15, LLV16, LD18, LPZZ12, NW11, RT11b, SvdAW14, XZ12, Zha19a, ZvdAW19]. rank-based [HPV10, SvdAW14, XZ12]. rank-one [CZ15a]. ranking [DMJ13, HSRW19]. rankings [HM10]. ranks [CC17, CGHH17, IP11]. Rao [DR11]. Rare [JLX17, QCN18]. Rare-event [JLX17]. Rate [CZ18, GLZ15, GZ15, GZ16a, BC15, Dic11, DHSM18, JM18b, LZ17, Liu13, NR12, SS13, WR13]. Rate-optimal [CZ18, GLZ15, GZ15]. Rates [CL13b, GN11, Han11, Rou10, AL19, ANW12b, BW19a, Bel17, CJ10, CZZ10, CZ12, CMW13, CLZ16, CLM16, CG17, Cas14, CLN19, Gun12, HW19, HK18, HN16, HRSH15, KS16, KLT11, LLL19, LH10b, NS17, PHW11, PS10, RT11a, RW10b, WLZ14, XQB13, YZ16, ZZ16, FLRB16]. ratio [CL13c, JY13, KR11b, NZL17, WWG19, YFG13]. ray [MNP19]. Rayleigh [FKLX15]. Realized [TT12b]. recombination [CMW14]. recommendation [BCR10]. recommender [BQS18]. recovering [LZ17]. recovery [ANW12a, BRS19, CZ15a, CHL17, JKNP12, JN18, LM18, LZ11, LW17, OWJ11, PR16, RT10]. rectangular [OR10, OB13]. recurrent [LTG16]. reduced [BSW11]. reduction [CZC10, FKLX15, KX14, LLC13, LH10a, LAL11, LS17, LLY14, MZ13, MZZ^{+}19, YL11]. reductions [CFR12]. refinements [CL13b]. Refining [ELL15]. refitted [CGZ18]. regimes [SFSL18]. region [SW10a]. regions [ABR10a, BCCW18, CEdH11, NL17, PR16]. Regression [ABT11, Arm15, AC11, BK19, BGvZ12, BGvZ21, Bel18b, BC11a, BCW14, BF12, BM14, BW19b, Bon11, BD13, BCZ10, BPE14, Bul13, CLX13, CG17, CG18, CCG16, CSHvdV15, CC18a, CHI14, CGS15, CHM11, CGTW16, CLZ19b, CS15, CJ12, CF19, CGZ18, DZ17, DGH^{+}19, DH12b, DHZ16, DS13, DPZ16, DPZ19, DW18, DPZ12, DSS11, Efr13, FJR15, FZ19, FGM17, GLN18, GH18, GZ16b, HPŠ10a, HW19, HWCS19, Hua10b, HMLZ11, Jac10, JYW14, JL14, JW15, JMR14, Kat12, KS11, KPS16, KLFL17, LMP12b, LS14, LH10b, LTG16, LZL18, LW12b, LLLH18, MH16, MRS12, MS10a, Mei11, MPL15, NvdG13, NZL17, OWJ11, PM19, Por12, QH19, RYC19, RWL10, SH14, SS11a, SS11b, SW16a, TZ19b, TBM11, TLT16, TGC^{+}19, Vog12, VCC19, WKL13, WCT11, XCL16, YH12, YT15, YD16, YPW17, YG16, YDS16, YC10]. regression [ZW12, Zha13, ZW15, ZPH15, ZPH18, Zhu17]. Regressions [Sch13, CFR12, EPFV18, Nor10, TZ19a]. regressors [Bon11]. regret [MP17]. regular [CEdH11]. Regularization [BFJ11, LM18, MN10, BCCW18, DHSM18, JY16, LW17, RYC19, WMZ18]. Regularized [BM15, XZ12, ACL19, Lam16, RWL10, YP15]. regularly [Des15]. reinforced [AZ12]. rejection [GS10]. Rejoinder [CPW12b, HPŠ10b, JW16b, LTTT14b, SvdVvZ15b, LTTT14a]. related [Gol10, JKW17]. Relative [FR13, DML10, HL13]. relatives [FGM17]. relaxation [ANW12b]. relaxations [AL18, KNV15]. relaxed [AcL17]. relevant [DW19a]. remedy [BK19]. Remembering [vZ11]. Rényi [HW16a, LPS08, LP10]. repeated [ZLST17]. repeatedly [JMW15]. replicates [KP11]. Reply [Fuh10]. representations [CM12, DS13]. reproducing [AW12, LLLH18, YC10]. Rerandomization [MR12]. Resampling [KP11, ABR10a, ABR10b, BT17, GCW19]. Residual [CL08, CL10]. residuals [BSV14, GH19]. resolution [CLN19, GX14]. resource [Kau18]. respect [LLY14]. response [AGR18, BW19b, DG14a, FSD^{+}17, KLFL17, TT18, ZH10]. response-adaptive [AGR18, ZH10]. responses [Xu17, ZLK12]. Restricted [EKDN18, Bel18b, CGS15, LJ11, MW16, Xu17]. results [ABR10a, ABR10b, DE17, GHS14, LX12]. retrieval [CLM16]. Reversible [HJY13, Bac11, BFT13]. revisited [BB18a, FG18b]. reward [CZvdL17]. Rho [BB18a]. Rho-estimators [BB18a]. Rice [TLT16]. Ridge [DW18, GPPVW14]. ridges [CGW15]. Riemannian [DM18, LY19]. Rietz [Ste11]. right [SW12]. right-censored [SW12]. Risk [El 10a, BJL15, CEdH11, CGS15, CL15b, DG14b, DML10, DKR18, KM10a, LN11, MBM18]. risks [KJ16]. RKHS [SSGF13]. RKHS-based [SSGF13]. RMT [BCF18b]. Robbins [BF12]. Roberts [PT10]. Robust [AC11, BBTW11, BK18, CL15a, CGR18, DMS13a, EvdG18, GWB18, HW17, LZ11, LPZZ12, SEC14, CR13, EW16, EBGG18, FFB14, FH14a, Loh17, MZZ^{+}19, ZBFL18]. Robustness [Des15, TS12, ZYL18]. ROC [KF11]. ROCKET [BK18]. Role [SB15]. Romano [DR15]. root [BCW14, Dic11, Por12, ZPG18]. root- [Por12]. roots [DS11, ZvdAW19]. ROP [CZ15a]. Rotation [AST12]. rotationally [CPV17]. rough [PL11]. rule [BS14, CZvdL17, GKM16]. rules [DLP12, EG12, JM18b, PDL12, QM11].

Saddlepoint [KR11b]. sample [ACF12, BWY17, BPZ15, BB16b, CW15, CLSY18, CQ10, CLZ19a, DLP12, FG18b, FLRB13, Gao13, JM18a, JPSZ10, KL17, LMNP15, LMNP17, LC12, LWY17, PTWZ18, QZ10, RWG19, Sae19, Spo12, TA17, XQB13, ZBY15, ZBFL18]. sample-based [BWY17]. Sampled [AW12, BG13, CY11]. sampler [DBCD19, RH12]. samples [CL18, DMS13b, KV17, Lv13]. Sampling [VR19, BFR19, BLRG17, CLSY18, Cha15b, EL10, FH14b, KF11, MH13, PS10, Roh19, SW13, SR13]. sandwich [KH11]. Saturated [HYS15]. scalable [FL16]. scalar [NS17]. scale [AST12, And10, CJ10, CKW^{+}16, CW19, Des15, Efr13, HQ11, MB15, ZFY11, dJvZ10]. scaled [ZIC12]. scaling [NW11]. scan [Wal10]. scanning [PWM18]. scatter [CGR18, PV18]. scattering [BMBM15]. scheme [JN18]. scope [CTW18]. score [NHM18, Zha19b, Jin15]. score-based [NHM18]. scores [LZW10]. scoring [DLP12, EG12, KMH14, PDL12]. scrambled [BM17, Dic11]. Screening [RT11a, CTW13, CTW16, CHLP14, FS10, FKLZ15, Han19, HWH13a, HWH13b, HJY16, KJF14, LPZZ12, MZ15, SEG19, Ste11, XCL16]. SDE [Mas13]. SDEs [GLM18]. search [BPE14]. Second [Li14, CGY18, DG14a, SH15, YL14]. Second-order [Li14, CGY18, DG14a, SH15, YL14]. section [BC11b]. segmentation [AL11, Fry14]. Selecting [CTT17]. selection [BLP19, BBFGD12, BBB^{+}13, BKM16, BSW11, BSW12, BNST18, CS12, CKG19, CPW12a, CPW12b, CHI14, CC18b, CZC10, CL15b, CZW11, CMKR12, CD12, DML10, FL12, FFB14, GT12, GL11, HIT19, HHW10, JJ12, JL14, Jir12, KLZ11, KJ16, KS11, LM12b, LSST16, LLL19, Ler11, LKZ15, LY11, NH14, RWL10, RZ12, SW10a, SB10, SQ17, SW19, TRTW18, VAC17, Wai12, WLLC11, WXQL14, WB17, WH19, YWJ16, Yen11, YDS16, Yua12, Zha10, ZW15]. Selective [TT18, BC19]. Self [CSWX16, Tod15, ZL11c]. Self-normalized [CSWX16, Tod15]. self-weighted [ZL11c]. Semi [ZBC19, CC18a, CS15, LKZ15]. semi-nonparametric [CS15]. Semi-supervised [ZBC19, CC18a]. semi-varying [LKZ15]. semidefinite [AL18, KNV15]. Semimartingale [Bul17]. semimartingales [Bul16, TT12b, Tod15, Tod17]. Semiparametric [CLLW15, JvdG18, SvdAW14, SW16b, TS12, BF12, BK12, CR15, CTW16, CH10, DN11, DML10, EPFV18, KLZ11, LHY^{+}12, NZL17, SYZL14, Zhu17]. Semiparametrically [IP11, ZvdAW19]. semisupervised [ASW13]. semivarying [CHL16, LXQ16]. sensing [Bul13, CJ11]. sensitive [ASW13]. sensitivity [BD11, DP18, TS12]. sensor [XS13]. separability [APT17]. separate [LLY18]. separately [CW14]. Separation [FLRB16, ATHW12, And10, BHM18, STD10]. sequence [YLZ16]. sequences [CvdV12]. Sequential [Che19a, PR12, SF19, XS13, ZH10, CGM^{+}13, CZvdL17, Cha17, GS10, Kau18, KF11, NJ13]. Sequentially [BDD10]. serial [FR13]. series [BM15, BYZ10, CL08, CL10, CI11, CHI13, CGY18, CXW13, GLN18, HKN18, HIT19, HR11, JP15, KP11, LY12, LXQ16, LS10, LW10, LAP15, NBL13, PT13, Pap18, Vog12, WZ18, XW12, ZS13, ZW17, ZPG18, Zho10, Zho14b]. set [CQ10, Gun12, MW16, MT13]. sets [Bel17, CGW18, Cas15, FLR^{+}14, FH14b, Gho15, LM15, NvdG13, Nic15, Rou15, SZ15, SvdVvZ15a, SvdVvZ15b, YG16]. settings [CC18a, LZW10, NMR17, WH19]. setup [WW18]. several [BD13]. shape [Bel18b, BG10, CLX13, CGS15, PV18, SHMD13, Vim10]. shapes [Huc11]. shared [SW16b]. shared-frailty [SW16b]. Sharp [AGL14, BW19a, Bel18b, DS12, Dob17, ACL19]. shifted [BG10]. Shiryaev [PT10]. short [CL08, CL10]. shrinkage [DGJ18, HMLZ11, LW12a, XKB16]. shrinking [NH14]. sided [JMR14, LK10, Wan10]. Sieve [CGT16, Pap18, XMW10, DN11, KPP11]. Sieve-based [CGT16]. Signal [CCC19, OMH14, ISS12, SQ17]. signals [CW15, CHL17, FG18a, HJ10, JKNP12, PV17, QCN18, Roč18]. signed [IP11]. significance [BMvdG14, BB14, CY14, FK14, LTTT14a, LTTT14b, LTTT14c, LZ14, Was14]. significant [BCF18a]. signs [BS14, CC17, CGHH17]. similarity [Liu17]. simple [HH13, SL19, YS12]. Simultaneous [FLSZ18, Jir12, LW10, LS10, WLCY14]. simultaneously [HL18]. Single [CHM11, DZ14, BLRG17, CLLW15, CHZ11, DGT16, JW11, JMW15, LS14, LYW^{+}11, LLLT10, MH16, WXZC10]. Single-index [DZ14, CLLW15, CHZ11, DGT16, LS14, LYW^{+}11, LLLT10, MH16, WXZC10]. single-stage [BLRG17]. Singular [Cha15a, CZ18, CTP19, Che19c, DG14b, GPPVW12, LWY17]. singularities [PV17, Pen13]. sixteenth [ZPMX11]. size [BGvZ12, BGvZ21, God18, JM18a]. sizes [DB16, Ma12]. sketches [YPW17]. slab [Roč18, Yen11]. sliced [JL14, LZL18, YDS16]. SLOPE [SC16, BLT18]. small [LM18]. small-ball [LM18]. smallest [LK10, Wan10]. smeary [EH19a]. Smirnov [FG18b]. Smirnov-type [FG18b]. Smooth [HP18, CR15, Dic11, GLN18, KR13, LMP10, LMP12a, Nor10]. smoothed [GJW10, Gro14]. smoothing [CDH13, JMW15, SC13, ZS13]. smoothness [Loh15, SS13]. Solution [Pen16, TT11]. solutions [LYL16]. solve [HL18, KNV15]. solved [ESW17]. Some [ABR10a, ABR10b, LX12, LK10, Ano10, Atc17, BCFG11, CL13a, Fuh10, HS17, HN16, Jen10, MMNP15]. Sorted [FZ19]. source [BHM18]. sources [Sae19]. Space [Ano10, Fuh10, Jen10, AL19, AST12, Fuh06, HQ14, LH10a, LLLH18, Nye11, PT13, PM16a, SLQ14, TW13, YC10]. space-filling [HQ14, SLQ14]. spaced [BLN15, Loh15]. spaces [AW12, CCGB17, CDO11, DLP12, DS13, GKM16, Lev15, LMY18]. SPADES [BTWB10]. Sparse [CMW13, GMZ17, KP15, LM18, Ma13, RT10, SWDW11, ACBL13, ACCP11, BM15, BFL^{+}18, BC11a, BR13, BJNP13, CZ12, CLZ16, CEL16, CLM16, CvdV12, CSHvdV15, CW15, CZC10, CHL17, EH19b, FJY13, FRW15, FLSZ18, GMRZ15, GZ15, HJ10, Han19, HJY13, HMLZ11, JBD17, JKNP12, KTV17, KNV15, LLL19, LV15, LZ17, LSR^{+}17, Liu17, LLLH18, MZZ^{+}19, MJ15, MPL15, MMS18, MMY18, NMR17, NvdG13, NL17, PBPD14, RT11a, Roč18, TWZ13, VR19, VAC17, VL13, WLZ14, WBS16, XZC12, Zha13, ZW16, ZLZ18, vdGB13]. sparse-change [LSR^{+}17]. Sparsistency [LV15]. Sparsity [KY10, BCFG11, CV19a, CCT17, CCTV18, HZ10, LZL18, LPvdGT11, NR12, Pen17, SC16, SS13, ZCX13]. spatial [BLN15, CC17, Rao18, SYZL14, Wal10]. Spatially [Bul13, Pen13]. Spatially-adaptive [Bul13]. Spearman [BLPZ15]. special [BC11b]. specification [HL13, LKZ15, SBK16, WP12]. Spectral [BLPZ15, Bel10, RCY11, AS11, DMS13b, FJ10, GD17, GNZ10, JPSZ10, JY16, KH11, KL17, LR15, RCL12, SB15, SWY15, XQB13, XZ18, ZBY15]. Spectrum [KV17, El 10c]. spheres [CPV17, DM18, EH19a]. spherical [HLLZ19, KNP11]. sphericity [OMH13]. spike [Roč18, Yen11]. spike-and-slab [Roč18]. Spiked [PWBM18, Che19b, DGJ18, WF17, WY17]. spline [JMW15, LZL19, Ma12, SC13, WZ12]. split [ZDMD18]. split-plot [ZDMD18]. splitting [RWG19]. spurious [FSZ18]. square [BCW14, Dic11]. square-root [BCW14]. squared [CHI13]. Squares [Bel18b, AC11, Cha14, CF19, DH12a, GZ16b, HW19, Jac10, SS11b]. stage [BGvZ12, BGvZ21, BLRG17, TBM11]. State [Ano10, Fuh10, Jen10, CKW^{+}16, CDO11, Fuh06, Wan13]. statements [SHMD13]. states [BGN18, Li16a]. stationarity [HBT16, Tru19]. stationary [CGY18, Ler11, PO10, PT13, RD19, Ver10, Vog12, VD15, XW12]. Statist [BGvZ21, Fuh10, Jen10]. statistic [LS13, SR15]. Statistical [BL12, BWY17, Bel11, BCR10, FH14a, Loh17, LvdL16, Rao18, WBS16, Zhu19, ANW12a, CLR17, Chi11, FLSZ18, LKA10, LLY14, WLZ14]. Statistics [CV19b, BLPZ15, CZ18, CTP19, Che18, DMS13b, ELL15, KR11b, Lah10, Rao18, SSGF13, Tod15, Wal10, ZBY15, Zho14b]. status [GJW10, GH18, TBK12, WZ12]. stem [HE18]. step [DB16, JPSZ10, Ma12]. step-sizes [DB16]. Stiefel [CV19b]. Stochastic [Bow10, WCT11, BCCZ13, CLX18, DP13, DB16, GWB18, Jac10, KRD10, LR15, Lei16, Lia10, RCY11, SB15, SS11a, TWT17, TA17, WB17, ZZ16, ZLZ12]. Stolarsky [HBZO19]. strategies [PS10]. strategy [LvdL16]. stratum [CC18b]. Straw [CvdV12]. stream [Cha17]. strength [HT14, HCT18]. strictly [Han19]. Strong [FXZ14, HCT18, HK18, EKDN18, HT14]. Structural [BD15, Liu17, DFS11, DFW19, FDD12, REB18, RBM19, WZ18]. Structure [KLZ16, LW13, ATHW12, BJR16, CJ11, CHLP14, CS15, Lep13, LW19, LKZ15, NHM18]. structured [RLM^{+}17]. structures [FLS^{+}17, ZCCL19, ZLST17]. studentized [Lah10]. study [ACF12, JLP^{+}16, KP15]. Sub [DLLO16, LM19, Min18, GRSP15, PWBM18]. Sub-Gaussian [DLLO16, LM19, Min18]. sub-linear [GRSP15]. sub-optimality [PWBM18]. subclass [Vim10]. subexponential [LM12c]. subgradient [FGM17]. subject [XH12, Zhe13b]. submatrix [CLR17, GL18, MW15]. submodularity [EKDN18]. Subsampling [BB15, FH14b, RS12, TPN19]. subset [BKM16, Jia13]. subspace [CTP19, Pen18, SC12, SEC14, VL13]. subspaces [CZ18, LZ11]. Substitution [ZBY15]. Successive [OR10, OB13, ASJ12]. Sufficient [YL11, CZC10, CFR12, LLC13, LAL11, LS17, MZ13, MZZ^{+}19, VR12]. sum [GH19]. sums [Bob19, CCK13, CHL17, LD18]. sup [Lep13, LMP18]. sup-norm [Lep13, LMP18]. Super [CLN19, BDS19, BFR19]. super-efficiency [BDS19]. super-efficient [BFR19]. Super-resolution [CLN19]. supersaturated [SLL11]. supervised [CC18a, FKLX15, ZBC19]. Support [LSR^{+}17, LW17, MJ18, OWJ11, Gun12, LAL11, PR16]. supported [KJ16]. suprema [CCK14b]. Supremum [YG16, Cas14]. Sure [FS10, CTW13]. surface [DG14a, JY11]. surfaces [APT17]. surrogate [DKR18]. survival [wCM19, SW12]. symmetric [CPV17, IP11]. symmetries [GL12]. systematic [Kon18]. systems [BQS18, HS17, MMNP15].

Table [Ano16d, Ano16e, Ano16f, Ano17b]. tables [BB10]. Tail [Fry18, CEKL15, EKS12]. Tail-greedy [Fry18]. tailed [BJL15, HW19, Min18]. tailored [Zha19b]. tangent [AL19]. Targeted [CZvdL17]. tau [Bao19, BK18]. techniques [WMZ18]. temporally [SN19]. Tensor [JBD17, BQS18, Che19b, RYC19, WZ12, Zha19a]. terms [FG18b]. Test [CGZ18, QC12, ZCGZ19, BDV11, BMvdG14, BB14, CY14, CQ10, FK14, HLLZ19, JP13, Lei16, LLY18, LTTT14a, LTTT14b, LTTT14c, LZ14, PTWZ18, SR15, WP12, Was14, YP15]. Testing [ASJ11, BDP19, CL11, CL18, CPV17, HKN18, Hua10a, KLJ15, LD18, Tod17, ABBDL10, ACCP11, Arm15, BW19a, BFL^{+}18, BCFG11, CJ10, CJ11, CEL16, CEKL15, wCM19, CQ10, CS17, Chi11, DH12b, DZ14, FLRB16, HPV10, HL13, Khm13, KSZ18, Li14, LLYY19, LLLT10, Liu17, MMB11, MPL15, MMY18, NJ13, NL19, NS11, PV17, PANS14, RBWJ19, SGA11, SSGF13, SSCL19, SSL19, SF19, WZHO17, WWG19, ZFY11, ZPG18, ZYS18, ZBY15, ZCCL19, ZBFL18, ZZX18]. Tests [APT17, CC17, ZPW13, ZCX13, ZLST17, ASFJ10, ABFRB15, ABR10b, Bul17, CLZ19a, CR13, CAGPFBGM19, EPFV18, FG18b, FLRB13, GRD13, HMV13, JY13, Jir15, Joh13, KLN12, KR11b, LC12, LS10, LS13, NL17, OMH13, PR12, TWT17, VV10, WWG19, ZvdAW19]. TFT [KP11]. TFT-bootstrap [KP11]. th [Pho12]. th-fractions [Pho12]. their [Che18, FSZ18, GZ16b, KH11, LMY18, LW13]. theorem [BK12, Bob19, BG14, CR15, DN11, EH19a, HQ14, LS13, MH11, RR12]. theorems [BLRG17, Bon11, CN13, DR10, JPV10, JY13, Ray17, TP18]. theoretic [AL11, VS19]. Theoretical [QP16, ZHCC11]. Theory [BL14, CV19b, Akr16, BB18a, CLOP19, CL13a, CHI14, CGW15, DH12a, FdH15, HA13, LLC13, MH14, MT13, NL17, OPP17, Pen17, PS18, Sae19, SF19, Spo12, TL13, TS12, VR12, ZBC19, ZLZ18, ZBFL18]. Think [WW18]. three [HT14]. threshold [DFKP19, LJ18, Roh19]. thresholded [CLM16]. thresholding [CY12, DG14b, KA10, Ma13, NR12, QCN18, SWDW11, Cha15a]. thresholds [MMS18]. Tight [CD12]. Time [ZW15, ASFJ10, BM15, BYZ10, Bel11, Bul14, CEL16, CL08, CL10, CI11, CHI13, CGY18, CXW13, CGC11, DPZ19, GRSP15, GLN18, HKN18, HIT19, HR11, JP15, KRD10, KP11, LY12, LMNP17, LXQ16, LJ18, LS10, LW10, LAP15, NBL13, PT13, Pap18, Tod17, Tru19, Vog12, WZ18, XW12, XMW10, ZJS11, ZW12, ZS13, ZW17, ZPG18, Zho10, Zho14b]. time-changed [Bel11]. time-changes [Bul14]. time-frequency [CEL16]. time-inhomogeneous [Tru19]. Time-varying [ZW15, Tod17, XMW10, ZW12]. times [LTT13, ZJS11]. tomography [CKW^{+}16, LMP18, Wan13]. tool [KMS12]. Total [FLS^{+}17, LUZ19]. trace [KH11]. trace-class [KH11]. Tracy [Bao19, HPZ16]. Trade [SS13, WBS16]. Trade-off [SS13]. trade-offs [WBS16]. trajectories [PPB16]. Trajectory [Lia10]. transelliptical [BK18]. transfer [KMS12]. transformation [JG12, JG13, MM18, PM16a]. transformations [KRD10]. transformed [SW10b, TGC^{+}19]. transforms [MNP19, TT12b]. transition [CY11, Che19b, WMZ18]. transitions [JKW17]. transport [CCG16]. treatment [CZvdL17, LvdL16, MP17, QM11, RBWJ19, SFSL18, SSL19, Zha18]. tree [BO17, HA13, Huc11, SW18, WM10]. trees [Che19c, Nye11]. Trek [STD10, FDD12]. trend [ST19, Tib14]. trials [AZ12, ZH10]. trigonometric [ZPMX11]. true [CTT17]. truncated [SW12]. tuning [SW19]. Twisted [WL14]. Two [BB10, CLZ19a, FG18b, LC12, Ma12, BGvZ12, BGvZ21, BMW18, CTP19, CC18b, CQ10, DMS12, EG12, FLRB13, God18, HCT18, LK10, PTWZ18, SW13, TBM11, Wan10, Zhe15]. two-level [CC18b]. two-phase [SW13]. Two-sample [CLZ19a, FG18b, CQ10, FLRB13]. two-stage [BGvZ12, BGvZ21, TBM11]. Two-step [Ma12]. two-to-infinity [CTP19]. two-way [BMW18]. type [CSWX16, DKL12, FG18b, HPZ16, HS17, LS10].

Uhlenbeck [HA13]. ultra [CHLP14, FL16, LKZ15, SW16a, WKL13, ZPH15]. ultra-high [CHLP14, LKZ15, SW16a, WKL13, ZPH15]. ultra-large [FL16]. ultrahigh [Han19, YDS16]. unbiased [JT15, Zha10]. uncertainty [BD15, RT10]. underestimation [El 10a]. understanding [CD13]. Unexpected [CDH13, PM16b]. unified [RBWJ19]. Uniform [CI11, Jir15, LH10b, SWX19, TXL12, TRTW18, BSV14, LN11, RS12, TWT17]. uniformity [CPV17]. Uniformly [BCCW18, Joh13]. Unifying [LS18]. union [OWJ11]. unique [LvdL16]. Unit [DS11, ZPG18, ZvdAW19]. universal [Gol10, Cha15a]. Universality [BPZ15, PY12]. Universally [TSP13, Zhe13b, Zhe15, ZAL17]. unknown [AGZ17, CW12, SC16, Zhu19]. unlabeled [SW18]. unseen [WY19]. upper [CGM^{+}13]. UPS [JJ12]. urn [AGR18, ZHCC11]. using [AL11, BFFP19, BPD14, BCDD14, CGZ18, DRZ12, Des15, ELL15, Hua10a, KRD10, LPK17, Pen16, RBWJ19, RWL10, WCT11, Yen11, ZYL18, dJvZ10].

Valid [BLP19, BBB^{+}13, BCCW18, HPWW11]. validation [CGZ18, RD19, XH12]. validity [BT17, KPP11, RS12]. Value [Cha15a, BDV11, DG14b, ELL15, FdH15, HBZO19]. valued [LKA10]. values [CZ10, EJ10, LWY17]. vanilla [DR11]. variability [HM10]. Variable [BNST18, FL12, HHW10, JL14, JG12, JG13, Bac11, BBFGD12, BSW12, CS12, CPW12a, CPW12b, CZC10, CZW11, CD12, DML10, Efr13, FFB14, GT12, HWH13a, HWH13b, JJ12, KLZ11, KS11, LM12b, NH14, RZ12, SB10, Wai12, WLLC11, WH19, XCL16, YWJ16, Yen11, Yua12, Zha10]. variable-order [Bac11]. variable-selection [SB10]. variables [AV13, CMKR12, DZ17, HP18, Kat13, RBM19]. variance [And10, AGL14, Atc11, CGT16, CGZ18, DE17, FFN10, FJ19, FJ10, HL18, Lop19, XKB16, YL11]. variances [MDO14]. variants [KH11]. variate [Zho14a]. variation [BS14, CEdH11, JT14, PZ16]. variational [BCCZ13, HPWW11]. variations [Loh15]. various [Ler11]. varying [Des15, GRSP15, KLZ11, KP15, LMP12b, LKZ15, LX12, SBK16, Tod17, XMW10, ZW12, ZW15, ZLK12]. varying-coefficient [KLZ11]. Vast [WZ10]. Vector [CCG16, CV19a, CGY18, Lev15, LAL11, LM19, ZPL^{+}17]. vectors [BHPZ19, BDP19, BKRS19, CCK13]. verification [HF19]. vertex [TSP13]. very [HP10]. via [ANW12b, BCF18b, BC15, BK18, BCW14, Bel11, BSY19, BM14, BKM16, BGK10, BSV14, CZ15a, CLM16, CS12, CPW12a, CPW12b, CK15, CL15b, DR15, DHSM18, DP18, FKLX15, FGM17, GT12, HW16a, HBZO19, Jam17, JL14, KLFL17, LM12b, LLV16, Lep15, Li13, LLX19, Loh15, Lop19, NC10, QCN18, RZ12, SY12, TLT16, Tib14, Tod15, Wai12, WH19, Yua12, ZFY11]. Volatility [LTT13, BJV17, JR13, JP13, JT14, KRD10, Kon18, LLX19, Rei11, TWZ13, WZ10]. vs [Men17].

walk [JG12, JG13, STRR15]. Walker [Jir12]. wavelet [LN11]. way [BMW18]. Weak [BB18b, SQ17, BSV14, EKDN18, Lah10, PV17]. Weakly [HK10, HN16]. weighing [CK15]. weight [AcL17]. weight-relaxed [AcL17]. Weighted [SW13, Bob19, DGP18, Sam12, Zho14b, ZL11c]. Wendland [BFFP19]. Westfall [MMB11]. whether [ASJ11]. white [BGN18, CN13, GNZ10, JLX17, LLYY19, Mei11, Ray17]. wide [JLP^{+}16]. Widom [Bao19, HPZ16]. Wiener [Koi19]. Wigner [LMP18]. Wild [Fry14, FLRB13]. Wirtinger [CLM16]. wise [PHW11, ZCX13, FLRB16]. Wishart [CW19, DMO08, DG12, HKKR11, JLX17, KR11a, ULR18]. within [CGC11]. without [LW17, Pen17]. Wolf [DR15]. work [Roj11, Sam16].

X [MNP19]. X-ray [MNP19].

Young [MMB11]. Yule [ESW17, Jir12].

Zag [BFR19]. Zig [BFR19].

References

Addario-Berry:2010:CTP

[ABBDL10] Louigi Addario-Berry, Nicolas Broutin, Luc Devroye, and Gábor Lugosi. On combinatorial testing problems. Annals of Statistics, 38(5):3063–3092, October 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

               //projecteuclid.org/euclid.aos/1283175989.

Albert:2015:BPT

[ABFRB15] Mélisande Albert, Yann Bouret, Magalie Fromont, and Patricia Reynaud-Bouret. Bootstrap and permutation tests of independence for point processes. Annals of Statistics, 43 (6):2537–2564, December 2015. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

               projecteuclid.org/euclid.aos/1444222084.

Arlot:2010:SNRa

[ABR10a] Sylvain Arlot, Gilles Blanchard, and Etienne Roquain. Some nonasymptotic results on resampling in high dimension, i: Confidence regions. Annals of Statistics, 38(1):51–82, February 2010. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1262271609.

Arlot:2010:SNRb

[ABR10b] Sylvain Arlot, Gilles Blanchard, and Etienne Roquain. Some nonasymptotic results on resampling in high dimension, II: Multiple tests. Annals of Statistics, 38(1):83–99, February 2010. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1262271610.

Aswani:2011:RME

[ABT11] Anil Aswani, Peter Bickel, and Claire Tomlin. Regression on manifolds: Estimation of the exterior derivative. Annals of Statistics, 39(1):48–81, February 2011. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1291388369.

Audibert:2011:RLL

[AC11] Jean-Yves Audibert and Olivier Catoni. Robust linear least squares regression. Annals of Statistics, 39(5):2766–2794, October 2011. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1324563355.

Arias-Castro:2012:DC

[ACBL12] Ery Arias-Castro, Sébastien Bubeck, and Gábor Lugosi. Detection of correlations. Annals of Statistics, 40(1):412–435, February 2012. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1334581748.

Amini:2013:PLM

[ACBL13] Arash A. Amini, Aiyou Chen, Peter J. Bickel, and Elizaveta Levina. Pseudo-likelihood methods for community detection in large sparse networks. Annals of Statistics, 41 (4):2097–2122, August 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1382547514.

Arias-Castro:2011:DAC

[ACCD11] Ery Arias-Castro, Emmanuel J. Candès, and Arnaud Durand. Detection of an anomalous cluster in a network. Annals of Statistics, 39(1):278–304, February 2011. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1291388376.

Arias-Castro:2011:GTU

[ACCP11] Ery Arias-Castro, Emmanuel J. Candès, and Yaniv Plan. Global testing under sparse alternatives: ANOVA, multiple comparisons and the higher criticism. Annals of Statistics, 39(5):2533–2556, October 2011. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1322663467.

Asgharian:2012:LSS

[ACF12] Masoud Asgharian, Marco Carone, and Vahid Fakoor. Largesample study of the kernel density estimators under multiplicative censoring. Annals of Statistics, 40(1):159–187, February 2012. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1331830778.

Ando:2017:WRM

[AcL17] Tomohiro Ando and Ker chau Li. A weight-relaxed model averaging approach for high-dimensional generalized linear models. Annals of Statistics, 45(6):2654–2679, December 2017. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

Alquier:2019:EBS

[ACL19] Pierre Alquier, Vincent Cottet, and Guillaume Lecué. Estimation bounds and sharp oracle inequalities of regularized procedures with Lipschitz loss functions. Annals of Statistics, 47(4):2117–2144, August 2019. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://

projecteuclid.org/euclid.aos/1558425641.

Arias-Castro:2014:CDD

[ACV14] Ery Arias-Castro and Nicolas Verzelen. Community detection in dense random networks. Annals of Statistics, 42(3):940– 969, June 2014. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1400592648.

Arias-Castro:2016:DIF

[ACV16] Ery Arias-Castro and Nicolas Verzelen. Discussion of “Influential features PCA for high dimensional clustering”. Annals of Statistics, 44(6):2360–2365, December 2016. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/

1479891618. See [JW16a, JW16b].

Aronow:2014:SBV

[AGL14] Peter M. Aronow, Donald P. Green, and Donald K. K. Lee. Sharp bounds on the variance in randomized experiments. Annals of Statistics, 42(3):850–871, June 2014. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1400592645.

Aletti:2018:NCA

[AGR18] Giacomo Aletti, Andrea Ghiglietti, and William F. Rosenberger. Nonparametric covariate-adjusted response-adaptive design based on a functional urn model. Annals of Statistics, 46(6B):3838–3866, December 2018. CODEN ASTSC7. ISSN

0090-5364. URL http://projecteuclid.org/euclid.aos/
1536631292.

Anevski:2017:EPM

[AGZ17] Dragi Anevski, Richard D. Gill, and Stefan Zohren. Estimating a probability mass function with unknown labels. Annals of Statistics, 45(6):2708–2735, December 2017. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

Akritas:2016:ATF

[Akr16] Michael G. Akritas. Asymptotic theory for the first projective direction. Annals of Statistics, 44(5):2161–2189, October 2016. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1473685272.

Aue:2011:ISU

[AL11] Alexander Aue and Thomas C. M. Lee. On image segmentation using information theoretic criteria. Annals of Statistics, 39(6):2912–2935, December 2011. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1327413773.

Amini:2018:SRB

[AL18] Arash A. Amini and Elizaveta Levina. On semidefinite relaxations for the block model. Annals of Statistics, 46(1): 149–179, February 2018. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

Aamari:2019:NRM

[AL19] Eddie Aamari and Clément Levrard. Nonasymptotic rates for manifold, tangent space and curvature estimation. Annals of Statistics, 47(1):177–204, February 2019. CODEN ASTSC7.

ISSN 0090-5364. URL http://projecteuclid.org/euclid.
aos/1543568586.

Anderes:2010:CSS

[And10] Ethan Anderes. On the consistent separation of scale and variance for Gaussian random fields. Annals of Statistics, 38(2):870–893, April 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1266586617.

Anonymous:2010:SPA

[Ano10] Anonymous. On some problems in the article efficient likelihood estimation in state space models. Annals of Statistics, 38(3):1936, June 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1271271283.

Anonymous:2016:EBa

[Ano16a] Anonymous. Editorial board. Annals of Statistics, 44(4):??, August 2016. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1467894701.

Anonymous:2016:EBb

[Ano16b] Anonymous. Editorial board. Annals of Statistics, 44(5):??, October 2016. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1473685256.

Anonymous:2016:EBc

[Ano16c] Anonymous. Editorial Board. Annals of Statistics, 44(6):??, December 2016. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1479891616.

Anonymous:2016:TCa

[Ano16d] Anonymous. Table of contents. Annals of Statistics, 44(4): ??, August 2016. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1467894700.

Anonymous:2016:TCb

[Ano16e] Anonymous. Table of contents. Annals of Statistics, 44(5):??, October 2016. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1473685255.

Anonymous:2016:TCc

[Ano16f] Anonymous. Table of contents. Annals of Statistics, 44(6):??, December 2016. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1479891615.

Anonymous:2017:EBd

[Ano17a] Anonymous. Editorial board. Annals of Statistics, 45(1):??, February 2017. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1487667615.

Anonymous:2017:TCd

[Ano17b] Anonymous. Table of contents. Annals of Statistics, 45(1):??, February 2017. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1487667614.

Agarwal:2012:FGC

[ANW12a] Alekh Agarwal, Sahand Negahban, and Martin J. Wainwright. Fast global convergence of gradient methods for high-dimensional statistical recovery. Annals of Statistics, 40(5):2452–2482, October 2012. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1359987527.

Agarwal:2012:NMD

[ANW12b] Alekh Agarwal, Sahand Negahban, and Martin J. Wainwright. Noisy matrix decomposition via convex relaxation: Optimal rates in high dimensions. Annals of Statistics, 40(2):1171–1197, April 2012. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1342625465.

Aston:2017:TSN

[APT17] John A. D. Aston, Davide Pigoli, and Shahin Tavakoli. Tests for separability in nonparametric covariance operators of random surfaces. Annals of Statistics, 45(4):1431–1461, August 2017. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1498636862.

Armstrong:2015:ATR

[Arm15] Timothy Armstrong. Adaptive testing on a regression function at a point. Annals of Statistics, 43(5):2086–2101, October 2015. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1438606854.

Anevski:2011:MSD

[AS11] Dragi Anevski and Philippe Soulier. Monotone spectral density estimation. Annals of Statistics, 39(1):418–438, February 2011. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1297779852.

Ait-Sahalia:2010:NTM

[ASFJ10] Yacine Aït-Sahalia, Jianqing Fan, and Jiancheng Jiang. Nonparametric tests of the Markov hypothesis in continuoustime models. Annals of Statistics, 38(5):3129–3163, October 2010. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1284391760.

Ait-Sahalia:2010:BMN

[ASJ10] Yacine Aït-Sahalia and Jean Jacod. Is Brownian motion necessary to model high-frequency data? Annals of Statistics, 38(5):3093–3128, October 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1284391759.

Ait-Sahalia:2011:TWJ

[ASJ11] Yacine Aït-Sahalia and Jean Jacod. Testing whether jumps have finite or infinite activity. Annals of Statistics, 39 (3):1689–1719, June 2011. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1311600280.

Ait-Sahalia:2012:ISB

[ASJ12] Yacine Aït-Sahalia and Jean Jacod. Identifying the successive Blumenthal–Getoor indices of a discretely observed process. Annals of Statistics, 40(3):1430–1464, June 2012. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1346850061.

Adler:2012:RSS

[AST12] Robert J. Adler, Eliran Subag, and Jonathan E. Taylor. Rotation and scale space random fields and the Gaussian kinematic formula. Annals of Statistics, 40(6):2910–2942, December 2012. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1360332188.

Azizyan:2013:DSS

[ASW13] Martin Azizyan, Aarti Singh, and Larry Wasserman. Densitysensitive semisupervised inference. Annals of Statistics, 41(2):751–771, April 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1368018172.

Atchade:2011:KEA

[Atc11] Yves F. Atchadé. Kernel estimators of asymptotic variance for adaptive Markov chain Monte Carlo. Annals of Statistics, 39(2):990–1011, April 2011. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1302268084.

Atchade:2017:CPS

[Atc17] Yves A. Atchadé. On the contraction properties of some highdimensional quasi-posterior distributions. Annals of Statistics, 45(5):2248–2273, October 2017. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

Anandkumar:2012:HDS

[ATHW12] Animashree Anandkumar, Vincent Y. F. Tan, Furong Huang, and Alan S. Willsky. High-dimensional structure estimation in Ising models: Local separation criterion. Annals of Statistics, 40(3):1346–1375, June 2012. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1344610586.

Athey:2019:GRF

[ATW19] Susan Athey, Julie Tibshirani, and Stefan Wager. Generalized random forests. Annals of Statistics, 47(2):1148–1178, April 2019. CODEN ASTSC7. ISSN 0090-5364. URL http://

projecteuclid.org/euclid.aos/1547197251.

Anandkumar:2013:LLG

[AV13] Animashree Anandkumar and Ragupathyraj Valluvan. Learning loopy graphical models with latent variables: Efficient methods and guarantees. Annals of Statistics, 41(2):401–435, April 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1366138196.

Amini:2012:SFF

[AW12] Arash A. Amini and Martin J. Wainwright. Sampled forms of functional PCA in reproducing kernel Hilbert spaces. Annals of Statistics, 40(5):2483–2510, October 2012. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1359987528.

Antognini:2012:MOO

[AZ12] Alessandro Baldi Antognini and Maroussa Zagoraiou. Multiobjective optimal designs in comparative clinical trials with covariates: The reinforced doubly adaptive biased coin design. Annals of Statistics, 40(3):1315–1345, June 2012. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1344610585.

Bacallado:2011:BAV

[Bac11] Sergio Bacallado. Bayesian analysis of variable-order, reversible Markov chains. Annals of Statistics, 39(2):838–864, April 2011. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1299680956.

Bao:2019:TWL

[Bao19] Zhigang Bao. Tracy–Widom limit for Kendall’s tau. Annals of Statistics, 47(6):3504–3532, December 2019. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/euclid.aos/
1572487401.

Brien:2010:DTE

[BB10] C. J. Brien and R. A. Bailey. Decomposition tables for experiments. II. two–one randomizations. Annals of Statistics, 38(5):3164–3190, October 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1284391761.

Buja:2014:DST

[BB14] A. Buja and L. Brown. Discussion: “A significance test for the lasso”. Annals of Statistics, 42(2):509–517, April 2014. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/

1400592167. See [LTTT14c].

Bhattacharyya:2015:SBC

[BB15] Sharmodeep Bhattacharyya and Peter J. Bickel. Subsampling bootstrap of count features of networks. Annals of Statistics, 43(6):2384–2411, December 2015. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1444222079.

Bailey:2016:RBM

[BB16a] R. A. Bailey and C. J. Brien. Randomization-based models for multitiered experiments: I. A chain of randomizations. Annals of Statistics, 44(3):1131–1164, June 2016. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1460381689.

Bhattacharjee:2016:LSB

[BB16b] Monika Bhattacharjee and Arup Bose. Large sample behaviour of high dimensional autocovariance matrices. Annals of Statistics, 44(2):598–628, April 2016. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1458245729.

Baraud:2018:RER

[BB18a] Yannick Baraud and Lucien Birgé. Rho-estimators revisited: General theory and applications. Annals of Statistics, 46 (6B):3767–3804, December 2018. CODEN ASTSC7. ISSN

0090-5364. URL http://projecteuclid.org/euclid.aos/
1536631290.

Berghaus:2018:WCP

[BB18b] Betina Berghaus and Axel Bücher. Weak convergence of a pseudo maximum likelihood estimator for the extremal index. Annals of Statistics, 46(5):2307–2335, October 2018. CODEN ASTSC7. ISSN 0090-5364. URL http://projecteuclid.

org/euclid.aos/1534492837.

Berk:2013:VPS

[BBB^{+}13] Richard Berk, Lawrence Brown, Andreas Buja, Kai Zhang, and Linda Zhao. Valid post-selection inference. Annals of Statistics, 41(2):802–837, April 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1369836961.

Bayarri:2012:CBM

[BBFGD12] M. J. Bayarri, J. O. Berger, A. Forte, and G. García-Donato. Criteria for Bayesian model choice with application to variable selection. Annals of Statistics, 40(3):1550– 1577, June 2012. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1346850065.

Bali:2011:RFP

[BBTW11] Juan Lucas Bali, Graciela Boente, David E. Tyler, and Jane-Ling Wang. Robust functional principal components: A projection-pursuit approach. Annals of Statistics, 39 (6):2852–2882, December 2011. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1327413771.

Belloni:2011:SPQ

[BC11a] Alexandre Belloni and Victor Chernozhukov. ℓ_{1}-penalized quantile regression in high-dimensional sparse models. Annals of Statistics, 39(1):82–130, February 2011. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1291388370.

Buhlmann:2011:ILS

[BC11b] Peter Bühlmann and Tony Cai. Introduction to the Lehmann special section. Annals of Statistics, 39(5):2243, October 2011. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1321020522.

Barber:2015:CFD

[BC15] Rina Foygel Barber and Emmanuel J. Candès. Controlling the false discovery rate via knockoffs. Annals of Statistics, 43(5):2055–2085, October 2015. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1438606853.

Barber:2019:KFH

[BC19] Rina Foygel Barber and Emmanuel J. Candès. A knockoff filter for high-dimensional selective inference. Annals of Statistics, 47(5):2504–2537, October 2019. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https:

//projecteuclid.org/euclid.aos/1564797855.

Belloni:2018:UVP

[BCCW18] Alexandre Belloni, Victor Chernozhukov, Denis Chetverikov, and Ying Wei. Uniformly valid post-regularization confidence regions for many functional parameters in z-estimation framework. Annals of Statistics, 46(6B):3643–3675, December 2018. CODEN ASTSC7. ISSN 0090-5364. URL http:/

/projecteuclid.org/euclid.aos/1536631286.

Bickel:2013:ANM

[BCCZ13] Peter Bickel, David Choi, Xiangyu Chang, and Hai Zhang. Asymptotic normality of maximum likelihood and its variational approximation for stochastic blockmodels. Annals of Statistics, 41(4):1922–1943, August 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1382547508.

Binev:2014:CAU

[BCDD14] Peter Binev, Albert Cohen, Wolfgang Dahmen, and Ronald DeVore. Classification algorithms using adaptive partitioning. Annals of Statistics, 42(6):2141–2163, December 2014. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1413810724.

Bai:2018:CAB

[BCF18a] Zhidong Bai, Kwok Pui Choi, and Yasunori Fujikoshi. Consistency of AIC and BIC in estimating the number of significant components in high-dimensional principal component analysis. Annals of Statistics, 46(3):1050–1076, June 2018. CODEN ASTSC7. ISSN 0090-5364. URL http://

projecteuclid.org/euclid.aos/1525313075.

Bai:2018:LBE

[BCF18b] Zhidong Bai, Kwok Pui Choi, and Yasunori Fujikoshi. Limiting behavior of eigenvalues in high-dimensional MANOVA via RMT. Annals of Statistics, 46(6A):2985–3013, December 2018. CODEN ASTSC7. ISSN 0090-5364. URL http://

projecteuclid.org/euclid.aos/1536307240.

Bogdan:2011:ABO

[BCFG11] Małgorzata Bogdan, Arijit Chakrabarti, Florian Frommlet, and Jayanta K. Ghosh. Asymptotic Bayes-optimality under sparsity of some multiple testing procedures. Annals of Statistics, 39(3):1551–1579, June 2011. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1307452128.

Bickel:2011:MMD

[BCL11] Peter J. Bickel, Aiyou Chen, and Elizaveta Levina. The method of moments and degree distributions for network models. Annals of Statistics, 39(5):2280–2301, October 2011. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1321020525.

Biau:2010:SAK

[BCR10] Gérard Biau, Benoît Cadre, and Laurent Rouvière. Statistical analysis of k-nearest neighbor collaborative recommendation. Annals of Statistics, 38(3):1568–1592, June 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1269452647.

Belloni:2014:PES

[BCW14] Alexandre Belloni, Victor Chernozhukov, and Lie Wang. Pivotal estimation via square-root Lasso in nonparametric regression. Annals of Statistics, 42(2):757–788, April 2014. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1400592177.

Brown:2010:NRE

[BCZ10] Lawrence D. Brown, T. Tony Cai, and Harrison H. Zhou. Nonparametric regression in exponential families. Annals of Statistics, 38(4):2005–2046, August 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1278861241.

Bickel:2015:CPC

[BCZ^{+}15] Peter J. Bickel, Aiyou Chen, Yunpeng Zhao, Elizaveta Levina, and Ji Zhu. Correction to the proof of consistency of community detection. Annals of Statistics, 43(1):462–466, February 2015. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1423230086.

Buta:2011:CAE

[BD11] Eugenia Buta and Hani Doss. Computational approaches for empirical Bayes methods and Bayesian sensitivity analysis. Annals of Statistics, 39(5):2658–2685, October 2011. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1324563351.

Braess:2013:ODD

[BD13] Dietrich Braess and Holger Dette. Optimal discriminating designs for several competing regression models. Annals of Statistics, 41(2):897–922, April 2013. CODEN ASTSC7.

ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1369836964.

Bailey:2014:OCD

[BD14] R. A. Bailey and P. Druilhet. Optimal cross-over designs for full interaction models. Annals of Statistics, 42(6):2282–2300, December 2014. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1413810728.

Byrne:2015:SMG

[BD15] Simon Byrne and A. Philip Dawid. Structural Markov graph laws for Bayesian model uncertainty. Annals of Statistics, 43(4):1647–1681, August 2015. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1434546218.

Brockwell:2010:SIM

[BDD10] Anthony Brockwell, Pierre Del Moral, and Arnaud Doucet. Sequentially interacting Markov chain Monte Carlo methods. Annals of Statistics, 38(6):3387–3411, December 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1291126961.

Bodnar:2019:TIL

[BDP19] Taras Bodnar, Holger Dette, and Nestor Parolya. Testing for independence of large dimensional vectors. Annals of Statistics, 47(5):2977–3008, October 2019. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https:

//projecteuclid.org/euclid.aos/1564797870.

Banerjee:2019:DCN

[BDS19] Moulinath Banerjee, Cécile Durot, and Bodhisattva Sen. Divide and conquer in nonstandard problems and the superefficiency phenomenon. Annals of Statistics, 47(2):720–757, April 2019. CODEN ASTSC7. ISSN 0090-5364. URL http:

//projecteuclid.org/euclid.aos/1547197236.

Bucher:2011:NEP

[BDV11] Axel Bücher, Holger Dette, and Stanislav Volgushev. New estimators of the Pickands dependence function and a test for extreme-value dependence. Annals of Statistics, 39 (4):1963–2006, August 2011. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1314190620.

Belomestny:2010:SEF

[Bel10] Denis Belomestny. Spectral estimation of the fractional order of a Lévy process. Annals of Statistics, 38(1):317–351, February 2010. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1262271617.

Belomestny:2011:SIT

[Bel11] Denis Belomestny. Statistical inference for time-changed Lévy processes via composite characteristic function estimation. Annals of Statistics, 39(4):2205–2242, August 2011. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1319595463.

Belitser:2017:CLR

[Bel17] Eduard Belitser. On coverage and local radial rates of credible sets. Annals of Statistics, 45(3):1124–1151, June 2017. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1497319690.

Bellec:2018:OBA

[Bel18a] Pierre C. Bellec. Optimal bounds for aggregation of affine estimators. Annals of Statistics, 46(1):30–59, February 2018. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

Bellec:2018:SOI

[Bel18b] Pierre C. Bellec. Sharp oracle inequalities for least squares estimators in shape restricted regression. Annals of Statistics, 46(2):745–780, April 2018. CODEN ASTSC7. ISSN

0090-5364. URL http://projecteuclid.org/euclid.aos/
1522742435.

Bercu:2012:RMP

[BF12] Bernard Bercu and Philippe Fraysse. A Robbins–Monro procedure for estimation in semiparametric regression models. Annals of Statistics, 40(2):666–693, April 2012. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1337268208.

Brouste:2018:LAN

[BF18] Alexandre Brouste and Masaaki Fukasawa. Local asymptotic normality property for fractional Gaussian noise under high-frequency observations. Annals of Statistics, 46 (5):2045–2061, October 2018. CODEN ASTSC7. ISSN

0090-5364. URL http://projecteuclid.org/euclid.aos/
1534492828.

Bevilacqua:2019:EPU

[BFFP19] Moreno Bevilacqua, Tarik Faouzi, Reinhard Furrer, and Emilio Porcu. Estimation and prediction using generalized Wendland covariance functions under fixed domain asymptotics. Annals of Statistics, 47(2):828–856, April 2019. CODEN ASTSC7. ISSN 0090-5364. URL http://

projecteuclid.org/euclid.aos/1547197240.

Bradic:2011:RCP

[BFJ11] Jelena Bradic, Jianqing Fan, and Jiancheng Jiang. Regularization for Cox’s proportional hazards model with NPdimensionality. Annals of Statistics, 39(6):3092–3120, December 2011. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1327672847.

Battey:2018:DTE

[BFL^{+}18] Heather Battey, Jianqing Fan, Han Liu, Junwei Lu, and Ziwei Zhu. Distributed testing and estimation under sparse high dimensional models. Annals of Statistics, 46(3):1352–1382, June 2018. CODEN ASTSC7. ISSN 0090-5364. URL http:

//projecteuclid.org/euclid.aos/1525313085.

Bierkens:2019:ZZP

[BFR19] Joris Bierkens, Paul Fearnhead, and Gareth Roberts. The Zig–Zag process and super-efficient sampling for Bayesian analysis of big data. Annals of Statistics, 47(3):1288–1320, June 2019. CODEN ASTSC7. ISSN 0090-5364. URL http:

//projecteuclid.org/euclid.aos/1550026838.

Bacallado:2013:BNA

[BFT13] Sergio Bacallado, Stefano Favaro, and Lorenzo Trippa. Bayesian nonparametric analysis of reversible Markov chains. Annals of Statistics, 41(2):870–896, April 2013. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1369836963.

Bigot:2010:DAE

[BG10] Jérémie Bigot and Sébastien Gadat. A deconvolution approach to estimation of a common shape in a shifted curves model. Annals of Statistics, 38(4):2422–2464, August 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1278861253.

Bigot:2013:MPF

[BG13] Jérémie Bigot and Xavier Gendre. Minimax properties of Fréchet means of discretely sampled curves. Annals of Statistics, 41(2):923–956, April 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1369836965.

Bochkina:2014:BMT

[BG14] Natalia A. Bochkina and Peter J. Green. The Bernstein–von Mises theorem and nonregular models. Annals of Statistics, 42(5):1850–1878, October 2014. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1410440627.

Botev:2010:KDE

[BGK10] Z. I. Botev, J. F. Grotowski, and D. P. Kroese. Kernel density estimation via diffusion. Annals of Statistics, 38 (5):2916–2957, October 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1281964340.

Butucea:2018:LAE

[BGN18] Cristina Butucea, Mădălin Guţă, and Michael Nussbaum. Local asymptotic equivalence of pure states ensembles and quantum Gaussian white noise. Annals of Statistics, 46 (6B):3676–3706, December 2018. CODEN ASTSC7. ISSN

0090-5364. URL http://projecteuclid.org/euclid.aos/
1536631287.

Belitser:2012:OTS

[BGvZ12] Eduard Belitser, Subhashis Ghosal, and Harry van Zanten. Optimal two-stage procedures for estimating location and size of the maximum of a multivariate regression function. Annals of Statistics, 40(6):2850–2876, December 2012. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1360332186. See correction [BGvZ21].

Belitser:2021:CNS

[BGvZ21] Eduard Belitser, Subhashis Ghosal, and Harry van Zanten. Correction note: “Optimal two-stage procedures for estimating location and size of the maximum of a multivariate regression function” ann. statist. 40 (2012) 2850-2876. Annals of Statistics, 49(1):612–613, February 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/journals/
annals-of-statistics/volume-49/issue-1/Correction-
note--Optimal-two-stage-procedures-for-estimating-
location/10.1214/20-AOS1993.full. See [BGvZ12].

Behr:2018:MBS

[BHM18] Merle Behr, Chris Holmes, and Axel Munk. Multiscale blind source separation. Annals of Statistics, 46(2):711–744, April 2018. CODEN ASTSC7. ISSN 0090-5364. URL http://

projecteuclid.org/euclid.aos/1522742434.

Bibinger:2014:EQC

[BHMR14] Markus Bibinger, Nikolaus Hautsch, Peter Malec, and Markus Reiß. Estimating the quadratic covariation matrix from noisy observations: Local method of moments and efficiency. Annals of Statistics, 42(4):1312–1346, August 2014. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1403715202.

Bao:2019:CCC

[BHPZ19] Zhigang Bao, Jiang Hu, Guangming Pan, and Wang Zhou. Canonical correlation coefficients of high-dimensional Gaussian vectors: Finite rank case. Annals of Statistics, 47(1):612– 640, February 2019. CODEN ASTSC7. ISSN 0090-5364. URL

http://projecteuclid.org/euclid.aos/1543568600.

Brownlees:2015:ERM

[BJL15] Christian Brownlees, Emilien Joly, and Gábor Lugosi. Empirical risk minimization for heavy-tailed losses. Annals of Statistics, 43(6):2507–2536, December 2015. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1444222083.

Birnbaum:2013:MBS

[BJNP13] Aharon Birnbaum, Iain M. Johnstone, Boaz Nadler, and Debashis Paul. Minimax bounds for sparse PCA with noisy high-dimensional data. Annals of Statistics, 41(3):1055–1084, June 2013. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1371150893.

Bonhomme:2016:EML

[BJR16] Stéphane Bonhomme, Koen Jochmans, and Jean-Marc Robin. Estimating multivariate latent-structure models. Annals of Statistics, 44(2):540–563, April 2016. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1458245727.

Bibinger:2017:NCP

[BJV17] Markus Bibinger, Moritz Jirak, and Mathias Vetter. Nonparametric change-point analysis of volatility. Annals of Statistics, 45(4):1542–1578, August 2017. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1498636866.

Bickel:2012:SBM

[BK12] P. J. Bickel and B. J. K. Kleijn. The semiparametric Bernstein–von Mises theorem. Annals of Statistics, 40 (1):206–237, February 2012. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1333029963.

Barber:2018:RRB

[BK18] Rina Foygel Barber and Mladen Kolar. ROCKET: Robust confidence intervals via Kendall’s tau for transelliptical graphical models. Annals of Statistics, 46(6B):3422–3450, December 2018. CODEN ASTSC7. ISSN 0090-5364. URL

http://projecteuclid.org/euclid.aos/1536631279.

Bauer:2019:DLR

[BK19] Benedikt Bauer and Michael Kohler. On deep learning as a remedy for the curse of dimensionality in nonparametric regression. Annals of Statistics, 47(4):2261–2285, August 2019. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/euclid.aos/
1558425645.

Bertsimas:2016:BSS

[BKM16] Dimitris Bertsimas, Angela King, and Rahul Mazumder. Best subset selection via a modern optimization lens. Annals of Statistics, 44(2):813–852, April 2016. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1458245736.

Bottcher:2019:DMN

[BKRS19] Björn Böttcher, Martin Keller-Ressel, and René L. Schilling. Distance multivariance: New dependence measures for random vectors. Annals of Statistics, 47(5):2757–2789, October 2019. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/euclid.
aos/1564797863.

Bai:2012:SAF

[BL12] Jushan Bai and Kunpeng Li. Statistical analysis of factor models of high dimension. Annals of Statistics, 40 (1):436–465, February 2012. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1334581749.

Bai:2014:TMP

[BL14] Jushan Bai and Kunpeng Li. Theory and methods of panel data models with interactive effects. Annals of Statistics, 42(1):142–170, February 2014. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1392733183.

Bandyopadhyay:2015:FDE

[BLN15] Soutir Bandyopadhyay, Soumendra N. Lahiri, and Daniel J. Nordman. A frequency domain empirical likelihood method for irregularly spaced spatial data. Annals of Statistics, 43(2):519–545, April 2015. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1424787427.

Bachoc:2019:VCI

[BLP19] François Bachoc, Hannes Leeb, and Benedikt M. Pötscher. Valid confidence intervals for post-model-selection predictors. Annals of Statistics, 47(3):1475–1504, June 2019. CODEN ASTSC7. ISSN 0090-5364. URL http://projecteuclid.

org/euclid.aos/1550026846.

Bao:2015:SSL

[BLPZ15] Zhigang Bao, Liang-Ching Lin, Guangming Pan, and Wang Zhou. Spectral statistics of large dimensional Spearman’s rank correlation matrix and its application. Annals of Statistics, 43(6):2588–2623, December 2015. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1444222086.

Boistard:2017:FCL

[BLRG17] Hélène Boistard, Hendrik P. Lopuhaä, and Anne Ruiz-Gazen. Functional central limit theorems for single-stage sampling designs. Annals of Statistics, 45(4):1728–1758, August 2017. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1498636872.

Bellec:2018:SML

[BLT18] Pierre C. Bellec, Guillaume Lecué, and Alexandre B. Tsybakov. Slope meets Lasso: Improved oracle bounds and optimality. Annals of Statistics, 46(6B):3603–3642, December 2018. CODEN ASTSC7. ISSN 0090-5364. URL http://

projecteuclid.org/euclid.aos/1536631285.

Bertsimas:2014:LQR

[BM14] Dimitris Bertsimas and Rahul Mazumder. Least quantile regression via modern optimization. Annals of Statistics, 42 (6):2494–2525, December 2014. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1415801781.

Basu:2015:RES

[BM15] Sumanta Basu and George Michailidis. Regularized estimation in sparse high-dimensional time series models. Annals of Statistics, 43(4):1535–1567, August 2015. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1434546214.

Basu:2017:ANS

[BM17] Kinjal Basu and Rajarshi Mukherjee. Asymptotic normality of scrambled geometric net quadrature. Annals of Statistics, 45(4):1759–1788, August 2017. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1498636873.

Bruna:2015:IPA

[BMBM15] Joan Bruna, Stéphane Mallat, Emmanuel Bacry, and Jean-François Muzy. Intermittent process analysis with scattering moments. Annals of Statistics, 43(1):323–351, February 2015. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1423230082.

Buhlmann:2014:DST

[BMvdG14] Peter Bühlmann, Lukas Meier, and Sara van de Geer. Discussion: “A significance test for the lasso”. Annals of Statistics, 42(2):469–477, April 2014. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1400592162. See

[LTTT14c].

Brown:2018:EBE

[BMW18] Lawrence D. Brown, Gourab Mukherjee, and Asaf Weinstein. Empirical Bayes estimates for a two-way crossclassified model. Annals of Statistics, 46(4):1693–1720, August 2018. CODEN ASTSC7. ISSN 0090-5364. URL http:

//projecteuclid.org/euclid.aos/1530086430.

Butucea:2018:VSH

[BNST18] Cristina Butucea, Mohamed Ndaoud, Natalia A. Stepanova, and Alexandre B. Tsybakov. Variable selection with Hamming loss. Annals of Statistics, 46(5):1837–1875, October 2018. CODEN ASTSC7. ISSN 0090-5364. URL http://

projecteuclid.org/euclid.aos/1534492821.

Belloni:2017:AGC

[BO17] Alexandre Belloni and Roberto I. Oliveira. Approximate group context tree. Annals of Statistics, 45(1):355–385, February 2017. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1487667626.

Bobkov:2019:KTE

[Bob19] Sergey G. Bobkov. Khinchine’s theorem and Edgeworth approximations for weighted sums. Annals of Statistics, 47(3):1616–1633, June 2019. CODEN ASTSC7. ISSN

0090-5364. URL http://projecteuclid.org/euclid.aos/
1550026851.

Bontemps:2011:BMT

[Bon11] Dominique Bontemps. Bernstein–von Mises theorems for Gaussian regression with increasing number of regressors. Annals of Statistics, 39(5):2557–2584, October 2011. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1322663468.

Bowsher:2010:SKM

[Bow10] Clive G. Bowsher. Stochastic kinetic models: Dynamic independence, modularity and graphs. Annals of Statistics, 38(4):2242–2281, August 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1278861248.

Bhattacharya:2014:AFE

[BPD14] Anirban Bhattacharya, Debdeep Pati, and David Dunson. Anisotropic function estimation using multi-bandwidth Gaussian processes. Annals of Statistics, 42(1):352–381, February 2014. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1395234981.

Buhlmann:2014:CCA

[BPE14] Peter Bühlmann, Jonas Peters, and Jan Ernest. CAM: Causal additive models, high-dimensional order search and penalized regression. Annals of Statistics, 42(6):2526–2556, December 2014. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1415801782.

Bhattacharya:2019:BFP

[BPY19] Anirban Bhattacharya, Debdeep Pati, and Yun Yang. Bayesian fractional posteriors. Annals of Statistics, 47(1):39– 66, February 2019. CODEN ASTSC7. ISSN 0090-5364. URL

http://projecteuclid.org/euclid.aos/1543568581.

Bao:2015:ULE

[BPZ15] Zhigang Bao, Guangming Pan, and Wang Zhou. Universality for the largest eigenvalue of sample covariance matrices with general population. Annals of Statistics, 43(1):382–421, February 2015. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1423230084.

Bi:2018:MTF

[BQS18] Xuan Bi, Annie Qu, and Xiaotong Shen. Multilayer tensor factorization with applications to recommender systems. Annals of Statistics, 46(6B):3308–3333, December 2018. CODEN ASTSC7. ISSN 0090-5364. URL http://

projecteuclid.org/euclid.aos/1536631275.

Berthet:2013:ODS

[BR13] Quentin Berthet and Philippe Rigollet. Optimal detection of sparse principal components in high dimension. Annals of Statistics, 41(4):1780–1815, August 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1378386239.

Berthet:2019:ERI

[BRS19] Quentin Berthet, Philippe Rigollet, and Piyush Srivastava. Exact recovery in the Ising blockmodel. Annals of Statistics, 47(4):1805–1834, August 2019. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://

projecteuclid.org/euclid.aos/1558425631.

Bhaskar:2014:DRS

[BS14] Anand Bhaskar and Yun S. Song. Descartes’ rule of signs and the identifiability of population demographic models from genomic variation data. Annals of Statistics, 42(6):2469–2493, December 2014. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1413810735.

Bucher:2014:WUW

[BSV14] Axel Bücher, Johan Segers, and Stanislav Volgushev. When uniform weak convergence fails: Empirical processes for dependence functions and residuals via epi- and hypographs. Annals of Statistics, 42(4):1598–1634, August 2014. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1407420010.

Bunea:2011:OSR

[BSW11] Florentina Bunea, Yiyuan She, and Marten H. Wegkamp. Optimal selection of reduced rank estimators of high-dimensional matrices. Annals of Statistics, 39(2):1282–1309, April 2011. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1304947051.

Bunea:2012:JVR

[BSW12] Florentina Bunea, Yiyuan She, and Marten H. Wegkamp. Joint variable and rank selection for parsimonious estimation of high-dimensional matrices. Annals of Statistics, 40 (5):2359–2388, October 2012. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1359987524.

Berrett:2019:EME

[BSY19] Thomas B. Berrett, Richard J. Samworth, and Ming Yuan. Efficient multivariate entropy estimation via k-nearest neighbour distances. Annals of Statistics, 47(1):288–318, February 2019. CODEN ASTSC7. ISSN 0090-5364. URL http://

projecteuclid.org/euclid.aos/1543568589.

Bai:2017:VRM

[BT17] Shuyang Bai and Murad S. Taqqu. On the validity of resampling methods under long memory. Annals of Statistics, 45(6):2365–2399, December 2017. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

Bien:2013:LHI

[BTT13] Jacob Bien, Jonathan Taylor, and Robert Tibshirani. A lasso for hierarchical interactions. Annals of Statistics, 41(3):1111– 1141, June 2013. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1371150895.

Bunea:2010:SMM

[BTWB10] Florentina Bunea, Alexandre B. Tsybakov, Marten H. Wegkamp, and Adrian Barbu. SPADES and mixture models. Annals of Statistics, 38(4):2525–2558, August 2010. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1278861256.

Bull:2013:SAS

[Bul13] Adam D. Bull. Spatially-adaptive sensing in nonparametric regression. Annals of Statistics, 41(1):41–62, February 2013. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1362493039.

Bull:2014:ETC

[Bul14] Adam D. Bull. Estimating time-changes in noisy Lévy models. Annals of Statistics, 42(5):2026–2057, October 2014. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1410440633.

Bull:2016:NOE

[Bul16] Adam D. Bull. Near-optimal estimation of jump activity in semimartingales. Annals of Statistics, 44(1):58–86, February 2016. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1449755957.

Bull:2017:SDG

[Bul17] Adam D. Bull. Semimartingale detection and goodness-of-fit tests. Annals of Statistics, 45(3):1254–1283, June 2017. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1497319694.

Bucher:2013:NIL

[BV13] Axel Bücher and Mathias Vetter. Nonparametric inference on Lévy measures and copulas. Annals of Statistics, 41(3):1485– 1515, June 2013. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1375362557.

Belloni:2011:MQU

[BW11] Alexandre Belloni and Robert L. Winkler. On multivariate quantiles under partial orders. Annals of Statistics, 39(2):1125–1179, April 2011. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1304947046.

Balakrishnan:2019:HTD

[BW19a] Sivaraman Balakrishnan and Larry Wasserman. Hypothesis testing for densities and high-dimensional multinomials: Sharp local minimax rates. Annals of Statistics, 47 (4):1893–1927, August 2019. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://

projecteuclid.org/euclid.aos/1558425634.

Bing:2019:AER

[BW19b] Xin Bing and Marten H. Wegkamp. Adaptive estimation of the rank of the coefficient matrix in high-dimensional multivariate response regression models. Annals of Statistics, 47(6):3157–3184, December 2019. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://

projecteuclid.org/euclid.aos/1572487389.

Balakrishnan:2017:SGE

[BWY17] Sivaraman Balakrishnan, Martin J. Wainwright, and Bin Yu. Statistical guarantees for the EM algorithm: From population to sample-based analysis. Annals of Statistics, 45(1):77–120, February 2017. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1487667618.

Bathia:2010:IFD

[BYZ10] Neil Bathia, Qiwei Yao, and Flavio Ziegelmann. Identifying the finite dimensionality of curve time series. Annals of Statistics, 38(6):3352–3386, December 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1291126960.

Cuesta-Albertos:2019:GFT

[CAGPFBGM19] Juan A. Cuesta-Albertos, Eduardo García-Portugués, Manuel Febrero-Bande, and Wenceslao González-Manteiga. Goodnessof-fit tests for the functional linear model based on randomly projected empirical processes. Annals of Statistics, 47(1):439– 467, February 2019. CODEN ASTSC7. ISSN 0090-5364. URL

http://projecteuclid.org/euclid.aos/1543568594.

Castillo:2014:BSN

[Cas14] Ismaël Castillo. On Bayesian supremum norm contraction rates. Annals of Statistics, 42(5):2058–2091, October 2014. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1410440634.

Castillo:2015:DFC

[Cas15] Ismaël Castillo. Discussion of “Frequentist coverage of adaptive nonparametric Bayesian credible sets”. Annals of Statistics, 43(4):1437–1443, August 2015. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1434546207. See

[SvdVvZ15a].

Chang:2011:AML

[CC11] Jinyuan Chang and Song Xi Chen. On the approximate maximum likelihood estimation for diffusion processes. Annals of Statistics, 39(6):2820–2851, December 2011. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1327413770.

Chakraborty:2017:THD

[CC17] Anirvan Chakraborty and Probal Chaudhuri. Tests for highdimensional data based on means, spatial signs and spatial ranks. Annals of Statistics, 45(2):771–799, April 2017. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1494921957.

Chakrabortty:2018:EAL

[CC18a] Abhishek Chakrabortty and Tianxi Cai. Efficient and adaptive linear regression in semi-supervised settings. Annals of Statistics, 46(4):1541–1572, August 2018. CODEN ASTSC7.

ISSN 0090-5364. URL http://projecteuclid.org/euclid.
aos/1530086425.

Chang:2018:BAS

[CC18b] Ming-Chung Chang and Ching-Shui Cheng. A Bayesian approach to the selection of two-level multi-stratum factorial designs. Annals of Statistics, 46(4):1779–1806, August 2018. CODEN ASTSC7. ISSN 0090-5364. URL http://

projecteuclid.org/euclid.aos/1530086433.

Chu:2019:ADF

[CC19] Lynna Chu and Hao Chen. Asymptotic distribution-free change-point detection for multivariate and non-Euclidean data. Annals of Statistics, 47(1):382–414, February 2019. CODEN ASTSC7. ISSN 0090-5364. URL http://

projecteuclid.org/euclid.aos/1543568592.

Chang:2019:SAG

[CCC19] Ming-Chung Chang, Shao-Wei Cheng, and Ching-Shui Cheng. Signal aliasing in Gaussian random fields for experiments with qualitative factors. Annals of Statistics, 47(2):909–935, April 2019. CODEN ASTSC7. ISSN

0090-5364. URL http://projecteuclid.org/euclid.aos/
1547197243.

Cholaquidis:2014:PCP

[CCF14] Alejandro Cholaquidis, Antonio Cuevas, and Ricardo Fraiman. On Poincaré cone property. Annals of Statistics, 42(1):255–284, February 2014. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1395234978.

Carlier:2016:VQR

[CCG16] Guillaume Carlier, Victor Chernozhukov, and Alfred Galichon. Vector quantile regression: An optimal transport approach. Annals of Statistics, 44(3):1165–1192, June 2016. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1460381690.

Cardot:2017:OEG

[CCGB17] Hervé Cardot, Peggy Cénac, and Antoine Godichon-Baggioni. Online estimation of the geometric median in Hilbert spaces: Nonasymptotic confidence balls. Annals of Statistics, 45(2):591–614, April 2017. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1494921951.

Chernozhukov:2013:GAM

[CCK13] Victor Chernozhukov, Denis Chetverikov, and Kengo Kato. Gaussian approximations and multiplier bootstrap for maxima of sums of high-dimensional random vectors. Annals of Statistics, 41(6):2786–2819, December 2013. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1387313390.

Chernozhukov:2014:ACH

[CCK14a] Victor Chernozhukov, Denis Chetverikov, and Kengo Kato. Anti-concentration and honest, adaptive confidence bands. Annals of Statistics, 42(5):1787–1818, October 2014. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1410440625.

Chernozhukov:2014:GAS

[CCK14b] Victor Chernozhukov, Denis Chetverikov, and Kengo Kato. Gaussian approximation of suprema of empirical processes. Annals of Statistics, 42(4):1564–1597, August 2014. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1407420009.

Collier:2017:MEL

[CCT17] Olivier Collier, Laëtitia Comminges, and Alexandre B. Tsybakov. Minimax estimation of linear and quadratic functionals on sparsity classes. Annals of Statistics, 45(3):923– 958, June 2017. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1497319684.

Collier:2018:OAE

[CCTV18] Olivier Collier, Laëtitia Comminges, Alexandre B. Tsybakov, and Nicolas Verzelen. Optimal adaptive estimation of linear functionals under sparsity. Annals of Statistics, 46 (6A):3130–3150, December 2018. CODEN ASTSC7. ISSN

0090-5364. URL http://projecteuclid.org/euclid.aos/
1536307245.

Comminges:2012:TCC

[CD12] Laëtitia Comminges and Arnak S. Dalalyan. Tight conditions for consistency of variable selection in the context of high dimensionality. Annals of Statistics, 40(5):2667–2696, October 2012. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1359987534.

Chatterjee:2013:EUE

[CD13] Sourav Chatterjee and Persi Diaconis. Estimating and understanding exponential random graph models. Annals of Statistics, 41(5):2428–2461, October 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1383661269.

Constantinou:2017:ECI

[CD17] Panayiota Constantinou and A. Philip Dawid. Extended conditional independence and applications in causal inference. Annals of Statistics, 45(6):2618–2653, December 2017. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

Carroll:2013:UPB

[CDH13] Raymond J. Carroll, Aurore Delaigle, and Peter Hall. Unexpected properties of bandwidth choice when smoothing discrete data for constructing a functional data classifier. Annals of Statistics, 41(6):2739–2767, December 2013. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1387313388.

Chen:2011:CMC

[CDO11] S. Chen, J. Dick, and A. B. Owen. Consistency of Markov chain quasi-Monte Carlo on continuous state spaces. Annals of Statistics, 39(2):673–701, April 2011. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1299680951.

Cai:2011:EER

[CEdH11] Juan-Juan Cai, John H. J. Einmahl, and Laurens de Haan. Estimation of extreme risk regions under multivariate regular variation. Annals of Statistics, 39(3):1803–1826, June 2011. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1311600284.

Can:2015:ADF

[CEKL15] Sami Umut Can, John H. J. Einmahl, Estate V. Khmaladze, and Roger J. A. Laeven. Asymptotically distributionfree goodness-of-fit testing for tail copulas. Annals of Statistics, 43(2):878–902, April 2015. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1427115290.

Cai:2016:GTA

[CEL16] T. Tony Cai, Yonina C. Eldar, and Xiaodong Li. Global testing against sparse alternatives in time-frequency analysis. Annals of Statistics, 44(4):1438–1466, August 2016. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1467894704.

Cheng:2016:PHC

[CF16] Ming-Yen Cheng and Jianqing Fan. Peter Hall’s contributions to nonparametric function estimation and modeling. Annals of Statistics, 44(5):1837–1853, October 2016. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1473685259.

Cook:2019:PLS

[CF19] R. Dennis Cook and Liliana Forzani. Partial least squares prediction in high-dimensional regression. Annals of Statistics, 47(2):884–908, April 2019. CODEN ASTSC7. ISSN

0090-5364. URL http://projecteuclid.org/euclid.aos/
1547197242.

Cook:2012:ESR

[CFR12] R. Dennis Cook, Liliana Forzani, and Adam J. Rothman. Estimating sufficient reductions of the predictors in abundant high-dimensional regressions. Annals of Statistics, 40 (1):353–384, February 2012. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1333567193.

Cai:2017:CIH

[CG17] T. Tony Cai and Zijian Guo. Confidence intervals for highdimensional linear regression: Minimax rates and adaptivity. Annals of Statistics, 45(2):615–646, April 2017. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1494921952.

Cai:2018:AAH

[CG18] T. Tony Cai and Zijian Guo. Accuracy assessment for high-dimensional linear regression. Annals of Statistics, 46 (4):1807–1836, August 2018. CODEN ASTSC7. ISSN

0090-5364. URL http://projecteuclid.org/euclid.aos/
1530086434.

Comte:2011:ELP

[CGC11] Fabienne Comte and Valentine Genon-Catalot. Estimation for Lévy processes from high frequency data within a long time interval. Annals of Statistics, 39(2):803–837, April 2011. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1299680955.

Chernozhukov:2017:MKD

[CGHH17] Victor Chernozhukov, Alfred Galichon, Marc Hallin, and Marc Henry. Monge–Kantorovich depth, quantiles, ranks and signs. Annals of Statistics, 45(1):223–256, February 2017. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1487667622.

Cappe:2013:KLU

[CGM^{+}13] Olivier Cappé, Aurélien Garivier, Odalric-Ambrym Maillard, Rémi Munos, and Gilles Stoltz. Kullback–Leibler upper confidence bounds for optimal sequential allocation. Annals of Statistics, 41(3):1516–1541, June 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1375362558.

Chen:2018:RCS

[CGR18] Mengjie Chen, Chao Gao, and Zhao Ren. Robust covariance and scatter matrix estimation under Huber’s contamination model. Annals of Statistics, 46(5):1932–1960, October 2018. CODEN ASTSC7. ISSN 0090-5364. URL http://

projecteuclid.org/euclid.aos/1534492824.

Chatterjee:2015:RBI

[CGS15] Sabyasachi Chatterjee, Adityanand Guntuboyina, and Bodhisattva Sen. On risk bounds in isotonic and other shape restricted regression problems. Annals of Statistics, 43 (4):1774–1800, August 2015. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1434546222.

Cavaliere:2016:SBI

[CGT16] Giuseppe Cavaliere, Iliyan Georgiev, and A. M. Robert Taylor. Sieve-based inference for infinite-variance linear processes. Annals of Statistics, 44(4):1467–1494, August 2016. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1467894705.

Chen:2016:NMR

[CGTW16] Yen-Chi Chen, Christopher R. Genovese, Ryan J. Tibshirani, and Larry Wasserman. Nonparametric modal regression. Annals of Statistics, 44(2):489–514, April 2016. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1458245725.

Chen:2015:ATD

[CGW15] Yen-Chi Chen, Christopher R. Genovese, and Larry Wasserman. Asymptotic theory for density ridges. Annals of Statistics, 43(5):1896–1928, October 2015. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1438606848.

Cai:2018:AEP

[CGW18] T. Tony Cai, Adityanand Guntuboyina, and Yuting Wei. Adaptive estimation of planar convex sets. Annals of Statistics, 46(3):1018–1049, June 2018. CODEN ASTSC7. ISSN

0090-5364. URL http://projecteuclid.org/euclid.aos/
1525313074.

Chang:2018:PCA

[CGY18] Jinyuan Chang, Bin Guo, and Qiwei Yao. Principal component analysis for second-order stationary vector time series.

Annals of Statistics, 46(5):2094–2124, October 2018. CODEN ASTSC7. ISSN 0090-5364. URL http://projecteuclid.

org/euclid.aos/1534492830.

Cui:2018:THD

[CGZ18] Hengjian Cui, Wenwen Guo, and Wei Zhong. Test for high-dimensional regression coefficients using refitted cross-validation variance estimation. Annals of Statistics, 46(3):958–988, June 2018. CODEN ASTSC7. ISSN

0090-5364. URL http://projecteuclid.org/euclid.aos/
1525313072.

Cheng:2010:BCG

[CH10] Guang Cheng and Jianhua Z. Huang. Bootstrap consistency for general semiparametric M-estimation. Annals of Statistics, 38(5):2884–2915, October 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1279638543.

Cisewski:2012:GFI

[CH12] Jessi Cisewski and Jan Hannig. Generalized fiducial inference for normal linear mixed models. Annals of Statistics, 40(4):2102–2127, August 2012. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1351602538.

Chatterjee:2014:NPL

[Cha14] Sourav Chatterjee. A new perspective on least squares under convex constraint. Annals of Statistics, 42(6):2340–2381, December 2014. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1413810730.

Chatterjee:2015:MEU

[Cha15a] Sourav Chatterjee. Matrix estimation by Universal Singular Value Thresholding. Annals of Statistics, 43(1):177–214, February 2015. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1418135619.

Chauvet:2015:CMM

[Cha15b] Guillaume Chauvet. Coupling methods for multistage sampling. Annals of Statistics, 43(6):2484–2506, December 2015.

CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1444222082.

Chan:2017:OSD

[Cha17] Hock Peng Chan. Optimal sequential detection in multistream data. Annals of Statistics, 45(6):2736–2763, December 2017. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

Chen:2016:PHC

[Che16] Song Xi Chen. Peter Hall’s contributions to the bootstrap. Annals of Statistics, 44(5):1821–1836, October 2016. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1473685258.

Chen:2018:GBA

[Che18] Xiaohui Chen. Gaussian and bootstrap approximations for high-dimensional U-statistics and their applications. Annals of Statistics, 46(2):642–678, April 2018. CODEN ASTSC7.

ISSN 0090-5364. URL http://projecteuclid.org/euclid.
aos/1522742432.

Chen:2019:SCP

[Che19a] Hao Chen. Sequential change-point detection based on nearest neighbors. Annals of Statistics, 47(3):1381–1407, June 2019. CODEN ASTSC7. ISSN 0090-5364. URL http://

projecteuclid.org/euclid.aos/1550026842.

Chen:2019:PTS

[Che19b] Wei-Kuo Chen. Phase transition in the spiked random tensor with Rademacher prior. Annals of Statistics, 47 (5):2734–2756, October 2019. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://

projecteuclid.org/euclid.aos/1564797862.

Chen:2019:GCT

[Che19c] Yen-Chi Chen. Generalized cluster trees and singular measures. Annals of Statistics, 47(4):2174–2203, August 2019. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/euclid.aos/
1558425642.

Chi:2011:ESD

[Chi11] Zhiyi Chi. Effects of statistical dependence on multiple testing under a hidden Markov model. Annals of Statistics, 39(1):439–473, February 2011. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1297779853.

Chan:2013:MBM

[CHI13] Ngai Hang Chan, Shih-Feng Huang, and Ching-Kang Ing. Moment bounds and mean squared prediction errors of longmemory time series. Annals of Statistics, 41(3):1268–1298, June 2013. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1371150901.

Chang:2014:ATG

[CHI14] Chih-Hao Chang, Hsin-Cheng Huang, and Ching-Kang Ing. Asymptotic theory of generalized information criterion for geostatistical regression model selection. Annals of Statistics, 42(6):2441–2468, December 2014. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1413810733.

Cheng:2016:EES

[CHL16] Ming-Yen Cheng, Toshio Honda, and Jialiang Li. Efficient estimation in semivarying coefficient models for longitudinal/clustered data. Annals of Statistics, 44(5):1988–2017, October 2016. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1473685266.

Chernozhukov:2017:LAR

[CHL17] Victor Chernozhukov, Christian Hansen, and Yuan Liao. A lava attack on the recovery of sums of dense and sparse signals. Annals of Statistics, 45(1):39–76, February 2017. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1487667617.

Cheng:2014:NIS

[CHLP14] Ming-Yen Cheng, Toshio Honda, Jialiang Li, and Heng Peng. Nonparametric independence screening and structure identification for ultra-high dimensional longitudinal data. Annals of Statistics, 42(5):1819–1849, October 2014. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1410440626.

Chen:2011:SMI

[CHM11] Dong Chen, Peter Hall, and Hans-Georg Müller. Single and multiple index functional regression models with nonparametric link. Annals of Statistics, 39(3):1720–1747, June 2011. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1311600281.

Choi:2017:CCN

[Cho17] David Choi. Co-clustering of nonsmooth graphons. Annals of Statistics, 45(4):1488–1515, August 2017. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1498636864.

Cui:2011:EAS

[CHZ11] Xia Cui, Wolfgang Karl Härdle, and Lixing Zhu. The EFM approach for single-index models. Annals of Statistics, 39(3):1658–1688, June 2011. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1311600279.

Chan:2011:UMB

[CI11] Ngai Hang Chan and Ching-Kang Ing. Uniform moment bounds of Fisher’s information with applications to time series. Annals of Statistics, 39(3):1526–1550, June 2011. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1307452127.

Cai:2010:ORCa

[CJ10] T. Tony Cai and Jiashun Jin. Optimal rates of convergence for estimating the null density and proportion of nonnull effects in large-scale multiple testing. Annals of Statistics, 38(1):100–145, February 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1262271611.

Cai:2011:LLC

[CJ11] T. Tony Cai and Tiefeng Jiang. Limiting laws of coherence of random matrices with applications to testing covariance structure and construction of compressed sensing matrices. Annals of Statistics, 39(3):1496–1525, June 2011. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1305292044.

Comte:2012:AFL

[CJ12] Fabienne Comte and Jan Johannes. Adaptive functional linear regression. Annals of Statistics, 40(6):2765–2797, December 2012. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1360332183.

Cheng:2015:OED

[CK15] Ching-Shui Cheng and Ming-Hung Kao. Optimal experimental designs for fMRI via circulant biased weighing designs. Annals of Statistics, 43(6):2565–2587, December 2015. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1444222085.

Cao:2019:PGS

[CKG19] Xuan Cao, Kshitij Khare, and Malay Ghosh. Posterior graph selection and estimation consistency for high-dimensional Bayesian DAG models. Annals of Statistics, 47(1):319–348, February 2019. CODEN ASTSC7. ISSN 0090-5364. URL

http://projecteuclid.org/euclid.aos/1543568590.

Cai:2016:OLS

[CKW^{+}16] Tony Cai, Donggyu Kim, Yazhen Wang, Ming Yuan, and Harrison H. Zhou. Optimal large-scale quantum state tomography with Pauli measurements. Annals of Statistics, 44(2):682–712, April 2016. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1458245732.

Chan:2008:REP

[CL08] Ngai Hang Chan and Shiqing Ling. Residual empirical processes for long and short memory time series. Annals of Statistics, 36(5):2453–2470, October 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:/

/projecteuclid.org/euclid.aos/1223908099. See correc-

tion [CL10].

Chan:2010:CRE

[CL10] Ngai Hang Chan and Shiqing Ling. Correction: “Residual empirical processes for long and short memory time series”. Annals of Statistics, 38(6):3839, December 2010. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/

1291126976. See [CL08].

Cai:2011:TCH

[CL11] T. Tony Cai and Mark G. Low. Testing composite hypotheses, Hermite polynomials and optimal estimation of a nonsmooth functional. Annals of Statistics, 39(2):1012–1041, April 2011. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1302268085.

Chan:2013:GTP

[CL13a] Hock Peng Chan and Tze Leung Lai. A general theory of particle filters in hidden Markov models and some applications. Annals of Statistics, 41(6):2877–2904, December 2013. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1387313393.

Chatterjee:2013:RCA

[CL13b] A. Chatterjee and S. N. Lahiri. Rates of convergence of the adaptive LASSO estimators to the oracle distribution and higher order refinements by the bootstrap. Annals of Statistics, 41(3):1232–1259, June 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1371150899.

Chen:2013:QQF

[CL13c] Jiahua Chen and Yukun Liu. Quantile and quantile-function estimations under density ratio model. Annals of Statistics, 41(3):1669–1692, June 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1375362563.

Cai:2015:RCF

[CL15a] T. Tony Cai and Xiaodong Li. Robust and computationally feasible community detection in the presence of arbitrary outlier nodes. Annals of Statistics, 43(3):1027–1059, June 2015. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1431695637.

Chichignoud:2015:BSK

[CL15b] Michaël Chichignoud and Sébastien Loustau. Bandwidth selection in kernel empirical risk minimization via the gradient. Annals of Statistics, 43(4):1617–1646, August 2015. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1434546217.

Chen:2018:TIH

[CL18] Xi Chen and Weidong Liu. Testing independence with high-dimensional correlated samples. Annals of Statistics, 46(2):866–894, April 2018. CODEN ASTSC7. ISSN

0090-5364. URL http://projecteuclid.org/euclid.aos/
1522742439.

Chen:2015:SGA

[CLLW15] Jia Chen, Degui Li, Hua Liang, and Suojin Wang. Semiparametric GEE analysis in partially linear single-index models for longitudinal data. Annals of Statistics, 43(4):1682–1715, August 2015. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1434546219.

Cai:2016:ORC

[CLM16] T. Tony Cai, Xiaodong Li, and Zongming Ma. Optimal rates of convergence for noisy sparse phase retrieval via thresholded Wirtinger flow. Annals of Statistics, 44(5):2221–2251, October 2016. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1473685274.

Chen:2019:SRE

[CLN19] Ningyuan Chen, Donald K. K. Lee, and Sahand N. Negahban. Super-resolution estimation of cyclic arrival rates. Annals of Statistics, 47(3):1754–1775, June 2019. CODEN ASTSC7.

ISSN 0090-5364. URL http://projecteuclid.org/euclid.
aos/1550026856.

Camerlenghi:2019:DTH

[CLOP19] Federico Camerlenghi, Antonio Lijoi, Peter Orbanz, and Igor Prünster. Distribution theory for hierarchical processes. Annals of Statistics, 47(1):67–92, February 2019. CODEN ASTSC7. ISSN 0090-5364. URL http://projecteuclid.

org/euclid.aos/1543568582.

Cai:2016:GIG

[CLR16] T. Tony Cai, Tengyuan Liang, and Alexander Rakhlin. Geometric inference for general high-dimensional linear inverse problems. Annals of Statistics, 44(4):1536–1563, August 2016. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1467894707.

Cai:2017:CSB

[CLR17] T. Tony Cai, Tengyuan Liang, and Alexander Rakhlin. Computational and statistical boundaries for submatrix localization in a large noisy matrix. Annals of Statistics, 45 (4):1403–1430, August 2017. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1498636861.

Chan:2018:EMD

[CLSY18] Kwun Chuen Gary Chan, Hok Kan Ling, Tony Sit, and Sheung Chi Phillip Yam. Estimation of a monotone density in s-sample biased sampling models. Annals of Statistics, 46(5):2125–2152, October 2018. CODEN ASTSC7. ISSN

0090-5364. URL http://projecteuclid.org/euclid.aos/
1534492831.

Cai:2013:ACI

[CLX13] T. Tony Cai, Mark G. Low, and Yin Xia. Adaptive confidence intervals for regression functions under shape constraints. Annals of Statistics, 41(2):722–750, April 2013. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1368018171.

Chen:2018:CMM

[CLX18] Yudong Chen, Xiaodong Li, and Jiaming Xu. Convexified modularity maximization for degree-corrected stochastic block models. Annals of Statistics, 46(4):1573–1602, August 2018. CODEN ASTSC7. ISSN 0090-5364. URL http://

projecteuclid.org/euclid.aos/1530086426.

Cai:2016:ESP

[CLZ16] T. Tony Cai, Weidong Liu, and Harrison H. Zhou. Estimating sparse precision matrix: Optimal rates of convergence and adaptive estimation. Annals of Statistics, 44(2):455– 488, April 2016. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1458245724.

Chen:2019:TSA

[CLZ19a] Song Xi Chen, Jun Li, and Ping-Shou Zhong. Two-sample and ANOVA tests for high dimensional means. Annals of Statistics, 47(3):1443–1474, June 2019. CODEN ASTSC7.

ISSN 0090-5364. URL http://projecteuclid.org/euclid.
aos/1550026845.

Chen:2019:QRU

[CLZ19b] Xi Chen, Weidong Liu, and Yichen Zhang. Quantile regression under memory constraint. Annals of Statistics, 47 (6):3244–3273, December 2019. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://

projecteuclid.org/euclid.aos/1572487392.

Chen:2012:NMR

[CM12] Dong Chen and Hans-Georg Müller. Nonlinear manifold representations for functional data. Annals of Statistics, 40(1):1–29, February 2012. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1331830772.

Colombo:2012:LHD

[CMKR12] Diego Colombo, Marloes H. Maathuis, Markus Kalisch, and Thomas S. Richardson. Learning high-dimensional directed acyclic graphs with latent and selection variables. Annals of Statistics, 40(1):294–321, February 2012. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1333567191.

Cai:2013:SPO

[CMW13] T. Tony Cai, Zongming Ma, and Yihong Wu. Sparse PCA: Optimal rates and adaptive estimation. Annals of Statistics, 41(6):3074–3110, December 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1388545679.

Chen:2014:MJP

[CMW14] Xian Chen, Zhi-Ming Ma, and Ying Wang. Markov jump processes in modeling coalescent with recombination. Annals of Statistics, 42(4):1361–1393, August 2014. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1403715204.

Cai:2019:CCH

[CMZ19] T. Tony Cai, Jing Ma, and Linjun Zhang. CHIME: Clustering of high-dimensional Gaussian mixtures with EM algorithm and its optimality. Annals of Statistics, 47(3):1234– 1267, June 2019. CODEN ASTSC7. ISSN 0090-5364. URL

http://projecteuclid.org/euclid.aos/1550026835.

Castillo:2013:NBM

[CN13] Ismaël Castillo and Richard Nickl. Nonparametric Bernstein– von Mises theorems in Gaussian white noise. Annals of Statistics, 41(4):1999–2028, August 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1382547511.

Castillo:2014:BMP

[CN14] Ismaël Castillo and Richard Nickl. On the Bernstein–von Mises phenomenon for nonparametric Bayes procedures. Annals of Statistics, 42(5):1941–1969, October 2014. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1410440630.

Cutting:2017:TUH

[CPV17] Christine Cutting, Davy Paindaveine, and Thomas Verdebout. Testing uniformity on high-dimensional spheres against monotone rotationally symmetric alternatives. Annals of Statistics, 45(3):1024–1058, June 2017. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1497319687.

Chandrasekaran:2012:LVG

[CPW12a] Venkat Chandrasekaran, Pablo A. Parrilo, and Alan S. Willsky. Latent variable graphical model selection via convex optimization. Annals of Statistics, 40(4):1935–1967, August 2012. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1351602527.

Chandrasekaran:2012:RLV

[CPW12b] Venkat Chandrasekaran, Pablo A. Parrilo, and Alan S. Willsky. Rejoinder: Latent variable graphical model selection via convex optimization. Annals of Statistics, 40(4):2005–2013, August 2012. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1351602534.

Chen:2010:TST

[CQ10] Song Xi Chen and Ying-Li Qin. A two-sample test for high-dimensional data with applications to gene-set testing. Annals of Statistics, 38(2):808–835, April 2010. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1266586615.

Chung:2013:EAR

[CR13] EunYi Chung and Joseph P. Romano. Exact and asymptotically robust permutation tests. Annals of Statistics, 41(2):484–507, April 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1366138199.

Castillo:2015:BMT

[CR15] Ismaël Castillo and Judith Rousseau. A Bernstein–von Mises theorem for smooth functionals in semiparametric models. Annals of Statistics, 43(6):2353–2383, December 2015. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1444222078.

Candes:2012:DLV

[CS12] Emmanuel J. Candés and Mahdi Soltanolkotabi. Discussion: Latent variable graphical model selection via convex optimization. Annals of Statistics, 40(4):1997–2004, August 2012. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1351602533.

Cheng:2015:JAS

[CS15] Guang Cheng and Zuofeng Shang. Joint asymptotics for semi-nonparametric regression models with partially linear structure. Annals of Statistics, 43(3):1351–1390, June 2015. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1431695647.

Cheng:2017:MTL

[CS17] Dan Cheng and Armin Schwartzman. Multiple testing of local maxima for detection of peaks in random fields. Annals of Statistics, 45(2):529–556, April 2017. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1494921949.

Castillo:2015:BLR

[CSHvdV15] Ismaël Castillo, Johannes Schmidt-Hieber, and Aad van der Vaart. Bayesian linear regression with sparse priors. Annals of Statistics, 43(5):1986–2018, October 2015. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1438606851.

Chen:2016:SNC

[CSWX16] Xiaohong Chen, Qi-Man Shao, Wei Biao Wu, and Lihu Xu. Self-normalized Cramér-type moderate deviations under dependence. Annals of Statistics, 44(4):1593–1617, August 2016. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1467894709.

Cape:2019:TIN

[CTP19] Joshua Cape, Minh Tang, and Carey E. Priebe. The twoto-infinity norm and singular subspace geometry with applications to high-dimensional statistics. Annals of Statistics, 47(5):2405–2439, October 2019. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://

projecteuclid.org/euclid.aos/1564797852.

Choi:2017:SNP

[CTT17] Yunjin Choi, Jonathan Taylor, and Robert Tibshirani. Selecting the number of principal components: Estimation of the true rank of a noisy matrix. Annals of Statistics, 45(6): 2590–2617, December 2017. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

Chang:2013:MEL

[CTW13] Jinyuan Chang, Cheng Yong Tang, and Yichao Wu. Marginal empirical likelihood and sure independence feature screening. Annals of Statistics, 41(4):2123–2148, August 2013. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1382547515.

Chang:2016:LIF

[CTW16] Jinyuan Chang, Cheng Yong Tang, and Yichao Wu. Local independence feature screening for nonparametric and semiparametric models by marginal empirical likelihood. Annals of Statistics, 44(2):515–539, April 2016. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1458245726.

Chang:2018:NSP

[CTW18] Jinyuan Chang, Cheng Yong Tang, and Tong Tong Wu. A new scope of penalized empirical likelihood with highdimensional estimating equations. Annals of Statistics, 46 (6B):3185–3216, December 2018. CODEN ASTSC7. ISSN

0090-5364. URL http://projecteuclid.org/euclid.aos/
1536631271.

Carpentier:2019:AES

[CV19a] Alexandra Carpentier and Nicolas Verzelen. Adaptive estimation of the sparsity in the Gaussian vector model. Annals of Statistics, 47(1):93–126, February 2019. CODEN ASTSC7.

ISSN 0090-5364. URL http://projecteuclid.org/euclid.
aos/1543568583.

Chakraborty:2019:SSM

[CV19b] Rudrasis Chakraborty and Baba C. Vemuri. Statistics on the Stiefel manifold: Theory and applications. Annals of Statistics, 47(1):415–438, February 2019. CODEN ASTSC7.

ISSN 0090-5364. URL http://projecteuclid.org/euclid.
aos/1543568593.

Castillo:2012:NSH

[CvdV12] Ismaël Castillo and Aad van der Vaart. Needles and straw in a haystack: Posterior concentration for possibly sparse sequences. Annals of Statistics, 40(4):2069–2101, August 2012. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1351602537.

Chetelat:2012:IMN

[CW12] Didier Chételat and Martin T. Wells. Improved multivariate normal mean estimation with unknown covariance when p is greater than n. Annals of Statistics, 40(6):3137–3160, December 2012. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1361542077.

Choi:2014:CCS

[CW14] David Choi and Patrick J. Wolfe. Co-clustering separately exchangeable network data. Annals of Statistics, 42(1):29–63, February 2014. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1389795744.

Chan:2015:ODM

[CW15] Hock Peng Chan and Guenther Walther. Optimal detection of multi-sample aligned sparse signals. Annals of Statistics, 43(5):1865–1895, October 2015. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1438606847.

Chetelat:2019:MSA

[CW19] Didier Chételat and Martin T. Wells. The middle-scale asymptotics of Wishart matrices. Annals of Statistics, 47 (5):2639–2670, October 2019. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://

projecteuclid.org/euclid.aos/1564797859.

Chen:2013:CPM

[CXW13] Xiaohui Chen, Mengyu Xu, and Wei Biao Wu. Covariance and precision matrix estimation for high-dimensional time series. Annals of Statistics, 41(6):2994–3021, December 2013. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1388545676.

Cai:2011:OEM

[CY11] T. Tony Cai and Ming Yuan. Optimal estimation of the mean function based on discretely sampled functional data: Phase transition. Annals of Statistics, 39(5):2330–2355, October 2011. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1322663460.

Cai:2012:ACM

[CY12] T. Tony Cai and Ming Yuan. Adaptive covariance matrix estimation through block thresholding. Annals of Statistics, 40(4):2014–2042, August 2012. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1351602535.

Cai:2014:DST

[CY14] T. Tony Cai and Ming Yuan. Discussion: “A significance test for the lasso”. Annals of Statistics, 42(2):478–482, April 2014. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1400592163. See [LTTT14c].

Chen:2010:ALD

[CZ10] Song Xi Chen and Ping-Shou Zhong. ANOVA for longitudinal data with missing values. Annals of Statistics, 38 (6):3630–3659, December 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1291126968.

Cai:2012:ORC

[CZ12] T. Tony Cai and Harrison H. Zhou. Optimal rates of convergence for sparse covariance matrix estimation. Annals of Statistics, 40(5):2389–2420, October 2012. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1359987525.

Cai:2015:RMR

[CZ15a] T. Tony Cai and Anru Zhang. ROP: Matrix recovery via rank-one projections. Annals of Statistics, 43(1):102–138, February 2015. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1416322038.

Chen:2015:GBC

[CZ15b] Hao Chen and Nancy Zhang. Graph-based change-point detection. Annals of Statistics, 43(1):139–176, February 2015. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1416322039.

Cai:2016:DIF

[CZ16] T. Tony Cai and Linjun Zhang. Discussion of “Influential features PCA for high dimensional clustering”. Annals of Statistics, 44(6):2372–2381, December 2016. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/

1479891620. See [JW16a, JW16b].

Cai:2018:ROP

[CZ18] T. Tony Cai and Anru Zhang. Rate-optimal perturbation bounds for singular subspaces with applications to highdimensional statistics. Annals of Statistics, 46(1):60–89, February 2018. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

Chen:2010:CIS

[CZC10] Xin Chen, Changliang Zou, and R. Dennis Cook. Coordinateindependent sparse sufficient dimension reduction and variable selection. Annals of Statistics, 38(6):3696–3723, December 2010. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1291126970.

Chambaz:2017:TSD

[CZvdL17] Antoine Chambaz, Wenjing Zheng, and Mark J. van der Laan. Targeted sequential design for targeted learning inference of the optimal treatment rule and its mean reward. Annals of Statistics, 45(6):2537–2564, December 2017. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

Chu:2011:PML

[CZW11] Tingjin Chu, Jun Zhu, and Haonan Wang. Penalized maximum likelihood estimation and variable selection in geostatistics. Annals of Statistics, 39(5):2607–2625, October 2011. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1324563349.

Cai:2010:ORCb

[CZZ10] T. Tony Cai, Cun-Hui Zhang, and Harrison H. Zhou. Optimal rates of convergence for covariance matrix estimation. Annals of Statistics, 38(4):2118–2144, August 2010. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1278861244.

Dieuleveut:2016:NSA

[DB16] Aymeric Dieuleveut and Francis Bach. Nonparametric stochastic approximation with large step-sizes. Annals of Statistics, 44(4):1363–1399, August 2016. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1467894702.

Deligiannidis:2019:EEB

[DBCD19] George Deligiannidis, Alexandre Bouchard-Côté, and Arnaud Doucet. Exponential ergodicity of the bouncy particle sampler. Annals of Statistics, 47(3):1268–1287, June 2019. CODEN ASTSC7. ISSN 0090-5364. URL http://

projecteuclid.org/euclid.aos/1550026836.

Dicker:2017:FRQ

[DE17] Lee H. Dicker and Murat A. Erdogdu. Flexible results for quadratic forms with applications to variance components estimation. Annals of Statistics, 45(1):386–414, February 2017. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1487667627.

Delaigle:2016:PHM

[Del16] Aurore Delaigle. Peter Hall’s main contributions to deconvolution. Annals of Statistics, 44(5):1854–1866, October 2016. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1473685260.

Desgagne:2015:ROL

[Des15] Alain Desgagné. Robustness to outliers in location-scale parameter model using log-regularly varying distributions. Annals of Statistics, 43(4):1568–1595, August 2015. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1434546215.

Drton:2019:MLT

[DFKP19] Mathias Drton, Christopher Fox, Andreas Käufl, and Guillaume Pouliot. The maximum likelihood threshold of a path diagram. Annals of Statistics, 47(3):1536–1553, June 2019. CODEN ASTSC7. ISSN 0090-5364. URL http://

projecteuclid.org/euclid.aos/1550026848.

Drton:2011:GIL

[DFS11] Mathias Drton, Rina Foygel, and Seth Sullivant. Global identifiability of linear structural equation models. Annals of Statistics, 39(2):865–886, April 2011. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1299680957.

Drton:2019:CML

[DFW19] Mathias Drton, Christopher Fox, and Y. Samuel Wang. Computation of maximum likelihood estimates in cyclic structural equation models. Annals of Statistics, 47(2):663–690, April 2019. CODEN ASTSC7. ISSN 0090-5364. URL http://

projecteuclid.org/euclid.aos/1547197234.

Drton:2012:CMM

[DG12] Mathias Drton and Aldo Goia. Correction on moments of minors of Wishart matrices. Annals of Statistics, 40(2):1283– 1284, April 2012. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1342625469. See [DMO08].

Dette:2014:ODS

[DG14a] Holger Dette and Yuri Grigoriev. E-optimal designs for second-order response surface models. Annals of Statistics, 42(4):1635–1656, August 2014. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1407420011.

Donoho:2014:MRM

[DG14b] David Donoho and Matan Gavish. Minimax risk of matrix denoising by singular value thresholding. Annals of Statistics, 42(6):2413–2440, December 2014. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1413810732.

DeCastro:2019:AOD

[DGH^{+}19] Yohann De Castro, Fabrice Gamboa, Didier Henrion, Roxana Hess, and Jean-Bernard Lasserre. Approximate optimal designs for multivariate polynomial regression. Annals of Statistics, 47(1):127–155, February 2019. CODEN ASTSC7. ISSN

0090-5364. URL http://projecteuclid.org/euclid.aos/
1543568584.

Dattner:2011:DDF

[DGJ11] I. Dattner, A. Goldenshluger, and A. Juditsky. On deconvolution of distribution functions. Annals of Statistics, 39(5):2477–2501, October 2011. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1322663465.

Donoho:2018:OSE

[DGJ18] David Donoho, Matan Gavish, and Iain Johnstone. Optimal shrinkage of eigenvalues in the spiked covariance model. Annals of Statistics, 46(4):1742–1778, August 2018. CODEN ASTSC7. ISSN 0090-5364. URL http://projecteuclid.

org/euclid.aos/1530086432.

Dasgupta:2019:FEK

[DGK19] Sayan Dasgupta, Yair Goldberg, and Michael R. Kosorok. Feature elimination in kernel machines in moderately high dimensions. Annals of Statistics, 47(1):497–526, February 2019. CODEN ASTSC7. ISSN 0090-5364. URL http://

projecteuclid.org/euclid.aos/1543568596.

Das:2019:PBA

[DGL19] Debraj Das, Karl Gregory, and S. N. Lahiri. Perturbation bootstrap in adaptive lasso. Annals of Statistics, 47 (4):2080–2116, August 2019. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://

projecteuclid.org/euclid.aos/1558425640.

Dalalyan:2018:EWA

[DGP18] Arnak S. Dalalyan, Edwin Grappin, and Quentin Paris. On the exponentially weighted aggregate with the Laplace prior. Annals of Statistics, 46(5):2452–2478, October 2018. CODEN ASTSC7. ISSN 0090-5364. URL http://projecteuclid.

org/euclid.aos/1534492841.

Dong:2016:ESI

[DGT16] Chaohua Dong, Jiti Gao, and Dag Tjøstheim. Estimation for single-index and partially linear single-index integrated models. Annals of Statistics, 44(1):425–453, February 2016. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1452004792.

Delaigle:2010:DPD

[DH10] Aurore Delaigle and Peter Hall. Defining probability density for a distribution of random functions. Annals of Statistics, 38(2):1171–1193, April 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1266586626.

Delaigle:2012:MTP

[DH12a] Aurore Delaigle and Peter Hall. Methodology and theory for partial least squares applied to functional data. Annals of Statistics, 40(1):322–352, February 2012. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1333567192.

Delaigle:2012:NRH

[DH12b] Aurore Delaigle and Peter Hall. Nonparametric regression with homogeneous group testing data. Annals of Statistics, 40(1):131–158, February 2012. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1331830777.

Dinh:2018:CCR

[DHSM18] Vu Dinh, Lam Si Tung Ho, Marc A. Suchard, and Frederick A. Matsen IV. Consistency and convergence rate of phylogenetic inference via regularization. Annals of Statistics, 46(4):1481–1512, August 2018. CODEN ASTSC7. ISSN

0090-5364. URL http://projecteuclid.org/euclid.aos/
1530086423.

Delaigle:2016:NCA

[DHZ16] Aurore Delaigle, Peter Hall, and Wen-Xin Zhou. Nonparametric covariate-adjusted regression. Annals of Statistics, 44(5):2190–2220, October 2016. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1473685273.

Dick:2011:HOS

[Dic11] Josef Dick. Higher order scrambled digital nets achieve the optimal rate of the root mean square error for smooth integrands. Annals of Statistics, 39(3):1372–1398, June 2011. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1304514657.

deJonge:2010:ANB

[dJvZ10] R. de Jonge and J. H. van Zanten. Adaptive nonparametric Bayesian inference using location-scale mixture priors. Annals of Statistics, 38(6):3300–3320, December 2010. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1284988407.

Durot:2012:LDE

[DKL12] Cécile Durot, Vladimir N. Kulikov, and Hendrik P. Lopuhaä. The limit distribution of the L_{∞}-error of Grenandertype estimators. Annals of Statistics, 40(3):1578–1608, June 2012. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1346850066.

Duchi:2018:MCI

[DKR18] John Duchi, Khashayar Khosravi, and Feng Ruan. Multiclass classification, information, divergence and surrogate risk. Annals of Statistics, 46(6B):3246–3275, December 2018. CODEN ASTSC7. ISSN 0090-5364. URL http://

projecteuclid.org/euclid.aos/1536631273.

Draisma:2013:GAG

[DKZ13] Jan Draisma, Sonja Kuhnt, and Piotr Zwiernik. Groups acting on Gaussian graphical models. Annals of Statistics, 41(4):1944–1969, August 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1382547509.

Dette:2017:NAO

[DKZ17] Holger Dette, Maria Konstantinou, and Anatoly Zhigljavsky. A new approach to optimal designs for correlated observations. Annals of Statistics, 45(4):1579–1608, August 2017. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1498636867.

Dereudre:2017:CLE

[DL17] David Dereudre and Frédéric Lavancier. Consistency of likelihood estimation for Gibbs point processes. Annals of Statistics, 45(2):744–770, April 2017. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1494921956.

Devroye:2016:SGM

[DLLO16] Luc Devroye, Matthieu Lerasle, Gabor Lugosi, and Roberto I. Oliveira. Sub-Gaussian mean estimators. Annals of Statistics, 44(6):2695–2725, December 2016. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1479891632.

Dawid:2012:PLS

[DLP12] A. Philip Dawid, Steffen Lauritzen, and Matthew Parry. Proper local scoring rules on discrete sample spaces. Annals of Statistics, 40(1):593–608, February 2012. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1336396184.

Dette:2011:NGP

[DM11] Holger Dette and Viatcheslav B. Melas. A note on the de la Garza phenomenon for locally optimal designs. Annals of Statistics, 39(2):1266–1281, April 2011. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1304947050.

Douc:2012:APM

[DM12] Randal Douc and Eric Moulines. Asymptotic properties of the maximum likelihood estimation in misspecified hidden Markov models. Annals of Statistics, 40(5):2697–2732, October 2012. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1359987535.

Dai:2018:PCA

[DM18] Xiongtao Dai and Hans-Georg Müller. Principal component analysis for functional data on Riemannian manifolds and spheres. Annals of Statistics, 46(6B):3334–3361, December 2018. CODEN ASTSC7. ISSN 0090-5364. URL http://

projecteuclid.org/euclid.aos/1536631276.

Dette:2015:BOD

[DMG15] Holger Dette, Viatcheslav B. Melas, and Roman Guchenko. Bayesian T -optimal discriminating designs. Annals of Statistics, 43(5):1959–1985, October 2015. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1438606850.

Duchi:2013:ARA

[DMJ13] John C. Duchi, Lester Mackey, and Michael I. Jordan. The asymptotics of ranking algorithms. Annals of Statistics, 41(5):2292–2323, October 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1383661265.

Du:2010:PVS

[DML10] Pang Du, Shuangge Ma, and Hua Liang. Penalized variable selection procedure for Cox models with semiparametric relative risk. Annals of Statistics, 38(4):2092–2117, August 2010. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1278861243.

Drton:2008:MMW

[DMO08] Mathias Drton, Hélène Massam, and Ingram Olkin. Moments of minors of Wishart matrices. Annals of Statistics, 36(5):2261–2283, October 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1223908092. See correc-

tion [DG12].

Douc:2011:CML

[DMOvH11] Randal Douc, Eric Moulines, Jimmy Olsson, and Ramon van Handel. Consistency of the maximum likelihood estimator for general hidden Markov models. Annals of Statistics, 39(1):474–513, February 2011. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1297779854.

Dette:2012:ODD

[DMS12] Holger Dette, Viatcheslav B. Melas, and Petr Shpilev. T - optimal designs for discrimination between two polynomial models. Annals of Statistics, 40(1):188–205, February 2012. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1331830779.

Dette:2013:ROD

[DMS13a] Holger Dette, Viatcheslav B. Melas, and Petr Shpilev. Robust T -optimal discriminating designs. Annals of Statistics, 41(4):1693–1715, August 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1378386236.

DiNardo:2013:NSS

[DMS13b] E. Di Nardo, P. McCullagh, and D. Senato. Natural statistics for spectral samples. Annals of Statistics, 41(2):982– 1004, April 2013. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1369836967.

Ding:2011:SMT

[DN11] Ying Ding and Bin Nan. A sieve M-theorem for bundled parameters in semiparametric models, with application to the efficient estimation in a linear model for censored data. Annals of Statistics, 39(6):3032–3061, December 2011. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1327413777.

Dobriban:2017:SDP

[Dob17] Edgar Dobriban. Sharp detection in PCA under correlations: All eigenvalues matter. Annals of Statistics, 45(4):1810–1833, August 2017. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1498636875.

Davidov:2013:LSO

[DP13] Ori Davidov and Shyamal Peddada. The linear stochastic order and directed inference for multivariate ordered distributions. Annals of Statistics, 41(1):1–40, February 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1362493038.

Doss:2018:MAE

[DP18] Hani Doss and Yeonhee Park. An MCMC approach to empirical Bayes inference and Bayesian sensitivity analysis via empirical processes. Annals of Statistics, 46(4):1630–1663, August 2018. CODEN ASTSC7. ISSN 0090-5364. URL

http://projecteuclid.org/euclid.aos/1530086428.

Descary:2019:FDA

[DP19] Marie-Hélène Descary and Victor M. Panaretos. Functional data analysis by matrix completion. Annals of Statistics, 47(1):1–38, February 2019. CODEN ASTSC7. ISSN

0090-5364. URL http://projecteuclid.org/euclid.aos/
1543568580.

Dou:2012:EFR

[DPZ12] Winston Wei Dou, David Pollard, and Harrison H. Zhou. Estimation in functional regression for general exponential families. Annals of Statistics, 40(5):2421–2451, October 2012. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1359987526.

Dette:2013:ODL

[DPZ13] Holger Dette, Andrey Pepelyshev, and Anatoly Zhigljavsky. Optimal design for linear models with correlated observations. Annals of Statistics, 41(1):143–176, February 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1362493043.

Dette:2016:ODR

[DPZ16] Holger Dette, Andrey Pepelyshev, and Anatoly Zhigljavsky. Optimal designs in regression with correlated errors. Annals of Statistics, 44(1):113–152, February 2016. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1449755959.

Dette:2019:BCT

[DPZ19] Holger Dette, Andrey Pepelyshev, and Anatoly Zhigljavsky. The BLUE in continuous-time regression models with correlated errors. Annals of Statistics, 47(4):1928–1959, August 2019. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/euclid.
aos/1558425635.

Drees:2010:LTE

[DR10] Holger Drees and Holger Rootzén. Limit theorems for empirical processes of cluster functionals. Annals of Statistics, 38(4):2145–2186, August 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1278861245. See correc-

tion [DR16].

Douc:2011:VRB

[DR11] Randal Douc and Christian P. Robert. A vanilla Rao– Blackwellization of Metropolis–Hastings algorithms. Annals of Statistics, 39(1):261–277, February 2011. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1291388375.

Delattre:2015:NPC

[DR15] Sylvain Delattre and Etienne Roquain. New procedures controlling the false discovery proportion via Romano–Wolf’s heuristic. Annals of Statistics, 43(3):1141–1177, June 2015. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1431695641.

Drees:2016:C

[DR16] Holger Drees and Holger Rootzén. Correction. Annals of Statistics, 44(3):1360–1361, June 2016. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1460381696. See

[DR10].

Dai:2012:DOL

[DRZ12] Dong Dai, Philippe Rigollet, and Tong Zhang. Deviation optimal learning using greedy Q-aggregation. Annals of Statistics, 40(3):1878–1905, June 2012. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1350394520.

Davis:2011:URM

[DS11] Richard A. Davis and Li Song. Unit roots in moving averages beyond first order. Annals of Statistics, 39(6):3062–3091, December 2011. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1327413778.

Dalalyan:2012:SOI

[DS12] Arnak S. Dalalyan and Joseph Salmon. Sharp oracle inequalities for aggregation of affine estimators. Annals of Statistics, 40(4):2327–2355, August 2012. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1358951384.

Dette:2013:CCD

[DS13] Holger Dette and Kirsten Schorning. Complete classes of designs for nonlinear regression models and principal representations of moment spaces. Annals of Statistics, 41(3):1260– 1267, June 2013. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1371150900.

Dette:2016:ODC

[DS16] Holger Dette and Kirsten Schorning. Optimal designs for comparing curves. Annals of Statistics, 44(3):1103–1130, June 2016. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1460381688.

Dumbgen:2011:ALC

[DSS11] Lutz Dümbgen, Richard Samworth, and Dominic Schuhmacher. Approximation by log-concave distributions, with applications to regression. Annals of Statistics, 39(2):702– 730, April 2011. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1299680952.

Dobriban:2018:HDA

[DW18] Edgar Dobriban and Stefan Wager. High-dimensional asymptotics of prediction: Ridge regression and classification. Annals of Statistics, 46(1):247–279, February 2018. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

Dette:2019:DRC

[DW19a] Holger Dette and Weichi Wu. Detecting relevant changes in the mean of nonstationary processes — a mass excess approach. Annals of Statistics, 47(6):3578–3608, December 2019. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/euclid.
aos/1572487403.

Doss:2019:IML

[DW19b] Charles R. Doss and Jon A. Wellner. Inference for the mode of a log-concave density. Annals of Statistics, 47(5):2950–2976, October 2019. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL https://projecteuclid.org/
euclid.aos/1564797869.

Du:2014:SIM

[DZ14] Lilun Du and Chunming Zhang. Single-index modulated multiple testing. Annals of Statistics, 42(4):1262–1311, August 2014. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1403715201.

Datta:2017:CHD

[DZ17] Abhirup Datta and Hui Zou. CoCoLasso for high-dimensional error-in-variables regression. Annals of Statistics, 45(6):2400– 2426, December 2017. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

Escobar-Bach:2018:LRE

[EBGG18] Mikael Escobar-Bach, Yuri Goegebeur, and Armelle Guillou. Local robust estimation of the Pickands dependence function. Annals of Statistics, 46(6A):2806–2843, December 2018. CODEN ASTSC7. ISSN 0090-5364. URL http://

projecteuclid.org/euclid.aos/1536307234.

Efromovich:2013:NRS

[Efr13] Sam Efromovich. Nonparametric regression with the scale depending on auxiliary variable. Annals of Statistics, 41(3):1542–1568, June 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1375362559.

Efron:2018:CIM

[Efr18] Bradley Efron. Curvature and inference for maximum likelihood estimates. Annals of Statistics, 46(4):1664–1692, August 2018. CODEN ASTSC7. ISSN 0090-5364. URL http:

//projecteuclid.org/euclid.aos/1530086429.

Ehm:2012:LPS

[EG12] Werner Ehm and Tilmann Gneiting. Local proper scoring rules of order two. Annals of Statistics, 40(1):609–637, February 2012. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1336396185.

Eltzner:2019:SCL

[EH19a] Benjamin Eltzner and Stephan F. Huckemann. A smeary central limit theorem for manifolds with application to highdimensional spheres. Annals of Statistics, 47(6):3360–3381, December 2019. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL https://projecteuclid.org/
euclid.aos/1572487396.

Enikeeva:2019:HDC

[EH19b] Farida Enikeeva and Zaid Harchaoui. High-dimensional change-point detection under sparse alternatives. Annals of Statistics, 47(4):2051–2079, August 2019. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https:

//projecteuclid.org/euclid.aos/1558425639.

Ehrlinger:2012:CB

[EI12] John Ehrlinger and Hemant Ishwaran. Characterizing L_{2}Boosting. Annals of Statistics, 40(2):1074–1101, April 2012. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.
aos/1342625462.

Evans:2010:IPV

[EJ10] Michael Evans and Gun Ho Jang. Invariant P -values for model checking. Annals of Statistics, 38(1):512–525, February 2010. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1262271622.

Elenberg:2018:RSC

[EKDN18] Ethan R. Elenberg, Rajiv Khanna, Alexandros G. Dimakis, and Sahand Negahban. Restricted strong convexity implies weak submodularity. Annals of Statistics, 46(6B):3539–3568, December 2018. CODEN ASTSC7. ISSN 0090-5364. URL

http://projecteuclid.org/euclid.aos/1536631283.

Einmahl:2012:MET

[EKS12] John H. J. Einmahl, Andrea Krajina, and Johan Segers. An M-estimator for tail dependence in arbitrary dimensions. Annals of Statistics, 40(3):1764–1793, June 2012. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1349196391.

ElKaroui:2010:HDE

[El 10a] Noureddine El Karoui. High-dimensionality effects in the Markowitz problem and other quadratic programs with linear constraints: Risk underestimation. Annals of Statistics, 38(6):3487–3566, December 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1291126965.

ElKaroui:2010:IPN

[El 10b] Noureddine El Karoui. On information plus noise kernel random matrices. Annals of Statistics, 38(5):3191–3216, October 2010. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1284391762.

Karoui:2010:SKR

[El 10c] Noureddine El Karoui. The spectrum of kernel random matrices. Annals of Statistics, 38(1):1–50, February 2010. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1262271608.

Egloff:2010:QEA

[EL10] Daniel Egloff and Markus Leippold. Quantile estimation with adaptive importance sampling. Annals of Statistics, 38(2):1244–1278, April 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1266586629.

Einmahl:2015:BCE

[ELL15] John H. J. Einmahl, Jun Li, and Regina Y. Liu. Bridging centrality and extremity: Refining empirical data depth using extreme value statistics. Annals of Statistics, 43(6):2738–2765, December 2015. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1444222091.

Escanciano:2018:ADF

[EPFV18] Juan Carlos Escanciano, Juan Carlos Pardo-Fernández, and Ingrid Van Keilegom. Asymptotic distribution-free tests for semiparametric regressions with dependent data. Annals of Statistics, 46(3):1167–1196, June 2018. CODEN ASTSC7.

ISSN 0090-5364. URL http://projecteuclid.org/euclid.
aos/1525313079.

Evans:2014:MAD

[ER14] Robin J. Evans and Thomas S. Richardson. Markovian acyclic directed mixed graphs for discrete data. Annals of Statistics, 42(4):1452–1482, August 2014. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1407420005.

Ernst:2017:YNC

[ESW17] Philip A. Ernst, Larry A. Shepp, and Abraham J. Wyner. Yule’s “nonsense correlation” solved! Annals of Statistics, 45(4):1789–1809, August 2017. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1498636874.

Evans:2018:MDB

[Eva18] Robin J. Evans. Margins of discrete Bayesian networks. Annals of Statistics, 46(6A):2623–2656, December 2018. CODEN ASTSC7. ISSN 0090-5364. URL http://

projecteuclid.org/euclid.aos/1536307228.

Elsener:2018:RLR

[EvdG18] Andreas Elsener and Sara van de Geer. Robust low-rank matrix estimation. Annals of Statistics, 46(6B):3481–3509, December 2018. CODEN ASTSC7. ISSN 0090-5364. URL

http://projecteuclid.org/euclid.aos/1536631281.

ElKaroui:2016:GCL

[EW16] Noureddine El Karoui and Hau-Tieng Wu. Graph connection Laplacian methods can be made robust to noise. Annals of Statistics, 44(1):346–372, February 2016. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1452004789.

Foygel:2012:HTC

[FDD12] Rina Foygel, Jan Draisma, and Mathias Drton. Half-trek criterion for generic identifiability of linear structural equation models. Annals of Statistics, 40(3):1682–1713, June 2012. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1349196388.

Ferreira:2015:BMM

[FdH15] Ana Ferreira and Laurens de Haan. On the block maxima method in extreme value theory: PWM estimators. Annals of Statistics, 43(1):276–298, February 2015. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1418135622.

Fougeres:2015:BCM

[FdHM15] Anne-Laure Fougères, Laurens de Haan, and Cécile Mercadier. Bias correction in multivariate extremes. Annals of Statistics, 43(2):903–934, April 2015. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1427115291.

Fellouris:2012:AOP

[Fel12] Georgios Fellouris. Asymptotically optimal parameter estimation under communication constraints. Annals of Statistics, 40(4):2239–2265, August 2012. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1358951381.

Fan:2014:ARV

[FFB14] Jianqing Fan, Yingying Fan, and Emre Barut. Adaptive robust variable selection. Annals of Statistics, 42(1):324–351, February 2014. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1395234980.

Fan:2010:NEG

[FFN10] Jianqing Fan, Yang Feng, and Yue S. Niu. Nonparametric estimation of genewise variance for microarray data. Annals of Statistics, 38(5):2723–2750, October 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1278861458.

Fan:2018:APE

[FG18a] Zhou Fan and Leying Guan. Approximate ℓ_{0}-penalized estimation of piecewise-constant signals on graphs. Annals of Statistics, 46(6B):3217–3245, December 2018. CODEN ASTSC7. ISSN 0090-5364. URL http://projecteuclid.

org/euclid.aos/1536631272.

Finner:2018:TSK

[FG18b] Helmut Finner and Veronika Gontscharuk. Two-sample Kolmogorov–Smirnov-type tests revisited: Old and new tests in terms of local levels. Annals of Statistics, 46 (6A):3014–3037, December 2018. CODEN ASTSC7. ISSN

0090-5364. URL http://projecteuclid.org/euclid.aos/
1536307241.

Freund:2017:NPB

[FGM17] Robert M. Freund, Paul Grigas, and Rahul Mazumder. A new perspective on boosting in linear regression via subgradient optimization and relatives. Annals of Statistics, 45(6):2328– 2364, December 2017. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

Feng:2014:SIB

[FH14a] Xingdong Feng and Xuming He. Statistical inference based on robust low-rank data matrix approximation. Annals of Statistics, 42(1):190–210, February 2014. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1392733185.

Fithian:2014:LCC

[FH14b] William Fithian and Trevor Hastie. Local case-control sampling: Efficient subsampling in imbalanced data sets. Annals of Statistics, 42(5):1693–1724, October 2014. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1410440622.

Flegal:2010:BMS

[FJ10] James M. Flegal and Galin L. Jones. Batch means and spectral variance estimators in Markov chain Monte Carlo. Annals of Statistics, 38(2):1034–1070, April 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1266586622.

Fan:2019:EDV

[FJ19] Zhou Fan and Iain M. Johnstone. Eigenvalue distributions of variance components estimators in high-dimensional random effects models. Annals of Statistics, 47(5):2855–2886, October 2019. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL https://projecteuclid.org/
euclid.aos/1564797866.

Fan:2015:FAR

[FJR15] Yingying Fan, Gareth M. James, and Peter Radchenko. Functional additive regression. Annals of Statistics, 43 (5):2296–2325, October 2015. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1442364153.

Fan:2013:OCS

[FJY13] Yingying Fan, Jiashun Jin, and Zhigang Yao. Optimal classification in sparse Gaussian graphic model. Annals of Statistics, 41(5):2537–2571, October 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1384871345.

Fan:2014:DST

[FK14] Jianqing Fan and Zheng Tracy Ke. Discussion: “A significance test for the lasso”. Annals of Statistics, 42(2):483– 492, April 2014. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1400592164. See [LTTT14c].

Fan:2015:QSD

[FKLX15] Jianqing Fan, Zheng Tracy Ke, Han Liu, and Lucy Xia. QUADRO: A supervised dimension reduction method via Rayleigh quotient optimization. Annals of Statistics, 43 (4):1498–1534, August 2015. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1434546213.

Fan:2015:IIS

[FKLZ15] Yingying Fan, Yinfei Kong, Daoji Li, and Zemin Zheng. Innovated interaction screening for high-dimensional nonlinear classification. Annals of Statistics, 43(3):1243–1272, June 2015. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1431695643.

Fan:2012:VSL

[FL12] Yingying Fan and Runze Li. Variable selection in linear mixed effects models. Annals of Statistics, 40(4):2043–2068, August 2012. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1351602536.

Fan:2014:EHD

[FL14] Jianqing Fan and Yuan Liao. Endogeneity in high dimensions. Annals of Statistics, 42(3):872–917, June 2014. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1400592646.

Fan:2016:ISE

[FL16] Yingying Fan and Jinchi Lv. Innovated scalable efficient estimation in ultra-large Gaussian graphical models. Annals of Statistics, 44(5):2098–2126, October 2016. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1473685270.

Fan:2011:HDC

[FLM11] Jianqing Fan, Yuan Liao, and Martina Mincheva. Highdimensional covariance matrix estimation in approximate factor models. Annals of Statistics, 39(6):3320–3356, December 2011. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1330958681.

Fasy:2014:CSP

[FLR^{+}14] Brittany Terese Fasy, Fabrizio Lecci, Alessandro Rinaldo, Larry Wasserman, Sivaraman Balakrishnan, and Aarti Singh. Confidence sets for persistence diagrams. Annals of Statistics, 42(6):2301–2339, December 2014. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1413810729.

Fromont:2013:TSP

[FLRB13] Magalie Fromont, Béatrice Laurent, and Patricia Reynaud-Bouret. The two-sample problem for Poisson processes: Adaptive tests with a nonasymptotic wild bootstrap approach. Annals of Statistics, 41(3):1431–1461, June 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1375362555.

Fromont:2016:FWS

[FLRB16] Magalie Fromont, Matthieu Lerasle, and Patricia Reynaud-Bouret. Family-Wise Separation Rates for multiple testing. Annals of Statistics, 44(6):2533–2563, December 2016. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1479891627.

Fallat:2017:TPM

[FLS^{+}17] Shaun Fallat, Steffen Lauritzen, Kayvan Sadeghi, Caroline Uhler, Nanny Wermuth, and Piotr Zwiernik. Total positivity in Markov structures. Annals of Statistics, 45(3):1152– 1184, June 2017. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1497319691.

Fan:2018:LSL

[FLSZ18] Jianqing Fan, Han Liu, Qiang Sun, and Tong Zhang. I-LAMM for sparse learning: Simultaneous control of algorithmic complexity and statistical error. Annals of Statistics, 46(2):814–841, April 2018. CODEN ASTSC7. ISSN

0090-5364. URL http://projecteuclid.org/euclid.aos/
1522742437.

Fan:2016:PPC

[FLW16] Jianqing Fan, Yuan Liao, and Weichen Wang. Projected principal component analysis in factor models. Annals of Statistics, 44(1):219–254, February 2016. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1449755962.

Fan:2018:LCE

[FLW18] Jianqing Fan, Han Liu, and Weichen Wang. Large covariance estimation through elliptical factor models. Annals of Statistics, 46(4):1383–1414, August 2018. CODEN ASTSC7. ISSN

0090-5364. URL http://projecteuclid.org/euclid.aos/
1530086420.

Fort:2011:CAI

[FMP11] G. Fort, E. Moulines, and P. Priouret. Convergence of adaptive and interacting Markov chain Monte Carlo algorithms. Annals of Statistics, 39(6):3262–3289, December 2011. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1330958679.

Fienberg:2012:MLE

[FR12] Stephen E. Fienberg and Alessandro Rinaldo. Maximum likelihood estimation in log-linear models. Annals of Statistics, 40(2):996–1023, April 2012. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1342625459.

Field:2013:REB

[FR13] Chris Field and John Robinson. Relative errors for bootstrap approximations of the serial correlation coefficient. Annals of Statistics, 41(2):1035–1053, April 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1369836969.

Fan:2015:EFS

[FRW15] Jianqing Fan, Philippe Rigollet, and Weichen Wang. Estimation of functionals of sparse covariance matrices. Annals of Statistics, 43(6):2706–2737, December 2015. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1444222090.

Fryzlewicz:2014:WBS

[Fry14] Piotr Fryzlewicz. Wild binary segmentation for multiple change-point detection. Annals of Statistics, 42(6): 2243–2281, December 2014. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1413810727.

Fryzlewicz:2018:TGB

[Fry18] Piotr Fryzlewicz. Tail-greedy bottom-up data decompositions and fast multiple change-point detection. Annals of Statistics, 46(6B):3390–3421, December 2018. CODEN ASTSC7. ISSN

0090-5364. URL http://projecteuclid.org/euclid.aos/
1536631278.

Fan:2010:SIS

[FS10] Jianqing Fan and Rui Song. Sure independence screening in generalized linear models with NP-dimensionality. Annals of Statistics, 38(6):3567–3604, December 2010. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1291126966.

Feller:2017:ODD

[FSD^{+}17] Chrystel Feller, Kirsten Schorning, Holger Dette, Georgina Bermann, and Björn Bornkamp. Optimal designs for dose response curves with common parameters. Annals of Statistics, 45(5):2102–2132, October 2017. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

Fan:2018:DSD

[FSZ18] Jianqing Fan, Qi-Man Shao, and Wen-Xin Zhou. Are discoveries spurious? Distributions of maximum spurious correlations and their applications. Annals of Statistics, 46(3):989– 1017, June 2018. CODEN ASTSC7. ISSN 0090-5364. URL

http://projecteuclid.org/euclid.aos/1525313073.

Fuh:2006:ELE

[Fuh06] Cheng-Der Fuh. Efficient likelihood estimation in state space models. Annals of Statistics, 34(4):2026–2068, August 2006. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1162567642. See comments [Jen10] and reply [Fuh10].

Fuh:2010:RSP

[Fuh10] Cheng-Der Fuh. Reply to “On some problems in the article Efficient Likelihood Estimation in State Space Models” by Cheng-Der Fuh [Ann. Statist. 34 (2006) 2026–2068]. Annals of Statistics, 38(2):1282–1285, April 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:/

/projecteuclid.org/euclid.aos/1266586631. See [Fuh06,

Jen10].

Fan:2019:DEP

[FWWZ19] Jianqing Fan, Dong Wang, Kaizheng Wang, and Ziwei Zhu. Distributed estimation of principal eigenspaces. Annals of Statistics, 47(6):3009–3031, December 2019. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/euclid.aos/
1572487381.

Fan:2014:SOO

[FXZ14] Jianqing Fan, Lingzhou Xue, and Hui Zou. Strong oracle optimality of folded concave penalized estimation. Annals of Statistics, 42(3):819–849, June 2014. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1400592644.

Fan:2015:C

[FXZ15] Jianqing Fan, Lingzhou Xue, and Hui Zou. Correction. Annals of Statistics, 43(2):935, April 2015. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1427115292.

Ferrari:2010:MQL

[FY10] Davide Ferrari and Yuhong Yang. Maximum L_{q}-likelihood estimation. Annals of Statistics, 38(2):753–783, April 2010.

CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1266586613.

Francq:2013:INA

[FZ13] Christian Francq and Jean-Michel Zakoïan. Inference in nonstationary asymmetric GARCH models. Annals of Statistics, 41(4):1970–1998, August 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1382547510.

Fissler:2016:HOE

[FZ16] Tobias Fissler and Johanna F. Ziegel. Higher order elicitability and Osband’s principle. Annals of Statistics, 44 (4):1680–1707, August 2016. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1467894712. See correc-

tion [FZ21].

Feng:2019:SCP

[FZ19] Long Feng and Cun-Hui Zhang. Sorted concave penalized regression. Annals of Statistics, 47(6):3069–3098, December 2019. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/euclid.
aos/1572487384.

Fissler:2021:CNH

[FZ21] Tobias Fissler and Johanna F. Ziegel. Correction note: Higher order elicitability and Osband’s principle. Annals of Statistics, 49(1):614, February 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https:

//projecteuclid.org/journals/annals-of-statistics/
volume-49/issue-1/Correction-note-Higher-order-elicitability-
and-Osbands-principle/10.1214/20-AOS2014.full. See

[FZ16].

Gao:2013:MDN

[Gao13] Fuqing Gao. Moderate deviations for a nonparametric estimator of sample coverage. Annals of Statistics, 41(2):641– 669, April 2013. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1366980560.

Gerber:2019:NAO

[GCW19] Mathieu Gerber, Nicolas Chopin, and Nick Whiteley. Negative association, ordering and convergence of resampling methods. Annals of Statistics, 47(4):2236–2260, August 2019. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/euclid.aos/
1558425644.

Ghoshdastidar:2017:CSH

[GD17] Debarghya Ghoshdastidar and Ambedkar Dukkipati. Consistency of spectral hypergraph partitioning under planted partition model. Annals of Statistics, 45(1):289–315, February 2017. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1487667624.

Groeneboom:2018:CSL

[GH18] Piet Groeneboom and Kim Hendrickx. Current status linear regression. Annals of Statistics, 46(4):1415–1444, August 2018. CODEN ASTSC7. ISSN 0090-5364. URL http://

projecteuclid.org/euclid.aos/1530086421.

Gronneberg:2019:PSP

[GH19] Steffen Grønneberg and Benjamin Holcblat. On partial-sum processes of ARMAX residuals. Annals of Statistics, 47 (6):3216–3243, December 2019. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://

projecteuclid.org/euclid.aos/1572487391.

Ghosal:2015:DFC

[Gho15] Subhashis Ghosal. Discussion of “Frequentist coverage of adaptive nonparametric Bayesian credible sets”. Annals of Statistics, 43(4):1455–1462, August 2015. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/

1434546210. See [SvdVvZ15a].

Guo:2014:FRC

[GHS14] Wenge Guo, Li He, and Sanat K. Sarkar. Further results on controlling the false discovery proportion. Annals of Statistics, 42(3):1070–1101, June 2014. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1400592652.

Groeneboom:2015:NCI

[GJ15] Piet Groeneboom and Geurt Jongbloed. Nonparametric confidence intervals for monotone functions. Annals of Statistics, 43(5):2019–2054, October 2015. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1438606852.

Groeneboom:2010:MSL

[GJW10] Piet Groeneboom, Geurt Jongbloed, and Birgit I. Witte. Maximum smoothed likelihood estimation and smoothed maximum likelihood estimation in the current status model. Annals of Statistics, 38(1):352–387, February 2010. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1262271618.

Goldberg:2012:QLC

[GK12] Yair Goldberg and Michael R. Kosorok. Q-learning with censored data. Annals of Statistics, 40(1):529–560, February 2012. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1336396182.

Gadat:2016:CGF

[GKM16] Sébastien Gadat, Thierry Klein, and Clément Marteau. Classification in general finite dimensional spaces with the knearest neighbor rule. Annals of Statistics, 44(3):982–1009, June 2016. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1460381684.

Goldenshluger:2011:BSK

[GL11] Alexander Goldenshluger and Oleg Lepski. Bandwidth selection in kernel density estimation: Oracle inequalities and adaptive minimax optimality. Annals of Statistics, 39(3):1608–1632, June 2011. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1307452130.

Gehrmann:2012:EMG

[GL12] Helene Gehrmann and Steffen L. Lauritzen. Estimation of means in graphical Gaussian models with symmetries. Annals of Statistics, 40(2):1061–1073, April 2012. CODEN ASTSC7.

ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1342625461.

Gamarnik:2018:FLS

[GL18] David Gamarnik and Quan Li. Finding a large submatrix of a Gaussian random matrix. Annals of Statistics, 46 (6A):2511–2561, December 2018. CODEN ASTSC7. ISSN

0090-5364. URL http://projecteuclid.org/euclid.aos/
1536307224.

Gloter:2018:JFE

[GLM18] Arnaud Gloter, Dasha Loukianova, and Hilmar Mai. Jump filtering and efficient drift estimation for Lévy-driven SDEs. Annals of Statistics, 46(4):1445–1480, August 2018. CODEN ASTSC7. ISSN 0090-5364. URL http://projecteuclid.

org/euclid.aos/1530086422.

Gregory:2018:SBB

[GLN18] Karl B. Gregory, Soumendra N. Lahiri, and Daniel J. Nordman. A smooth block bootstrap for quantile regression with time series. Annals of Statistics, 46(3):1138–1166, June 2018. CODEN ASTSC7. ISSN 0090-5364. URL http://

projecteuclid.org/euclid.aos/1525313078.

Gao:2015:ROG

[GLZ15] Chao Gao, Yu Lu, and Harrison H. Zhou. Rate-optimal graphon estimation. Annals of Statistics, 43(6):2624–2652, December 2015. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1444222087.

Gao:2015:MES

[GMRZ15] Chao Gao, Zongming Ma, Zhao Ren, and Harrison H. Zhou. Minimax estimation in sparse canonical correlation analysis. Annals of Statistics, 43(5):2168–2197, October 2015. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1442364149.

Gao:2017:SCA

[GMZ17] Chao Gao, Zongming Ma, and Harrison H. Zhou. Sparse CCA: Adaptive estimation and computational barriers. Annals of Statistics, 45(5):2074–2101, October 2017. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

Gao:2018:CDD

[GMZZ18] Chao Gao, Zongming Ma, Anderson Y. Zhang, and Harrison H. Zhou. Community detection in degree-corrected block models. Annals of Statistics, 46(5):2153–2185, October 2018. CODEN ASTSC7. ISSN 0090-5364. URL http:/

/projecteuclid.org/euclid.aos/1534492832.

Gine:2010:CBD

[GN10] Evarist Giné and Richard Nickl. Confidence bands in density estimation. Annals of Statistics, 38(2):1122–1170, April 2010. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1266586625.

Gine:2011:RCP

[GN11] Evarist Giné and Richard Nickl. Rates of contraction for posterior distributions in L^{r}-metrics, 1 ≤ r ≤ ∞. Annals of Statistics, 39(6):2883–2911, December 2011. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1327413772.

Golubev:2010:AES

[GNZ10] Georgi K. Golubev, Michael Nussbaum, and Harrison H. Zhou. Asymptotic equivalence of spectral density estimation and Gaussian white noise. Annals of Statistics, 38 (1):181–214, February 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1262271613.

Godolphin:2018:DBS

[God18] Janet Godolphin. Designs with blocks of size two and applications to microarray experiments. Annals of Statistics, 46 (6A):2775–2805, December 2018. CODEN ASTSC7. ISSN

0090-5364. URL http://projecteuclid.org/euclid.aos/
1536307233.

Golubev:2010:UOI

[Gol10] Yuri Golubev. On universal oracle inequalities related to high-dimensional linear models. Annals of Statistics, 38 (5):2751–2780, October 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1279638539.

Genovese:2012:MES

[GPPVW12] Christopher R. Genovese, Marco Perone-Pacifico, Isabella Verdinelli, and Larry Wasserman. Manifold estimation and singular deconvolution under Hausdorff loss. Annals of Statistics, 40(2):941–963, April 2012. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1338515143.

Genovese:2014:NRE

[GPPVW14] Christopher R. Genovese, Marco Perone-Pacifico, Isabella Verdinelli, and Larry Wasserman. Nonparametric ridge estimation. Annals of Statistics, 42(4):1511–1545, August 2014. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1407420007.

Gandy:2013:ACP

[GRD13] Axel Gandy and Patrick Rubin-Delanchy. An algorithm to compute the power of Monte Carlo tests with guaranteed precision. Annals of Statistics, 41(1):125–142, February 2013. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1362493042.

Groeneboom:2014:MSL

[Gro14] Piet Groeneboom. Maximum smoothed likelihood estimators for the interval censoring model. Annals of Statistics, 42(5):2092–2137, October 2014. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1410440635.

Giraud:2015:APN

[GRSP15] Christophe Giraud, François Roueff, and Andres Sanchez-Perez. Aggregation of predictors for nonstationary sub-linear processes and online adaptive forecasting of time varying autoregressive processes. Annals of Statistics, 43(6):2412–2450, December 2015. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1444222080.

Goeman:2010:SRP

[GS10] Jelle J. Goeman and Aldo Solari. The sequential rejection principle of familywise error control. Annals of Statistics, 38(6):3782–3810, December 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1291126973.

Giraud:2012:DLV

[GT12] Christophe Giraud and Alexandre Tsybakov. Discussion: Latent variable graphical model selection via convex optimization. Annals of Statistics, 40(4):1984–1988, August 2012. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1351602531.

Guntuboyina:2012:ORC

[Gun12] Adityanand Guntuboyina. Optimal rates of convergence for convex set estimation from support functions. Annals of Statistics, 40(1):385–411, February 2012. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1334581747.

Gu:2018:RGS

[GWB18] Mengyang Gu, Xiaojing Wang, and James O. Berger. Robust Gaussian stochastic process emulation. Annals of Statistics, 46(6A):3038–3066, December 2018. CODEN ASTSC7. ISSN

0090-5364. URL http://projecteuclid.org/euclid.aos/
1536307242.

Gromping:2014:GRO

[GX14] Ulrike Grömping and Hongquan Xu. Generalized resolution for orthogonal arrays. Annals of Statistics, 42(3):918– 939, June 2014. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1400592647.

Gao:2018:MDN

[GXZ18] Fuqing Gao, Jie Xiong, and Xingqiu Zhao. Moderate deviations and nonparametric inference for monotone functions. Annals of Statistics, 46(3):1225–1254, June 2018. CODEN ASTSC7. ISSN 0090-5364. URL http://projecteuclid.

org/euclid.aos/1525313081.

Gao:2011:DML

[GZ11] Fuqing Gao and Xingqiu Zhao. Delta method in large deviations and moderate deviations for estimators. Annals of Statistics, 39(2):1211–1240, April 2011. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1304947048.

Gao:2015:ROP

[GZ15] Chao Gao and Harrison H. Zhou. Rate-optimal posterior contraction for sparse PCA. Annals of Statistics, 43(2):785– 818, April 2015. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1427115287.

Gao:2016:REB

[GZ16a] Chao Gao and Harrison H. Zhou. Rate exact Bayesian adaptation with modified block priors. Annals of Statistics, 44(1):318–345, February 2016. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1449755965.

Gu:2016:HDG

[GZ16b] Yuwen Gu and Hui Zou. High-dimensional generalizations of asymmetric least squares regression and their applications. Annals of Statistics, 44(6):2661–2694, December 2016. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1479891631.

Ho:2013:ATH

[HA13] Lam Si Tung Ho and Cécile Ané. Asymptotic theory with hierarchical autocorrelation: Ornstein–Uhlenbeck tree models. Annals of Statistics, 41(2):957–981, April 2013. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1369836966.

Hanneke:2011:RCA

[Han11] Steve Hanneke. Rates of convergence in active learning. Annals of Statistics, 39(1):333–361, February 2011. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1291388378.

Han:2019:NSU

[Han19] Xu Han. Nonparametric screening under conditional strictly convex loss for ultrahigh dimensional sparse data. Annals of Statistics, 47(4):1995–2022, August 2019. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https:

//projecteuclid.org/euclid.aos/1558425637.

Hsing:2016:LIS

[HBT16] Tailen Hsing, Thomas Brown, and Brian Thelen. Local intrinsic stationarity and its inference. Annals of Statistics, 44(5):2058–2088, October 2016. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1473685268.

He:2019:PVA

[HBZO19] Hera Y. He, Kinjal Basu, Qingyuan Zhao, and Art B. Owen. Permutation p-value approximation via generalized Stolarsky invariance. Annals of Statistics, 47(1):583–611, February 2019. CODEN ASTSC7. ISSN 0090-5364. URL http://

projecteuclid.org/euclid.aos/1543568599.

He:2018:SOA

[HCT18] Yuanzhen He, Ching-Shui Cheng, and Boxin Tang. Strong orthogonal arrays of strength two plus. Annals of Statistics, 46(2):457–468, April 2018. CODEN ASTSC7. ISSN

0090-5364. URL http://projecteuclid.org/euclid.aos/
1522742425.

Huckemann:2018:BND

[HE18] Stephan F. Huckemann and Benjamin Eltzner. Backward nested descriptors asymptotics with inference on stem cell differentiation. Annals of Statistics, 46(5):1994–2019, October 2018. CODEN ASTSC7. ISSN 0090-5364. URL http:/

/projecteuclid.org/euclid.aos/1534492826.

Hung:2019:RVE

[HF19] Kenneth Hung and William Fithian. Rank verification for exponential families. Annals of Statistics, 47(2):758–782, April 2019. CODEN ASTSC7. ISSN 0090-5364. URL http://

projecteuclid.org/euclid.aos/1547197237.

Hu:2012:APC

[HH12] Yanqing Hu and Feifang Hu. Asymptotic properties of covariate-adaptive randomization. Annals of Statistics, 40 (3):1794–1815, June 2012. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1350394517.

Hall:2013:SBM

[HH13] Peter Hall and Joel Horowitz. A simple bootstrap method for constructing nonparametric confidence bands for functions. Annals of Statistics, 41(4):1892–1921, August 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1378386242.

Huang:2010:VSN

[HHW10] Jian Huang, Joel L. Horowitz, and Fengrong Wei. Variable selection in nonparametric additive models. Annals of Statistics, 38(4):2282–2313, August 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1278861249.

Hsu:2019:MSF

[HIT19] Hsiang-Ling Hsu, Ching-Kang Ing, and Howell Tong. On model selection from a finite family of possibly misspecified time series models. Annals of Statistics, 47(2):1061–1087, April 2019. CODEN ASTSC7. ISSN 0090-5364. URL http:

//projecteuclid.org/euclid.aos/1547197248.

Hall:2010:IHC

[HJ10] Peter Hall and Jiashun Jin. Innovated higher criticism for detecting sparse signals in correlated noise. Annals of Statistics, 38(3):1686–1732, June 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1269452652.

He:2013:RMM

[HJY13] Yangbo He, Jinzhu Jia, and Bin Yu. Reversible MCMC on Markov equivalence classes of sparse directed acyclic graphs. Annals of Statistics, 41(4):1742–1779, August 2013. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1378386238.

Huang:2016:PCS

[HJY16] Shiqiong Huang, Jiashun Jin, and Zhigang Yao. Partial correlation screening for estimating large precision matrices, with applications to classification. Annals of Statistics, 44(5):2018–2057, October 2016. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1473685267.

Hormann:2010:WDF

[HK10] Siegfried Hörmann and Piotr Kokoszka. Weakly dependent functional data. Annals of Statistics, 38(3):1845–1884, June 2010. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1269452656.

Heinrich:2018:SIO

[HK18] Philippe Heinrich and Jonas Kahn. Strong identifiability and optimal minimax rates for finite mixture estimation. Annals of Statistics, 46(6A):2844–2870, December 2018. CODEN ASTSC7. ISSN 0090-5364. URL http://projecteuclid.

org/euclid.aos/1536307235.

Huckemann:2010:MDH

[HKKM10] Stephan F. Huckemann, Peter T. Kim, Ja-Yong Koo, and Axel Munk. Möbius deconvolution on the hyperbolic plane with application to impedance density estimation. Annals of Statistics, 38(4):2465–2498, August 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1278861254.

Haff:2011:MEM

[HKKR11] L. R. Haff, P. T. Kim, J.-Y. Koo, and D. St. P. Richards. Minimax estimation for mixtures of Wishart distributions. Annals of Statistics, 39(6):3417–3440, December 2011. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1330958685.

Hormann:2018:TPF

[HKN18] Siegfried Hörmann, Piotr Kokoszka, and Gilles Nisol. Testing for periodicity in functional time series. Annals of Statistics, 46(6A):2960–2984, December 2018. CODEN ASTSC7. ISSN

0090-5364. URL http://projecteuclid.org/euclid.aos/
1536307239.

Hong:2013:LFA

[HL13] Yongmiao Hong and Yoon-Jin Lee. A loss function approach to model specification testing and its relative efficiency. Annals of Statistics, 41(3):1166–1203, June 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1371150897.

Hirose:2018:EVR

[HL18] Masayo Yoshimori Hirose and Partha Lahiri. Estimating variance of random effects to solve multiple problems simultaneously. Annals of Statistics, 46(4):1721–1741, August 2018. CODEN ASTSC7. ISSN 0090-5364. URL http://

projecteuclid.org/euclid.aos/1530086431.

Hu:2019:HDC

[HLLZ19] Jiang Hu, Weiming Li, Zhi Liu, and Wang Zhou. Highdimensional covariance matrices in elliptical distributions with application to spherical test. Annals of Statistics, 47(1):527–555, February 2019. CODEN ASTSC7. ISSN

0090-5364. URL http://projecteuclid.org/euclid.aos/
1543568597.

Hall:2010:MVR

[HM10] Peter Hall and Hugh Miller. Modeling the variability of rankings. Annals of Statistics, 38(5):2652–2677, October 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1278861456.

Huang:2011:SLS

[HMLZ11] Jian Huang, Shuangge Ma, Hongzhe Li, and Cun-Hui Zhang. The sparse Laplacian shrinkage estimator for highdimensional regression. Annals of Statistics, 39(4):2021–2046, August 2011. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1314190622.

Hilgert:2013:MAT

[HMV13] Nadine Hilgert, André Mas, and Nicolas Verzelen. Minimax adaptive tests for the functional linear model. Annals of Statistics, 41(2):838–869, April 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1369836962.

Hoffmann:2011:AIC

[HN11] Marc Hoffmann and Richard Nickl. On adaptive inference and confidence bands. Annals of Statistics, 39(5):2383–2409, October 2011. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1322663462.

Ho:2016:CRP

[HN16] Nhat Ho and XuanLong Nguyen. Convergence rates of parameter estimation for some weakly identifiable finite mixtures. Annals of Statistics, 44(6):2726–2755, December 2016. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1479891633.

Hall:2010:OPC

[HP10] Peter Hall and Tung Pham. Optimal properties of centroidbased classifiers for very high-dimensional data. Annals of Statistics, 38(2):1071–1093, April 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1266586623.

Han:2018:SBE

[HP18] Kyunghee Han and Byeong U. Park. Smooth backfitting for errors-in-variables additive models. Annals of Statistics, 46(5):2216–2250, October 2018. CODEN ASTSC7. ISSN

0090-5364. URL http://projecteuclid.org/euclid.aos/
1534492834.

Hallin:2010:MQM

[HPŠ10a] Marc Hallin, Davy Paindaveine, and Miroslav Šiman. Multivariate quantiles and multiple-output regression quantiles: From L_{1} optimization to halfspace depth. Annals of Statistics, 38(2):635–669, April 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1266586607.

Hallin:2010:R

[HPŠ10b] Marc Hallin, Davy Paindaveine, and Miroslav Šiman. Rejoinder. Annals of Statistics, 38(2):694–703, April 2010. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1266586611.

Hallin:2010:ORB

[HPV10] Marc Hallin, Davy Paindaveine, and Thomas Verdebout. Optimal rank-based testing for principal components. Annals of Statistics, 38(6):3245–3299, December 2010. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1284988406.

Hall:2011:ANV

[HPWW11] Peter Hall, Tung Pham, M. P. Wand, and S. S. J. Wang. Asymptotic normality and valid inference for Gaussian variational approximation. Annals of Statistics, 39(5):2502–2532, October 2011. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1322663466.

Han:2016:TWL

[HPZ16] Xiao Han, Guangming Pan, and Bo Zhang. The Tracy– Widom law for the largest eigenvalue of F type matrices. Annals of Statistics, 44(4):1564–1592, August 2016. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1467894708.

Haaland:2011:AEL

[HQ11] Ben Haaland and Peter Z. G. Qian. Accurate emulators for large-scale computer experiments. Annals of Statistics, 39 (6):2974–3002, December 2011. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1327413775.

He:2014:CLT

[HQ14] Xu He and Peter Z. G. Qian. A central limit theorem for general orthogonal array based space-filling designs. Annals of Statistics, 42(5):1725–1750, October 2014. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1410440623.

Hualde:2011:GPM

[HR11] Javier Hualde and Peter M. Robinson. Gaussian pseudomaximum likelihood estimation of fractional time series models. Annals of Statistics, 39(6):3152–3181, December 2011.

CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1330958676.

Hoffmann:2015:APC

[HRSH15] Marc Hoffmann, Judith Rousseau, and Johannes Schmidt-Hieber. On adaptive posterior concentration rates. Annals of Statistics, 43(5):2259–2295, October 2015. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1442364152.

Hang:2017:BTI

[HS17] Hanyuan Hang and Ingo Steinwart. A Bernstein-type inequality for some mixing processes and dynamical systems with an application to learning. Annals of Statistics, 45(2):708– 743, April 2017. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1494921955.

Heckel:2019:ARP

[HSRW19] Reinhard Heckel, Nihar B. Shah, Kannan Ramchandran, and Martin J. Wainwright. Active ranking from pairwise comparisons and when parametric assumptions do not help. Annals of Statistics, 47(6):3099–3126, December 2019. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/euclid.aos/
1572487385.

Huang:2013:OIL

[HSY^{+}13] Jian Huang, Tingni Sun, Zhiliang Ying, Yi Yu, and Cun-Hui Zhang. Oracle inequalities for the lasso in the Cox model. Annals of Statistics, 41(3):1142–1165, June 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1371150896.

He:2014:CSO

[HT14] Yuanzhen He and Boxin Tang. A characterization of strong orthogonal arrays of strength three. Annals of Statistics, 42(4):1347–1360, August 2014. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1403715203.

Han:2017:OBC

[HTX17] Dong Han, Fugee Tsung, and Jinguo Xian. On the optimality of Bayesian change-point detection. Annals of Statistics, 45(4):1375–1402, August 2017. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1498636860.

Huang:2010:TCI

[Hua10a] Tzee-Ming Huang. Testing conditional independence using maximal nonlinear conditional correlation. Annals of Statistics, 38(4):2047–2091, August 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1278861242.

Huang:2010:QCC

[Hua10b] Yijian Huang. Quantile calculus and censored regression. Annals of Statistics, 38(3):1607–1637, June 2010. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1269452649.

Huckemann:2011:IIM

[Huc11] Stephan F. Huckemann. Intrinsic inference on the mean geodesic of planar shapes and tree discrimination by leaf growth. Annals of Statistics, 39(2):1098–1124, April 2011. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1304947045.

Han:2016:AEC

[HW16a] Qiyang Han and Jon A. Wellner. Approximation and estimation of s-concave densities via Rényi divergences. Annals of Statistics, 44(3):1332–1359, June 2016. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1460381695.

Hayashi:2016:IGA

[HW16b] Masahito Hayashi and Shun Watanabe. Information geometry approach to parameter estimation in Markov chains. Annals of Statistics, 44(4):1495–1535, August 2016. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1467894706.

Hu:2017:RDD

[HW17] Rui Hu and Douglas P. Wiens. Robust discrimination designs over Hellinger neighbourhoods. Annals of Statistics, 45(4):1638–1663, August 2017. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1498636869.

Han:2019:CRL

[HW19] Qiyang Han and Jon A. Wellner. Convergence rates of least squares regression estimators with heavy-tailed errors. Annals of Statistics, 47(4):2286–2319, August 2019. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/euclid.aos/
1558425646.

Han:2019:IRG

[HWCS19] Qiyang Han, Tengyao Wang, Sabyasachi Chatterjee, and Richard J. Samworth. Isotonic regression in general dimensions. Annals of Statistics, 47(5):2440–2471, October 2019. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/euclid.aos/
1564797853.

He:2013:CQA

[HWH13a] Xuming He, Lan Wang, and Hyokyoung Grace Hong. Correction: “Quantile-adaptive model-free variable screening for high-dimensional heterogeneous data”. Annals of Statistics, 41(5):2699, October 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1386598636. See

[HWH13b].

He:2013:QAM

[HWH13b] Xuming He, Lan Wang, and Hyokyoung Grace Hong. Quantile-adaptive model-free variable screening for highdimensional heterogeneous data. Annals of Statistics, 41 (1):342–369, February 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1364302746. See correc-

tion [HWH13a].

Hu:2015:SLO

[HYS15] Linwei Hu, Min Yang, and John Stufken. Saturated locally optimal designs under differentiable optimality criteria. Annals of Statistics, 43(1):30–56, February 2015. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1416322035.

Huang:2010:BGS

[HZ10] Junzhou Huang and Tong Zhang. The benefit of group sparsity. Annals of Statistics, 38(4):1978–2004, August 2010. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1278861240.

Ionides:2011:IF

[IBAK11] Edward L. Ionides, Anindya Bhadra, Yves Atchadé, and Aaron King. Iterated filtering. Annals of Statistics, 39(3):1776–1802, June 2011. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1311600283.

Ilmonen:2011:SEI

[IP11] Pauliina Ilmonen and Davy Paindaveine. Semiparametrically efficient inference based on signed ranks in symmetric independent component models. Annals of Statistics, 39(5):2448–2476, October 2011. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1322663464.

Ingster:2012:MSD

[ISS12] Yuri I. Ingster, Theofanis Sapatinas, and Irina A. Suslina. Minimax signal detection in ill-posed inverse problems. Annals of Statistics, 40(3):1524–1549, June 2012. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1346850064.

Jacob:2010:CLS

[Jac10] Christine Jacob. Conditional least squares estimation in nonstationary nonlinear stochastic regression models. Annals of Statistics, 38(1):566–597, February 2010. CODEN ASTSC7.

ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1262271624.

James:2017:BPC

[Jam17] Lancelot F. James. Bayesian Poisson calculus for latent feature modeling via generalized Indian Buffet Process priors. Annals of Statistics, 45(5):2016–2045, October 2017. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

Jankowski:2014:CLF

[Jan14] Hanna Jankowski. Convergence of linear functionals of the Grenander estimator under misspecification. Annals of Statistics, 42(2):625–653, April 2014. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1400592172.

Johndrow:2017:TDS

[JBD17] James E. Johndrow, Anirban Bhattacharya, and David B. Dunson. Tensor decompositions and sparse log-linear models. Annals of Statistics, 45(1):1–38, February 2017. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1487667616.

Janzing:2013:QCI

[JBGWS13] Dominik Janzing, David Balduzzi, Moritz Grosse-Wentrup, and Bernhard Schölkopf. Quantifying causal influences. Annals of Statistics, 41(5):2324–2358, October 2013. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1383661266.

Jensen:2010:SPA

[Jen10] Jens Ledet Jensen. On some problems in the article “Efficient Likelihood Estimation in State Space Models” by Cheng–Der Fuh [Ann. Statist. 34 (2006) 2026–2068]. Annals of Statistics, 38(2):1279–1281, April 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1266586630. See [Fuh06].

Jiang:2010:EAM

[JFF10] Jiancheng Jiang, Yingying Fan, and Jianqing Fan. Estimation in additive models with highly or nonhighly correlated covariates. Annals of Statistics, 38(3):1403–1432, June 2010. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1268056621.

Johnson:2012:VTO

[JG12] Leif T. Johnson and Charles J. Geyer. Variable transformation to obtain geometric ergodicity in the random-walk Metropolis algorithm. Annals of Statistics, 40(6):3050–3076, December 2012. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1361542074. See correction [JG13].

Johnson:2013:CVT

[JG13] Leif T. Johnson and Charles J. Geyer. Correction: “Variable transformation to obtain geometric ergodicity in the randomwalk Metropolis algorithm”. Annals of Statistics, 41(5):2698, October 2013. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1386598635. See [JG12].

Jiang:2013:SAC

[Jia13] Jiming Jiang. The subset argument and consistency of MLE in GLMM: Answer to an open problem and beyond. Annals of Statistics, 41(1):177–195, February 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1362493044.

Jin:2015:FCD

[Jin15] Jiashun Jin. Fast community detection by SCORE. Annals of Statistics, 43(1):57–89, February 2015. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1416322036.

Jirak:2012:SCB

[Jir12] Moritz Jirak. Simultaneous confidence bands for Yule– Walker estimators and order selection. Annals of Statistics, 40(1):494–528, February 2012. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1336396181.

Jirak:2015:UCP

[Jir15] Moritz Jirak. Uniform change point tests in high dimension. Annals of Statistics, 43(6):2451–2483, December 2015.

CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1444222081.

Ji:2012:UDO

[JJ12] Pengsheng Ji and Jiashun Jin. UPS delivers optimal phase diagram in high-dimensional variable selection. Annals of Statistics, 40(1):73–103, February 2012. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1331830775.

Jing:2012:MHF

[JKL12] Bing-Yi Jing, Xin-Bing Kong, and Zhi Liu. Modeling highfrequency financial data by pure jump processes. Annals of Statistics, 40(2):759–784, April 2012. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1337268211.

Juditsky:2012:AGR

[JKNP12] Anatoli Juditsky, Fatma Kılınç Karzan, Arkadi Nemirovski, and Boris Polyak. Accuracy guarantees for ℓ_{1} recovery of block-sparse signals. Annals of Statistics, 40(6):3077–3107, December 2012. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1361542075.

Jin:2017:PTH

[JKW17] Jiashun Jin, Zheng Tracy Ke, and Wanjie Wang. Phase transitions for high dimensional clustering and related problems. Annals of Statistics, 45(5):2151–2189, October 2017. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

Jiang:2014:VSG

[JL14] Bo Jiang and Jun S. Liu. Variable selection for general index models via sliced inverse regression. Annals of Statistics, 42(5):1751–1786, October 2014. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1410440624.

Jiang:2016:HDM

[JLP^{+}16] Jiming Jiang, Cong Li, Debashis Paul, Can Yang, and Hongyu Zhao. On high-dimensional misspecified mixed model analysis in genome-wide association study. Annals of Statistics, 44(5):2127–2160, October 2016. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1473685271.

Jiang:2017:REA

[JLX17] Tiefeng Jiang, Kevin Leder, and Gongjun Xu. Rare-event analysis for extremal eigenvalues of white Wishart matrices. Annals of Statistics, 45(4):1609–1637, August 2017. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1498636868.

Javanmard:2018:DLO

[JM18a] Adel Javanmard and Andrea Montanari. Debiasing the lasso: Optimal sample size for Gaussian designs. Annals of Statistics, 46(6A):2593–2622, December 2018. CODEN ASTSC7.

ISSN 0090-5364. URL http://projecteuclid.org/euclid.
aos/1536307227.

Javanmard:2018:ORC

[JM18b] Adel Javanmard and Andrea Montanari. Online rules for control of false discovery rate and false discovery exceedance. Annals of Statistics, 46(2):526–554, April 2018. CODEN ASTSC7. ISSN 0090-5364. URL http://projecteuclid.

org/euclid.aos/1522742428.

Jirak:2014:AFE

[JMR14] Moritz Jirak, Alexander Meister, and Markus Reiß. Adaptive function estimation in nonparametric regression with one-sided errors. Annals of Statistics, 42(5):1970–2002, October 2014. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1410440631.

Jiang:2015:FKS

[JMW15] Fei Jiang, Yanyuan Ma, and Yuanjia Wang. Fused kernelspline smoothing for repeatedly measured outcomes in a generalized partially linear model with functional single index. Annals of Statistics, 43(5):1929–1958, October 2015. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1438606849.

Juditsky:2018:NOL

[JN18] Anatoli Juditsky and Arkadi Nemirovski. Near-optimality of linear recovery in Gaussian observation scheme under ∥ · ∥_{2}^{2}-loss. Annals of Statistics, 46(4):1603–1629, August 2018. CODEN ASTSC7. ISSN 0090-5364. URL http://

projecteuclid.org/euclid.aos/1530086427.

Johnson:2013:UMP

[Joh13] Valen E. Johnson. Uniformly most powerful Bayesian tests. Annals of Statistics, 41(4):1716–1741, August 2013. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1378386237.

Jacod:2013:TRV

[JP13] Jean Jacod and Mark Podolskij. A test for the rank of the volatility process: The random perturbation approach. Annals of Statistics, 41(5):2391–2427, October 2013. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1383661268.

Jentsch:2015:CME

[JP15] Carsten Jentsch and Dimitris N. Politis. Covariance matrix estimation and linear process bootstrap for multivariate time series of possibly increasing dimension. Annals of Statistics, 43(3):1117–1140, June 2015. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1431695640.

Jing:2010:NES

[JPSZ10] Bing-Yi Jing, Guangming Pan, Qi-Man Shao, and Wang Zhou. Nonparametric estimate of spectral density functions of sample covariance matrices: A first step. Annals of Statistics, 38(6):3724–3750, December 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1291126971.

Jacod:2010:LTM

[JPV10] Jean Jacod, Mark Podolskij, and Mathias Vetter. Limit theorems for moving averages of discretized processes plus noise. Annals of Statistics, 38(3):1478–1545, June 2010. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1269452645.

Jacod:2013:QOF

[JR13] Jean Jacod and Mathieu Rosenbaum. Quarticity and other functionals of volatility: Efficient estimation. Annals of Statistics, 41(3):1462–1484, June 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1375362556.

Jacod:2014:EEI

[JT14] Jean Jacod and Viktor Todorov. Efficient estimation of integrated volatility in presence of infinite variation jumps. Annals of Statistics, 42(3):1029–1069, June 2014. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1400592651.

Jacob:2015:NUE

[JT15] Pierre E. Jacob and Alexandre H. Thiery. On nonnegative unbiased estimators. Annals of Statistics, 43(2):769– 784, April 2015. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1425398508.

Jankova:2018:SEB

[JvdG18] Jana Janková and Sara van de Geer. Semiparametric efficiency bounds for high-dimensional models. Annals of Statistics, 46(5):2336–2359, October 2018. CODEN ASTSC7. ISSN

0090-5364. URL http://projecteuclid.org/euclid.aos/
1534492838.

Jiang:2010:CAF

[JW10] Ci-Ren Jiang and Jane-Ling Wang. Covariate adjusted functional principal components analysis for longitudinal data. Annals of Statistics, 38(2):1194–1226, April 2010. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1266586627.

Jiang:2011:FSI

[JW11] Ci-Ren Jiang and Jane-Ling Wang. Functional single index models for longitudinal data. Annals of Statistics, 39(1):362–388, February 2011. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1291388379.

Jiang:2015:CIR

[JW15] Ci-Ren Jiang and Jane-Ling Wang. Correction to “Inverse regression for longitudinal data”. Annals of Statistics, 43(5):2326–2329, October 2015. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1442364154. See

[JYW14].

Jin:2016:IFP

[JW16a] Jiashun Jin and Wanjie Wang. Influential features PCA for high dimensional clustering. Annals of Statistics, 44 (6):2323–2359, December 2016. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1479891617. See discus-

sion [ACV16, Nad16, CZ16, ST16b] and rejoinder [JW16b].

Jin:2016:RIF

[JW16b] Jiashun Jin and Wanjie Wang. Rejoinder: “Influential features PCA for high dimensional clustering”. Annals of Statistics, 44(6):2387–2400, December 2016. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1479891622.

Jiang:2018:CCD

[JWJW18] Bai Jiang, Tung-Yu Wu, Yifan Jin, and Wing H. Wong. Convergence of contrastive divergence algorithm in exponential family. Annals of Statistics, 46(6A):3067–3098, December 2018. CODEN ASTSC7. ISSN 0090-5364. URL http://

projecteuclid.org/euclid.aos/1536307243.

Jimenez:2011:NES

[JY11] Raúl Jiménez and J. E. Yukich. Nonparametric estimation of surface integrals. Annals of Statistics, 39(1):232–260, February 2011. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1291388374.

Jiang:2013:CLT

[JY13] Tiefeng Jiang and Fan Yang. Central limit theorems for classical likelihood ratio tests for high-dimensional normal distributions. Annals of Statistics, 41(4):2029–2074, August 2013. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1382547512.

Joseph:2016:IRS

[JY16] Antony Joseph and Bin Yu. Impact of regularization on spectral clustering. Annals of Statistics, 44(4):1765–1791, August 2016. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1467894715.

Jiang:2014:IRL

[JYW14] Ci-Ren Jiang, Wei Yu, and Jane-Ling Wang. Inverse regression for longitudinal data. Annals of Statistics, 42(2):563– 591, April 2014. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1400592170. See correction [JW15].

Kim:2010:OT

[KA10] Min Hee Kim and Michael G. Akritas. Order thresholding. Annals of Statistics, 38(4):2314–2350, August 2010. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1278861250.

Kato:2012:EFL

[Kat12] Kengo Kato. Estimation in functional linear quantile regression. Annals of Statistics, 40(6):3108–3136, December 2012. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1361542076.

Kato:2013:QBA

[Kat13] Kengo Kato. Quasi-Bayesian analysis of nonparametric instrumental variables models. Annals of Statistics, 41 (5):2359–2390, October 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1383661267.

Kaufmann:2018:BIP

[Kau18] Emilie Kaufmann. On Bayesian index policies for sequential resource allocation. Annals of Statistics, 46(2):842–865, April 2018. CODEN ASTSC7. ISSN 0090-5364. URL http://

projecteuclid.org/euclid.aos/1522742438.

Koopmeiners:2011:APS

[KF11] Joseph S. Koopmeiners and Ziding Feng. Asymptotic properties of the sequential empirical ROC, PPV and NPV curves under case-control sampling. Annals of Statistics, 39 (6):3234–3261, December 2011. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1330958678.

Kyung:2010:EDR

[KGC10] Minjung Kyung, Jeff Gill, and George Casella. Estimation in Dirichlet random effects models. Annals of Statistics, 38(2):979–1009, April 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1266586620.

Kim:2018:ALC

[KGS18] Arlene K. H. Kim, Adityanand Guntuboyina, and Richard J. Samworth. Adaptation in log-concave density estimation. Annals of Statistics, 46(5):2279–2306, October 2018. CODEN ASTSC7. ISSN 0090-5364. URL http://projecteuclid.

org/euclid.aos/1534492836.

Khare:2011:SAC

[KH11] Kshitij Khare and James P. Hobert. A spectral analytic comparison of trace-class data augmentation algorithms and their sandwich variants. Annals of Statistics, 39(5):2585–2606, October 2011. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1324563348.

Khmaladze:2013:NDF

[Khm13] Estate Khmaladze. Note on distribution free testing for discrete distributions. Annals of Statistics, 41(6):2979–2993, December 2013. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1388545675.

Kim:2016:CMS

[KJ16] Yongdai Kim and Jong-June Jeon. Consistent model selection criteria for quadratically supported risks. Annals of Statistics, 44(6):2467–2496, December 2016. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1479891625.

Ke:2014:CAS

[KJF14] Zheng Tracy Ke, Jiashun Jin, and Jianqing Fan. Covariate assisted screening and estimation. Annals of Statistics, 42(6):2202–2242, December 2014. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1413810726.

Koltchinskii:2017:NAC

[KL17] Vladimir Koltchinskii and Karim Lounici. Normal approximation and concentration of spectral projectors of sample covariance. Annals of Statistics, 45(1):121–157, February 2017. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1487667619.

Kong:2017:IPH

[KLFL17] Yinfei Kong, Daoji Li, Yingying Fan, and Jinchi Lv. Interaction pursuit in high-dimensional multi-response regression via distance correlation. Annals of Statistics, 45(2):897– 922, April 2017. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1494921961.

Kong:2015:TPJ

[KLJ15] Xin-Bing Kong, Zhi Liu, and Bing-Yi Jing. Testing for purejump processes for high-frequency data. Annals of Statistics, 43(2):847–877, April 2015. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1427115289.

Kaiser:2012:GFT

[KLN12] Mark S. Kaiser, Soumendra N. Lahiri, and Daniel J. Nordman. Goodness of fit tests for a class of Markov random field models. Annals of Statistics, 40(1):104–130, February 2012. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1331830776.

Koltchinskii:2011:NNP

[KLT11] Vladimir Koltchinskii, Karim Lounici, and Alexandre B. Tsybakov. Nuclear-norm penalization and optimal rates for noisy low-rank matrix completion. Annals of Statistics, 39 (5):2302–2329, October 2011. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1322663459.

Kai:2011:NEE

[KLZ11] Bo Kai, Runze Li, and Hui Zou. New efficient estimation and variable selection methods for semiparametric varyingcoefficient partially linear models. Annals of Statistics, 39(1):305–332, February 2011. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1291388377.

Ke:2016:SIP

[KLZ16] Yuan Ke, Jialiang Li, and Wenyang Zhang. Structure identification in panel data analysis. Annals of Statistics, 44(3):1193–1233, June 2016. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1460381691.

Klemela:2010:ERM

[KM10a] Jussi Klemelä and Enno Mammen. Empirical risk minimization in inverse problems. Annals of Statistics, 38(1):482–511, February 2010. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1262271621.

Koenker:2010:QCD

[KM10b] Roger Koenker and Ivan Mizera. Quasi-concave density estimation. Annals of Statistics, 38(5):2998–3027, October 2010. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1282315406.

Kong:2010:D

[KM10c] Linglong Kong and Ivan Mizera. Discussion. Annals of Statistics, 38(2):685–693, April 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1266586610.

Kuipers:2014:ASG

[KMH14] Jack Kuipers, Giusi Moffa, and David Heckerman. Addendum on the scoring of Gaussian directed acyclic graphical models. Annals of Statistics, 42(4):1689–1691, August 2014. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1407420013.

Koul:2012:TPT

[KMS12] Hira L. Koul, Ursula U. Müller, and Anton Schick. The transfer principle: A tool for complete case analysis. Annals of Statistics, 40(6):3031–3049, December 2012. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1360332192.

Kerkyacharian:2011:LSD

[KNP11] Gérard Kerkyacharian, Thanh Mai Pham Ngoc, and Dominique Picard. Localized spherical deconvolution. Annals of Statistics, 39(2):1042–1068, April 2011. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1302268086.

Krauthgamer:2015:DSR

[KNV15] Robert Krauthgamer, Boaz Nadler, and Dan Vilenchik. Do semidefinite relaxations solve sparse PCA up to the information limit? Annals of Statistics, 43(3):1300–1322, June 2015. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1431695645.

Koike:2019:GAM

[Koi19] Yuta Koike. Gaussian approximation of maxima of Wiener functionals and its application to high-frequency data. Annals of Statistics, 47(3):1663–1687, June 2019. CODEN ASTSC7.

ISSN 0090-5364. URL http://projecteuclid.org/euclid.
aos/1550026853.

Koltchinskii:2011:NEP

[Kol11] Vladimir Koltchinskii. Von Neumann entropy penalization and low-rank matrix estimation. Annals of Statistics, 39 (6):2936–2973, December 2011. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1327413774.

Kong:2018:SIV

[Kon18] Xin-Bing Kong. On the systematic and idiosyncratic volatility with large panel high-frequency data. Annals of Statistics, 46(3):1077–1108, June 2018. CODEN ASTSC7. ISSN

0090-5364. URL http://projecteuclid.org/euclid.aos/
1525313076.

Kirch:2011:TBR

[KP11] Claudia Kirch and Dimitris N. Politis. TFT-bootstrap: Resampling time series in the frequency domain to obtain replicates in the time domain. Annals of Statistics, 39(3):1427– 1470, June 2011. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1305292042.

Klopp:2015:SHD

[KP15] Olga Klopp and Marianna Pensky. Sparse high-dimensional varying coefficient model: Nonasymptotic minimax study. Annals of Statistics, 43(3):1273–1299, June 2015. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1431695644.

Kontorovich:2019:ELB

[KP19] Aryeh Kontorovich and Iosif Pinelis. Exact lower bounds for the agnostic probably-approximately-correct (PAC) machine learning model. Annals of Statistics, 47(5):2822–2854, October 2019. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL https://projecteuclid.org/
euclid.aos/1564797865.

Kreiss:2011:RVA

[KPP11] Jens-Peter Kreiss, Efstathios Paparoditis, and Dimitris N. Politis. On the range of validity of the autoregressive sieve bootstrap. Annals of Statistics, 39(4):2103–2130, August 2011. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1319595460.

Kneip:2016:FLR

[KPS16] Alois Kneip, Dominik Poß, and Pascal Sarda. Functional linear regression with points of impact. Annals of Statistics, 44(1):1–30, February 2016. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1449755955.

Khare:2017:BAE

[KPS17] Kshitij Khare, Subhadip Pal, and Zhihua Su. A Bayesian approach for envelope models. Annals of Statistics, 45 (1):196–222, February 2017. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1487667621.

Khare:2011:WDD

[KR11a] Kshitij Khare and Bala Rajaratnam. Wishart distributions for decomposable covariance graph models. Annals of Statistics, 39(1):514–555, February 2011. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1297779855.

Kolassa:2011:SAL

[KR11b] John Kolassa and John Robinson. Saddlepoint approximations for likelihood ratio like statistics with applications to permutation tests. Annals of Statistics, 39(6):3357–3368, December 2011. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1330958682.

Koltchinskii:2013:LRE

[KR13] Vladimir Koltchinskii and Pedro Rangel. Low rank estimation of smooth kernels on graphs. Annals of Statistics, 41(2):604–640, April 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1366980559.

Kalogeropoulos:2010:ISV

[KRD10] Konstantinos Kalogeropoulos, Gareth O. Roberts, and Petros Dellaportas. Inference for stochastic volatility models using time change transformations. Annals of Statistics, 38(2):784– 807, April 2010. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1266586614.

Kubjas:2015:FPE

[KRS15] Kaie Kubjas, Elina Robeva, and Bernd Sturmfels. Fixed points EM algorithm and nonnegative rank boundaries. Annals of Statistics, 43(1):422–461, February 2015. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1423230085.

Kneip:2011:FMV

[KS11] Alois Kneip and Pascal Sarda. Factor models and variable selection in high-dimensional regression analysis. Annals of Statistics, 39(5):2410–2447, October 2011. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1322663463.

Kim:2016:GRC

[KS16] Arlene K. H. Kim and Richard J. Samworth. Global rates of convergence in log-concave density estimation. Annals of Statistics, 44(6):2756–2779, December 2016. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1479891634.

Koul:2018:GFT

[KSZ18] Hira L. Koul, Weixing Song, and Xiaoqing Zhu. Goodnessof-fit testing of error distribution in linear measurement error models. Annals of Statistics, 46(5):2479–2510, October 2018. CODEN ASTSC7. ISSN 0090-5364. URL http://

projecteuclid.org/euclid.aos/1534492842.

Klopp:2017:OIN

[KTV17] Olga Klopp, Alexandre B. Tsybakov, and Nicolas Verzelen. Oracle inequalities for network models and sparse graphon estimation. Annals of Statistics, 45(1):316–354, February 2017. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1487667625.

Kuelbs:2010:AIH

[KV10] Jim Kuelbs and Anand N. Vidyashankar. Asymptotic inference for high-dimensional data. Annals of Statistics, 38(2):836–869, April 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1266586616.

Kong:2017:SES

[KV17] Weihao Kong and Gregory Valiant. Spectrum estimation from samples. Annals of Statistics, 45(5):2218–2247, October 2017. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

Knapik:2011:BIP

[KvdVvZ11] B. T. Knapik, A. W. van der Vaart, and J. H. van Zanten. Bayesian inverse problems with Gaussian priors. Annals of Statistics, 39(5):2626–2657, October 2011. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1324563350.

Kong:2014:ACQ

[KX14] Efang Kong and Yingcun Xia. An adaptive composite quantile approach to dimension reduction. Annals of Statistics, 42(4):1657–1688, August 2014. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1407420012.

Koltchinskii:2010:SMK

[KY10] Vladimir Koltchinskii and Ming Yuan. Sparsity in multiple kernel learning. Annals of Statistics, 38(6):3660–3695, December 2010. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1291126969.

Lahiri:2010:EES

[Lah10] S. N. Lahiri. Edgeworth expansions for studentized statistics under weak dependence. Annals of Statistics, 38(1):388–434, February 2010. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1262271619.

Li:2011:PSV

[LAL11] Bing Li, Andreas Artemiou, and Lexin Li. Principal support vector machines for linear and nonlinear sufficient dimension reduction. Annals of Statistics, 39(6):3182–3210, December 2011. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1330958677.

Lam:2016:NER

[Lam16] Clifford Lam. Nonparametric eigenvalue-regularized precision or covariance matrix estimator. Annals of Statistics, 44(3):928–953, June 2016. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1460381682.

Liu:2015:MPL

[LAP15] Haoyang Liu, Alexander Aue, and Debashis Paul. On the Marčenko–Pastur law for linear time series. Annals of Statistics, 43(2):675–712, April 2015. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1425398505.

Lin:2010:NFM

[LBST10] C. Devon Lin, Derek Bingham, Randy R. Sitter, and Boxin Tang. A new and flexible method for constructing designs for computer experiments. Annals of Statistics, 38(3):1460– 1477, June 2010. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1268056623.

Liu:2010:AEL

[LC10] Yukun Liu and Jiahua Chen. Adjusted empirical likelihood with high-order precision. Annals of Statistics, 38(3):1341– 1362, June 2010. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1268056619.

Li:2012:TST

[LC12] Jun Li and Song Xi Chen. Two sample tests for highdimensional covariance matrices. Annals of Statistics, 40(2):908–940, April 2012. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1338515142.

Leung:2018:TIH

[LD18] Dennis Leung and Mathias Drton. Testing independence in high dimensions with sums of rank correlations. Annals of Statistics, 46(1):280–307, February 2018. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

Leeb:2013:CDL

[Lee13] Hannes Leeb. On the conditional distributions of lowdimensional projections from high-dimensional data. Annals of Statistics, 41(2):464–483, April 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1366138198.

Lei:2016:GFT

[Lei16] Jing Lei. A goodness-of-fit test for stochastic block models. Annals of Statistics, 44(1):401–424, February 2016. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1452004791.

Lepski:2013:MDE

[Lep13] Oleg Lepski. Multivariate density estimation under sup-norm loss: Oracle approach, adaptation and independence structure. Annals of Statistics, 41(2):1005–1034, April 2013. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1369836968.

Lepski:2015:AEA

[Lep15] Oleg Lepski. Adaptive estimation over anisotropic functional classes via oracle approach. Annals of Statistics, 43(3):1178– 1242, June 2015. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1431695642.

Lerasle:2011:OMS

[Ler11] Matthieu Lerasle. Optimal model selection for density estimation of stationary data under various mixing conditions. Annals of Statistics, 39(4):1852–1877, August 2011. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1311688538.

Levrard:2015:NBV

[Lev15] Clément Levrard. Nonasymptotic bounds for vector quantization in Hilbert spaces. Annals of Statistics, 43(2):592– 619, April 2015. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1424787429.

Li:2017:BDI

[LG17] Meng Li and Subhashis Ghosal. Bayesian detection of image boundaries. Annals of Statistics, 45(5):2190–2217, October 2017. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

Lai:2011:EPF

[LGS11] Tze Leung Lai, Shulamith T. Gross, and David Bo Shen. Evaluating probability forecasts. Annals of Statistics, 39 (5):2356–2382, October 2011. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1322663461.

Li:2010:DDE

[LH10a] Yehua Li and Tailen Hsing. Deciding the dimension of effective dimension reduction space for functional and highdimensional data. Annals of Statistics, 38(5):3028–3062, October 2010. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1283175988.

Li:2010:UCR

[LH10b] Yehua Li and Tailen Hsing. Uniform convergence rates for nonparametric regression and principal component analysis in functional/longitudinal data. Annals of Statistics, 38 (6):3321–3351, December 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1284988408.

Liu:2012:HDS

[LHY^{+}12] Han Liu, Fang Han, Ming Yuan, John Lafferty, and Larry Wasserman. High-dimensional semiparametric Gaussian copula graphical models. Annals of Statistics, 40(4):2293–2326, August 2012. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1358951383.

Li:2013:MLE

[Li13] Chenxu Li. Maximum-likelihood estimation for diffusion processes via closed-form density expansions. Annals of Statistics, 41(3):1350–1380, June 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1372979641.

Li:2014:SOA

[Li14] Ke Li. Second-order asymptotics for quantum hypothesis testing. Annals of Statistics, 42(1):171–189, February 2014. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1392733184.

Li:2016:DQS

[Li16a] Ke Li. Discriminating quantum states: The multiple Chernoff distance. Annals of Statistics, 44(4):1661–1679, August 2016. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1467894711.

Li:2016:E

[Li16b] Runze Li. Editorial. Annals of Statistics, 44(5):1817–1820, October 2016. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1473685257.

Liang:2010:TAS

[Lia10] Faming Liang. Trajectory averaging for stochastic approximation MCMC algorithms. Annals of Statistics, 38 (5):2823–2856, October 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1279638541.

Liu:2013:GGM

[Liu13] Weidong Liu. Gaussian graphical model estimation with false discovery rate control. Annals of Statistics, 41(6):2948–2978, December 2013. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1388545674.

Liu:2017:SSD

[Liu17] Weidong Liu. Structural similarity and difference testing on multiple sparse Gaussian graphical models. Annals of Statistics, 45(6):2680–2707, December 2017. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

Liao:2010:BAM

[LJ10] Yuan Liao and Wenxin Jiang. Bayesian analysis in moment inequality models. Annals of Statistics, 38(1):275–316, February 2010. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1262271616.

Liao:2011:PCN

[LJ11] Yuan Liao and Wenxin Jiang. Posterior consistency of nonparametric conditional moment restricted models. Annals of Statistics, 39(6):3003–3031, December 2011. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1327413776.

Li:2018:MTA

[LJ18] Jialiang Li and Baisuo Jin. Multi-threshold accelerated failure time model. Annals of Statistics, 46(6A):2657–2682, December 2018. CODEN ASTSC7. ISSN 0090-5364. URL http:

//projecteuclid.org/euclid.aos/1536307229.

Lloyd:2010:LES

[LK10] Chris Lloyd and Paul Kabaila. Letter to the Editor: Some comments on “On construction of the smallest one-sided confidence interval for the difference of two proportions”. Annals of Statistics, 38(6):3840–3841, December 2010. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1291126977.

Li:2010:DFM

[LKA10] Bing Li, Min Kyung Kim, and Naomi Altman. On dimension folding of matrix- or array-valued statistical objects. Annals of Statistics, 38(2):1094–1121, April 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1266586624.

Li:2015:MSS

[LKZ15] Degui Li, Yuan Ke, and Wenyang Zhang. Model selection and structure specification in ultra-high dimensional generalised semi-varying coefficient models. Annals of Statistics, 43(6):2676–2705, December 2015. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1444222089.

Levy-Leduc:2011:APU

[LLBM^{+}11] C. Lévy-Leduc, H. Boistard, E. Moulines, M. S. Taqqu, and V. A. Reisen. Asymptotic properties of U-processes under long-range dependence. Annals of Statistics, 39(3):1399– 1426, June 2011. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1305292041.

Lee:2013:GTN

[LLC13] Kuang-Yao Lee, Bing Li, and Francesca Chiaromonte. A general theory for nonlinear sufficient dimension reduction: Formulation and estimation. Annals of Statistics, 41(1):221–249, February 2013. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1364302741.

Lee:2019:MPC

[LLL19] Kyoungjae Lee, Jaeyong Lee, and Lizhen Lin. Minimax posterior convergence rates and model selection consistency in high-dimensional DAG models based on sparse Cholesky factors. Annals of Statistics, 47(6):3413–3437, December 2019. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/euclid.aos/
1572487398.

Lv:2018:OIS

[LLLH18] Shaogao Lv, Huazhen Lin, Heng Lian, and Jian Huang. Oracle inequalities for sparse additive quantile regression in reproducing kernel Hilbert space. Annals of Statistics, 46(2):781–813, April 2018. CODEN ASTSC7. ISSN

0090-5364. URL http://projecteuclid.org/euclid.aos/
1522742436.

Liang:2010:ETP

[LLLT10] Hua Liang, Xiang Liu, Runze Li, and Chih-Ling Tsai. Estimation and testing for partially linear single-index models. Annals of Statistics, 38(6):3811–3836, December 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1291126974.

Le:2016:OLR

[LLV16] Can M. Le, Elizaveta Levina, and Roman Vershynin. Optimization via low-rank approximation for community detection in networks. Annals of Statistics, 44(1):373–400, February 2016. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1452004790.

Li:2019:EEI

[LLX19] Jia Li, Yunxiao Liu, and Dacheng Xiu. Efficient estimation of integrated volatility functionals via multiscale jackknife. Annals of Statistics, 47(1):156–176, February 2019. CODEN ASTSC7. ISSN 0090-5364. URL http://projecteuclid.

org/euclid.aos/1543568585.

Luo:2014:EDR

[LLY14] Wei Luo, Bing Li, and Xiangrong Yin. On efficient dimension reduction with respect to a statistical functional of interest. Annals of Statistics, 42(1):382–412, February 2014. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1395234982.

Li:2018:CIC

[LLY18] Xiaoou Li, Jingchen Liu, and Zhiliang Ying. Chernoff index for Cox test of separate parametric families. Annals of Statistics, 46(1):1–29, February 2018. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

Li:2019:THD

[LLYY19] Zeng Li, Clifford Lam, Jianfeng Yao, and Qiwei Yao. On testing for high-dimensional white noise. Annals of Statistics, 47(6):3382–3412, December 2019. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https:

//projecteuclid.org/euclid.aos/1572487397.

Lahiri:2012:PEL

[LM12a] Soumendra N. Lahiri and Subhadeep Mukhopadhyay. A penalized empirical likelihood method in high dimensions. Annals of Statistics, 40(5):2511–2540, October 2012. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1359987529.

Lauritzen:2012:DLV

[LM12b] Steffen Lauritzen and Nicolai Meinshausen. Discussion: Latent variable graphical model selection via convex optimization. Annals of Statistics, 40(4):1973–1977, August 2012. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1351602529.

Lecue:2012:GNO

[LM12c] Guillaume Lecué and Shahar Mendelson. General nonexact oracle inequalities for classes with a subexponential envelope. Annals of Statistics, 40(2):832–860, April 2012. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1338515139.

Letac:2012:BFG

[LM12d] Gérard Letac and Hélène Massam. Bayes factors and the geometry of discrete hierarchical loglinear models. Annals of Statistics, 40(2):861–890, April 2012. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1338515140.

Low:2015:DFC

[LM15] Mark G. Low and Zongming Ma. Discussion of “Frequentist coverage of adaptive nonparametric Bayesian credible sets”. Annals of Statistics, 43(4):1448–1454, August 2015. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/

1434546209. See [SvdVvZ15a].

Lecue:2018:RSB

[LM18] Guillaume Lecué and Shahar Mendelson. Regularization and the small-ball method I: Sparse recovery. Annals of Statistics, 46(2):611–641, April 2018. CODEN ASTSC7. ISSN

0090-5364. URL http://projecteuclid.org/euclid.aos/
1522742431.

Lugosi:2019:SGE

[LM19] Gábor Lugosi and Shahar Mendelson. Sub-Gaussian estimators of the mean of a random vector. Annals of Statistics, 47(2):783–794, April 2019. CODEN ASTSC7. ISSN

0090-5364. URL http://projecteuclid.org/euclid.aos/
1547197238.

Lee:2015:ASD

[LMNP15] Young K. Lee, Enno Mammen, Jens P. Nielsen, and Byeong U. Park. Asymptotics for in-sample density forecasting. Annals of Statistics, 43(2):620–651, April 2015. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1425398503.

Lee:2017:OTS

[LMNP17] Young K. Lee, Enno Mammen, Jens P. Nielsen, and Byeong U. Park. Operational time and in-sample density forecasting. Annals of Statistics, 45(3):1312–1341, June 2017. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1497319696.

Lee:2010:BSB

[LMP10] Young Kyung Lee, Enno Mammen, and Byeong U. Park. Backfitting and smooth backfitting for additive quantile models. Annals of Statistics, 38(5):2857–2883, October 2010. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1279638542.

Lee:2012:BSB

[LMP12a] Young K. Lee, Enno Mammen, and Byeong U. Park. Backfitting and smooth backfitting for additive quantile models. Annals of Statistics, 40(4):2356–2357, August 2012. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1358951385.

Lee:2012:FGV

[LMP12b] Young K. Lee, Enno Mammen, and Byeong U. Park. Flexible generalized varying coefficient regression models. Annals of Statistics, 40(3):1906–1933, June 2012. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1350394521.

Lounici:2018:ASN

[LMP18] Karim Lounici, Katia Meziani, and Gabriel Peyré. Adaptive sup-norm estimation of the Wigner function in noisy quantum homodyne tomography. Annals of Statistics, 46(3):1318– 1351, June 2018. CODEN ASTSC7. ISSN 0090-5364. URL

http://projecteuclid.org/euclid.aos/1525313084.

Lin:2018:MIP

[LMY18] Zhenhua Lin, Hans-Georg Müller, and Fang Yao. Mixture inner product spaces and their application to functional data analysis. Annals of Statistics, 46(1):370–400, February 2018. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

Lin:2019:ODN

[LMY19] Yi Lin, Ryan Martin, and Min Yang. On optimal designs for nonregular models. Annals of Statistics, 47(6):3335–3359, December 2019. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL https://projecteuclid.org/
euclid.aos/1572487395.

Lounici:2011:GUR

[LN11] Karim Lounici and Richard Nickl. Global uniform risk bounds for wavelet deconvolution estimators. Annals of Statistics, 39(1):201–231, February 2011. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1291388373.

Loh:2015:ESG

[Loh15] Wei-Liem Loh. Estimating the smoothness of a Gaussian random field from irregularly spaced data via higher-order quadratic variations. Annals of Statistics, 43(6):2766–2794, December 2015. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1444222092.

Loh:2017:SCA

[Loh17] Po-Ling Loh. Statistical consistency and asymptotic normality for high-dimensional robust M-estimators. Annals of Statistics, 45(2):866–896, April 2017. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1494921960.

Lok:2017:MCO

[Lok17] Judith J. Lok. Mimicking counterfactual outcomes to estimate causal effects. Annals of Statistics, 45(2):461–499, April 2017. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1494921947.

Lopes:2019:EAV

[Lop19] Miles E. Lopes. Estimating the algorithmic variance of randomized ensembles via the bootstrap. Annals of Statistics, 47(2):1088–1112, April 2019. CODEN ASTSC7. ISSN

0090-5364. URL http://projecteuclid.org/euclid.aos/
1547197249.

Leonenko:2010:CCR

[LP10] Nikolai Leonenko and Luc Pronzato. Correction: “A class of Rényi information estimators for multidimensional densities”. Annals of Statistics, 38(6):3837–3838, December 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/

1291126975. See [LPS08].

Lin:2017:ODF

[LPK17] Yuan-Lung Lin, Frederick Kin Hing Phoa, and Ming-Hung Kao. Optimal design of fMRI experiments using circulant (almost-)orthogonal arrays. Annals of Statistics, 45(6):2483– 2510, December 2017. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

Leonenko:2008:CRI

[LPS08] Nikolai Leonenko, Luc Pronzato, and Vippal Savani. A class of Rényi information estimators for multidimensional densities. Annals of Statistics, 36(5):2153–2182, October 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/

1223908088. See correction [LP10].

Lounici:2011:OIO

[LPvdGT11] Karim Lounici, Massimiliano Pontil, Sara van de Geer, and Alexandre B. Tsybakov. Oracle inequalities and optimal inference under group sparsity. Annals of Statistics, 39(4):2164–2204, August 2011. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1319595462.

Li:2012:RRC

[LPZZ12] Gaorong Li, Heng Peng, Jun Zhang, and Lixing Zhu. Robust rank correlation based screening. Annals of Statistics, 40(3):1846–1877, June 2012. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1350394519.

Lecue:2014:OLQ

[LR14] Guillaume Lecué and Philippe Rigollet. Optimal learning with Q-aggregation. Annals of Statistics, 42(1):211–224, February 2014. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1392733186.

Lei:2015:CSC

[LR15] Jing Lei and Alessandro Rinaldo. Consistency of spectral clustering in stochastic block models. Annals of Statistics, 43(1):215–237, February 2015. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1418135620.

Liu:2010:CTM

[LS10] Weidong Liu and Qi-Man Shao. Cramér-type moderate deviation for the maximum of the periodogram with application to simultaneous tests in gene expression time series. Annals of Statistics, 38(3):1913–1935, June 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1271271282.

Liu:2013:CMD

[LS13] Weidong Liu and Qi-Man Shao. A Cramér moderate deviation theorem for Hotelling’s T^{2}-statistic with applications to global tests. Annals of Statistics, 41(1):296–322, February 2013. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1364302744.

Lepski:2014:AEU

[LS14] Oleg Lepski and Nora Serdyukova. Adaptive estimation under single-index constraint in a regression model. Annals of Statistics, 42(1):1–28, February 2014. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1389795743.

Li:2017:NSD

[LS17] Bing Li and Jun Song. Nonlinear sufficient dimension reduction for functional data. Annals of Statistics, 45(3):1059– 1095, June 2017. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1497319688.

Lauritzen:2018:UMP

[LS18] Steffen Lauritzen and Kayvan Sadeghi. Unifying Markov properties for graphical models. Annals of Statistics, 46 (5):2251–2278, October 2018. CODEN ASTSC7. ISSN

0090-5364. URL http://projecteuclid.org/euclid.aos/
1534492835.

Liu:2017:SCD

[LSR^{+}17] Song Liu, Taiji Suzuki, Raissa Relator, Jun Sese, Masashi Sugiyama, and Kenji Fukumizu. Support consistency of direct sparse-change learning in Markov networks. Annals of Statistics, 45(3):959–990, June 2017. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1497319685.

Lee:2016:EPS

[LSST16] Jason D. Lee, Dennis L. Sun, Yuekai Sun, and Jonathan E. Taylor. Exact post-selection inference, with application to the lasso. Annals of Statistics, 44(3):907–927, June 2016. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1460381681.

Li:2016:ENR

[LTG16] Degui Li, Dag Tjøstheim, and Jiti Gao. Estimation in nonlinear regression with Harris recurrent Markov chains. Annals of Statistics, 44(5):1957–1987, October 2016. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1473685265.

Li:2013:VOT

[LTT13] Jia Li, Viktor Todorov, and George Tauchen. Volatility occupation times. Annals of Statistics, 41(4):1865–1891, August 2013. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1378386241.

Lockhart:2014:CRS

[LTTT14a] Richard Lockhart, Jonathan Taylor, Ryan J. Tibshirani, and Robert Tibshirani. Correction to rejoinder to “A significance test for the Lasso”. Annals of Statistics, 42(5):2138–2139, October 2014. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1413810450.

Lockhart:2014:RST

[LTTT14b] Richard Lockhart, Jonathan Taylor, Ryan J. Tibshirani, and Robert Tibshirani. Rejoinder: “A significance test for the lasso”. Annals of Statistics, 42(2):518–531, April 2014. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/

1400592168. See [LTTT14c].

Lockhart:2014:STL

[LTTT14c] Richard Lockhart, Jonathan Taylor, Ryan J. Tibshirani, and Robert Tibshirani. A significance test for the lasso. Annals of Statistics, 42(2):413–468, April 2014. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:/

/projecteuclid.org/euclid.aos/1400592161. See discus-

sionBuhlmann:2014:DST,Cai:2014:DST,Fan:2014:DST,Lv:2014:DST,Wasserman:2014:DST,Buja:2014:DST,Lockhart:2014:RST.

Lauritzen:2019:MLE

[LUZ19] Steffen Lauritzen, Caroline Uhler, and Piotr Zwiernik. Maximum likelihood estimation in Gaussian models under total positivity. Annals of Statistics, 47(4):1835–1863, August 2019. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/euclid.
aos/1558425632.

Lv:2013:IHD

[Lv13] Jinchi Lv. Impacts of high dimensionality in finite samples. Annals of Statistics, 41(4):2236–2262, August 2013. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1382547520.

Lei:2015:SAI

[LV15] Jing Lei and Vincent Q. Vu. Sparsistency and agnostic inference in sparse PCA. Annals of Statistics, 43(1):299–322, February 2015. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1423230081.

Luedtke:2016:SIM

[LvdL16] Alexander R. Luedtke and Mark J. van der Laan. Statistical inference for the mean outcome under a possibly non-unique optimal treatment strategy. Annals of Statistics, 44(2):713– 742, April 2016. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1458245733.

Liu:2010:SNI

[LW10] Weidong Liu and Wei Biao Wu. Simultaneous nonparametric inference of time series. Annals of Statistics, 38(4):2388–2421, August 2010. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1278861252.

Ledoit:2012:NSE

[LW12a] Olivier Ledoit and Michael Wolf. Nonlinear shrinkage estimation of large-dimensional covariance matrices. Annals of Statistics, 40(2):1024–1060, April 2012. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1342625460.

Loh:2012:HDR

[LW12b] Po-Ling Loh and Martin J. Wainwright. High-dimensional regression with noisy and missing data: Provable guarantees with nonconvexity. Annals of Statistics, 40(3):1637– 1664, June 2012. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1346850068.

Loh:2013:SED

[LW13] Po-Ling Loh and Martin J. Wainwright. Structure estimation for discrete graphical models: Generalized covariance matrices and their inverses. Annals of Statistics, 41(6):3022–3049, December 2013. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1388545677.

Loh:2017:SRI

[LW17] Po-Ling Loh and Martin J. Wainwright. Support recovery without incoherence: a case for nonconvex regularization. Annals of Statistics, 45(6):2455–2482, December 2017. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

Lepski:2019:OIA

[LW19] O. V. Lepski and T. Willer. Oracle inequalities and adaptive estimation in the convolution structure density model. Annals of Statistics, 47(1):233–287, February 2019. CODEN ASTSC7. ISSN 0090-5364. URL http://projecteuclid.

org/euclid.aos/1543568588.

Li:2017:INF

[LWY17] Zeng Li, Qinwen Wang, and Jianfeng Yao. Identifying the number of factors from singular values of a large sample auto-covariance matrix. Annals of Statistics, 45(1):257–288, February 2017. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1487667623.

Liu:2012:SAR

[LX12] Jingchen Liu and Gongjun Xu. Some asymptotic results of Gaussian random fields with varying mean functions and the associated processes. Annals of Statistics, 40(1):262–293, February 2012. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1333029965.

Lei:2016:ESC

[LXQ16] Huang Lei, Yingcun Xia, and Xu Qin. Estimation of semivarying coefficient time series models with ARMA errors. Annals of Statistics, 44(4):1618–1660, August 2016. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1467894710.

Liu:2011:PNP

[LY11] Wei Liu and Yuhong Yang. Parametric or nonparametric? A parametricness index for model selection. Annals of Statistics, 39(4):2074–2102, August 2011. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1319595459.

Lam:2012:FMH

[LY12] Clifford Lam and Qiwei Yao. Factor modeling for highdimensional time series: Inference for the number of factors. Annals of Statistics, 40(2):694–726, April 2012. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1337268209.

Lin:2019:IRF

[LY19] Zhenhua Lin and Fang Yao. Intrinsic Riemannian functional data analysis. Annals of Statistics, 47(6):3533–3577, December 2019. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL https://projecteuclid.org/
euclid.aos/1572487402.

Liu:2016:GSF

[LYL16] Hongcheng Liu, Tao Yao, and Runze Li. Global solutions to folded concave penalized nonconvex learning. Annals of Statistics, 44(2):629–659, April 2016. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1458245730.

Li:2011:EPL

[LYW^{+}11] Ting-Ting Li, Hu Yang, Jane-Ling Wang, Liu-Gen Xue, and Li-Xing Zhu. Estimation for a partial-linear singleindex model. Annals of Statistics, 39(6):3441–3443, December 2011. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1330958686.

Lerman:2011:RRM

[LZ11] Gilad Lerman and Teng Zhang. Robust recovery of multiple subspaces by geometric l_{p} minimization. Annals of Statistics, 39(5):2686–2715, October 2011. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1324563352.

Lv:2014:DST

[LZ14] Jinchi Lv and Zemin Zheng. Discussion: “A significance test for the lasso”. Annals of Statistics, 42(2):493–500, April 2014. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1400592165. See [LTTT14c].

Li:2017:ROP

[LZ17] Jun Li and Ping-Shou Zhong. A rate optimal procedure for recovering sparse differences between high-dimensional means under dependence. Annals of Statistics, 45(2):557– 590, April 2017. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1494921950.

Li:2015:ODP

[LZA15] Kang Li, Wei Zheng, and Mingyao Ai. Optimal designs for the proportional interference model. Annals of Statistics, 43(4):1596–1616, August 2015. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1434546216.

Lin:2018:CSS

[LZL18] Qian Lin, Zhigen Zhao, and Jun S. Liu. On consistency and sparsity for sliced inverse regression in high dimensions. Annals of Statistics, 46(2):580–610, April 2018. CODEN ASTSC7. ISSN 0090-5364. URL http://projecteuclid.

org/euclid.aos/1522742430.

Lian:2019:PSE

[LZL19] Heng Lian, Kaifeng Zhao, and Shaogao Lv. Projected spline estimation of the nonparametric function in high-dimensional partially linear models for massive data. Annals of Statistics, 47(5):2922–2949, October 2019. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://

projecteuclid.org/euclid.aos/1564797868.

Lee:2010:CPP

[LZW10] Seunggeun Lee, Fei Zou, and Fred A. Wright. Convergence and prediction of principal component scores in highdimensional settings. Annals of Statistics, 38(6):3605–3629, December 2010. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1291126967.

Ma:2012:TSS

[Ma12] Shujie Ma. Two-step spline estimating equations for generalized additive partially linear models with large cluster sizes. Annals of Statistics, 40(6):2943–2972, December 2012. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1360332189.

Ma:2013:SPC

[Ma13] Zongming Ma. Sparse principal component analysis and iterative thresholding. Annals of Statistics, 41(2):772–801, April 2013. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1368018173.

Masuda:2013:CGQ

[Mas13] Hiroki Masuda. Convergence of Gaussian quasi-likelihood random fields for ergodic Lévy driven SDE observed at high frequency. Annals of Statistics, 41(3):1593–1641, June 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1375362561.

Meinshausen:2015:MEI

[MB15] Nicolai Meinshausen and Peter Bühlmann. Maximin effects in inhomogeneous large-scale data. Annals of Statistics, 43(4):1801–1830, August 2015. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1434546223.

Mei:2018:LER

[MBM18] Song Mei, Yu Bai, and Andrea Montanari. The landscape of empirical risk for nonconvex losses. Annals of Statistics, 46 (6A):2747–2774, December 2018. CODEN ASTSC7. ISSN

0090-5364. URL http://projecteuclid.org/euclid.aos/
1536307232.

Maathuis:2015:GBD

[MC15] Marloes H. Maathuis and Diego Colombo. A generalized back-door criterion. Annals of Statistics, 43(3):1060–1088, June 2015. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1431695638.

Ma:2015:EIG

[MCLX15] Shujie Ma, Raymond J. Carroll, Hua Liang, and Shizhong Xu. Estimation and inference in generalized additive coefficient models for nonlinear interactions with high-dimensional covariates. Annals of Statistics, 43(5):2102–2131, October 2015. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1438606855.

Maire:2014:CAV

[MDO14] Florian Maire, Randal Douc, and Jimmy Olsson. Comparison of asymptotic variances of inhomogeneous Markov chains with application to Markov chain Monte Carlo methods. Annals of Statistics, 42(4):1483–1510, August 2014. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1407420006.

Meister:2011:AEF

[Mei11] Alexander Meister. Asymptotic equivalence of functional linear regression and a white noise inverse problem. Annals of Statistics, 39(3):1471–1495, June 2011. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1305292043.

Mendelson:2017:LVG

[Men17] Shahar Mendelson. “Local” vs. “global” parameters — breaking the Gaussian complexity barrier. Annals of Statistics, 45 (5):1835–1862, October 2017. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

Maruyama:2011:FBF

[MG11] Yuzo Maruyama and Edward I. George. Fully Bayes factors with a generalized g-prior. Annals of Statistics, 39 (5):2740–2765, October 2011. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1324563354.

Moreno:2010:COB

[MGC10] Elías Moreno, F. Javier Girón, and George Casella. Consistency of objective Bayes factors as the model dimension grows. Annals of Statistics, 38(4):1937–1952, August 2010. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1278861238.

McCullagh:2011:BTI

[MH11] Peter McCullagh and Han Han. On Bayes’s theorem for improper mixtures. Annals of Statistics, 39(4):2007–2020, August 2011. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1314190621.

Miller:2013:ESC

[MH13] Jeffrey W. Miller and Matthew T. Harrison. Exact sampling and counting for fixed-margin matrices. Annals of Statistics, 41(3):1569–1592, June 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1375362560.

McElroy:2014:ATC

[MH14] Tucker S. McElroy and Scott H. Holan. Asymptotic theory of cepstral random fields. Annals of Statistics, 42(1):64–86, February 2014. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1389795745.

Ma:2016:ISI

[MH16] Shujie Ma and Xuming He. Inference for single-index quantile regression models with profile optimization. Annals of Statistics, 44(3):1234–1268, June 2016. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1460381692.

Minsker:2018:SGE

[Min18] Stanislav Minsker. Sub-Gaussian estimators of the mean of a random matrix with heavy-tailed entries. Annals of Statistics, 46(6A):2871–2903, December 2018. CODEN ASTSC7. ISSN

0090-5364. URL http://projecteuclid.org/euclid.aos/
1536307236.

Mukherjee:2015:EME

[MJ15] Gourab Mukherjee and Iain M. Johnstone. Exact minimax estimation of the predictive density in sparse Gaussian models. Annals of Statistics, 43(3):937–961, June 2015. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1431695634.

Mak:2018:SP

[MJ18] Simon Mak and V. Roshan Joseph. Support points. Annals of Statistics, 46(6A):2562–2592, December 2018. CODEN ASTSC7. ISSN 0090-5364. URL http://projecteuclid.

org/euclid.aos/1536307226.

Molina:2018:EBP

[MM18] Isabel Molina and Nirian Martín. Empirical best prediction under a nested error model with log transformation. Annals of Statistics, 46(5):1961–1993, October 2018. CODEN ASTSC7.

ISSN 0090-5364. URL http://projecteuclid.org/euclid.
aos/1534492825.

Meinshausen:2011:AOW

[MMB11] Nicolai Meinshausen, Marloes H. Maathuis, and Peter Bühlmann. Asymptotic optimality of the Westfall–Young permutation procedure for multiple testing under dependence. Annals of Statistics, 39(6):3369–3391, December 2011. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1330958683.

Mousavi:2017:CPE

[MMB17] Ali Mousavi, Arian Maleki, and Richard G. Baraniuk. Consistent parameter estimation for LASSO and approximate message passing. Annals of Statistics, 45(6):2427–2454, December 2017. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

Mousavi:2018:CPE

[MMB18] Ali Mousavi, Arian Maleki, and Richard G. Baraniuk. Consistent parameter estimation for LASSO and approximate message passing. Annals of Statistics, 46(1):119–148, February 2018. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

McGoff:2015:CML

[MMNP15] Kevin McGoff, Sayan Mukherjee, Andrew Nobel, and Natesh Pillai. Consistency of maximum likelihood estimation for some dynamical systems. Annals of Statistics, 43(1):1–29, February 2015. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1416322034.

Mukherjee:2018:DTM

[MMS18] Rajarshi Mukherjee, Sumit Mukherjee, and Subhabrata Sen. Detection thresholds for the β-model on sparse graphs. Annals of Statistics, 46(3):1288–1317, June 2018. CODEN ASTSC7. ISSN 0090-5364. URL http://projecteuclid.

org/euclid.aos/1525313083.

Mukherjee:2018:GTA

[MMY18] Rajarshi Mukherjee, Sumit Mukherjee, and Ming Yuan. Global testing against sparse alternatives under Ising models. Annals of Statistics, 46(5):2062–2093, October 2018. CODEN ASTSC7. ISSN 0090-5364. URL http://projecteuclid.

org/euclid.aos/1534492829.

Mendelson:2010:RKL

[MN10] Shahar Mendelson and Joseph Neeman. Regularization in kernel learning. Annals of Statistics, 38(1):526–565, February 2010. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1262271623.

Monard:2019:ENB

[MNP19] François Monard, Richard Nickl, and Gabriel P. Paternain. Efficient nonparametric Bayesian inference for X-ray transforms. Annals of Statistics, 47(2):1113–1147, April 2019. CODEN ASTSC7. ISSN 0090-5364. URL http://

projecteuclid.org/euclid.aos/1547197250.

Metelkina:2017:IRC

[MP17] Asya Metelkina and Luc Pronzato. Information-regret compromise in covariate-adaptive treatment allocation. Annals of Statistics, 45(5):2046–2073, October 2017. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

Mukherjee:2015:HTH

[MPL15] Rajarshi Mukherjee, Natesh S. Pillai, and Xihong Lin. Hypothesis testing for high-dimensional sparse binary regression. Annals of Statistics, 43(1):352–381, February 2015. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1423230083.

Markaryan:2010:EPE

[MR10] Tigran Markaryan and William F. Rosenberger. Exact properties of Efron’s biased coin randomization procedure. Annals of Statistics, 38(3):1546–1567, June 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1269452646.

Morgan:2012:RIC

[MR12] Kari Lock Morgan and Donald B. Rubin. Rerandomization to improve covariate balance in experiments. Annals of Statistics, 40(2):1263–1282, April 2012. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1342625468.

Mammen:2012:NRN

[MRS12] Enno Mammen, Christoph Rothe, and Melanie Schienle. Nonparametric regression with nonparametrically generated covariates. Annals of Statistics, 40(2):1132–1170, April 2012. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1342625464.

McKeague:2010:FPI

[MS10a] Ian W. McKeague and Bodhisattva Sen. Fractals with point impact in functional linear regression. Annals of Statistics, 38(4):2559–2586, August 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1278861257.

Mijatovic:2010:GOP

[MS10b] Aleksandar Mijatović and Paul Schneider. Globally optimal parameter estimates for nonlinear diffusions. Annals of Statistics, 38(1):215–245, February 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1262271614.

Mukerjee:2013:CST

[MT13] Rahul Mukerjee and Boxin Tang. A complementary set theory for quaternary code designs. Annals of Statistics, 41 (6):2768–2785, December 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1387313389.

Mukherjee:2016:EEF

[Muk16] Sumit Mukherjee. Estimation in exponential families on permutations. Annals of Statistics, 44(2):853–875, April 2016. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1458245737.

Muller:2016:PHF

[Mül16] Hans-Georg Müller. Peter Hall, functional data analysis and random objects. Annals of Statistics, 44(5):1867–1887, October 2016. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1473685261.

Ma:2015:CBM

[MW15] Zongming Ma and Yihong Wu. Computational barriers in minimax submatrix detection. Annals of Statistics, 43(3):1089–1116, June 2015. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1431695639.

Massam:2016:NPD

[MW16] Hélène Massam and Jacek Wesolowski. A new prior for discrete DAG models with a restricted set of directions. Annals of Statistics, 44(3):1010–1037, June 2016. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1460381685.

Muller:2010:EDL

[MY10] Hans-Georg Müller and Fang Yao. Empirical dynamics for longitudinal data. Annals of Statistics, 38(6):3458–3486, December 2010. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1291126964.

Ma:2013:EES

[MZ13] Yanyuan Ma and Liping Zhu. Efficient estimation in sufficient dimension reduction. Annals of Statistics, 41(1):250–268, February 2013. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1364302742.

Mai:2015:FKF

[MZ15] Qing Mai and Hui Zou. The fused Kolmogorov filter: A nonparametric model-free screening method. Annals of Statistics, 43(4):1471–1497, August 2015. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1434546212.

Ma:2019:REA

[MZZ^{+}19] Shujie Ma, Liping Zhu, Zhiwei Zhang, Chih-Ling Tsai, and Raymond J. Carroll. A robust and efficient approach to causal inference based on sparse sufficient dimension reduction. Annals of Statistics, 47(3):1505–1535, June 2019. CODEN ASTSC7. ISSN 0090-5364. URL http://projecteuclid.

org/euclid.aos/1550026847.

Nadler:2016:DIF

[Nad16] Boaz Nadler. Discussion of “Influential features PCA for high dimensional clustering”. Annals of Statistics, 44(6): 2366–2371, December 2016. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1479891619. See [JW16a,

JW16b].

Nordman:2013:NEL

[NBL13] Daniel J. Nordman, Helle Bunzel, and Soumendra N. Lahiri. A nonstandard empirical likelihood for time series. Annals of Statistics, 41(6):3050–3073, December 2013. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1388545678.

Newton:2010:GBC

[NC10] Michael A. Newton and Lisa M. Chung. Gamma-based clustering via ordered means with application to gene-expression analysis. Annals of Statistics, 38(6):3217–3244, December 2010. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1284988405.

Nguyen:2013:CLM

[Ngu13] XuanLong Nguyen. Convergence of latent mixing measures in finite and infinite mixture models. Annals of Statistics, 41(1):370–400, February 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1364302747.

Narisetty:2014:BVS

[NH14] Naveen Naidu Narisetty and Xuming He. Bayesian variable selection with shrinking and diffusing priors. Annals of Statistics, 42(2):789–817, April 2014. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1400592178.

Nandy:2018:HDC

[NHM18] Preetam Nandy, Alain Hauser, and Marloes H. Maathuis. High-dimensional consistency in score-based and hybrid structure learning. Annals of Statistics, 46(6A):3151–3183, December 2018. CODEN ASTSC7. ISSN 0090-5364. URL

http://projecteuclid.org/euclid.aos/1536307246.

Nickl:2015:DFC

[Nic15] Richard Nickl. Discussion of “Frequentist coverage of adaptive nonparametric Bayesian credible sets”. Annals of Statistics, 43(4):1429–1436, August 2015. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1434546206. See

[SvdVvZ15a].

Naghshvar:2013:ASH

[NJ13] Mohammad Naghshvar and Tara Javidi. Active sequential hypothesis testing. Annals of Statistics, 41(6):2703–2738, December 2013. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1387313387.

Neumann:2018:MOC

[NK18] Christoph Neumann and Joachim Kunert. On MSE-optimal crossover designs. Annals of Statistics, 46(6A):2939–2959, December 2018. CODEN ASTSC7. ISSN 0090-5364. URL

http://projecteuclid.org/euclid.aos/1536307238.

Ning:2017:GTH

[NL17] Yang Ning and Han Liu. A general theory of hypothesis tests and confidence regions for sparse high dimensional models. Annals of Statistics, 45(1):158–195, February 2017. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1487667620.

Neykov:2019:PTH

[NL19] Matey Neykov and Han Liu. Property testing in highdimensional Ising models. Annals of Statistics, 47(5): 2472–2503, October 2019. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://

projecteuclid.org/euclid.aos/1564797854.

Neykov:2019:CIG

[NLL19] Matey Neykov, Junwei Lu, and Han Liu. Combinatorial inference for graphical models. Annals of Statistics, 47(2):795– 827, April 2019. CODEN ASTSC7. ISSN 0090-5364. URL

http://projecteuclid.org/euclid.aos/1547197239.

Nandy:2017:EEJ

[NMR17] Preetam Nandy, Marloes H. Maathuis, and Thomas S. Richardson. Estimating the effect of joint interventions from observational data in sparse high-dimensional settings. Annals of Statistics, 45(2):647–674, April 2017. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1494921953.

Norets:2010:ACD

[Nor10] Andriy Norets. Approximation of conditional densities by smooth mixtures of regressions. Annals of Statistics, 38(3):1733–1766, June 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1269452653.

Neuvial:2012:FDR

[NR12] Pierre Neuvial and Etienne Roquain. On false discovery rate thresholding for classification under sparsity. Annals of Statistics, 40(5):2572–2600, October 2012. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1359987531.

Nussbaum:2011:AEB

[NS11] Michael Nussbaum and Arleta Szkoła. An asymptotic error bound for testing multiple quantum hypotheses. Annals of Statistics, 39(6):3211–3233, December 2011. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1336568601.

Nickl:2017:NBP

[NS17] Richard Nickl and Jakob Söhl. Nonparametric Bayesian posterior contraction rates for discretely observed scalar diffusions. Annals of Statistics, 45(4):1664–1693, August 2017. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1498636870.

Nickl:2013:CSS

[NvdG13] Richard Nickl and Sara van de Geer. Confidence sets in sparse regression. Annals of Statistics, 41(6):2852–2876, December 2013. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1387313392.

Negahban:2011:ENL

[NW11] Sahand Negahban and Martin J. Wainwright. Estimation of (near) low-rank matrices with noise and high-dimensional scaling. Annals of Statistics, 39(2):1069–1097, April 2011. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1304947044.

Nye:2011:PCA

[Nye11] Tom M. W. Nye. Principal components analysis in the space of phylogenetic trees. Annals of Statistics, 39(5):2716–2739, October 2011. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1324563353.

Ning:2017:LRF

[NZL17] Yang Ning, Tianqi Zhao, and Han Liu. A likelihood ratio framework for high-dimensional semiparametric regression. Annals of Statistics, 45(6):2299–2327, December 2017. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

Olshen:2013:CSN

[OB13] Richard A. Olshen and Bala Brajaratnam. Correction: “Successive normalization of rectangular arrays”. Annals of Statistics, 41(5):2700–2702, October 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1386598637. See [OR10].

Onatski:2013:APS

[OMH13] Alexei Onatski, Marcelo J. Moreira, and Marc Hallin. Asymptotic power of sphericity tests for high-dimensional data. Annals of Statistics, 41(3):1204–1231, June 2013. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1371150898.

Onatski:2014:SDH

[OMH14] Alexei Onatski, Marcelo J. Moreira, and Marc Hallin. Signal detection in high dimension: The multispiked case. Annals of Statistics, 42(1):225–254, February 2014. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1395234977.

Overgaard:2017:ATG

[OPP17] Morten Overgaard, Erik Thorlund Parner, and Jan Pedersen. Asymptotic theory of generalized estimating equations based on jack-knife pseudo-observations. Annals of Statistics, 45(5): 1988–2015, October 2017. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

Olshen:2010:SNR

[OR10] Richard A. Olshen and Bala Rajaratnam. Successive normalization of rectangular arrays. Annals of Statistics, 38(3):1638– 1664, June 2010. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1269452650. See correction [OB13].

Obozinski:2011:SUR

[OWJ11] Guillaume Obozinski, Martin J. Wainwright, and Michael I. Jordan. Support union recovery in high-dimensional multivariate regression. Annals of Statistics, 39(1):1–47, February 2011. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1291388368.

Pomatto:2014:MTO

[PANS14] Luciano Pomatto, Nabil Al-Najjar, and Alvaro Sandroni. Merging and testing opinions. Annals of Statistics, 42 (3):1003–1028, June 2014. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1400592650.

Paparoditis:2018:SBF

[Pap18] Efstathios Paparoditis. Sieve bootstrap for functional time series. Annals of Statistics, 46(6B):3510–3538, December 2018. CODEN ASTSC7. ISSN 0090-5364. URL http://

projecteuclid.org/euclid.aos/1536631282.

Pati:2014:PCS

[PBPD14] Debdeep Pati, Anirban Bhattacharya, Natesh S. Pillai, and David Dunson. Posterior contraction in sparse Bayesian factor models for massive covariance matrices. Annals of Statistics, 42(3):1102–1130, June 2014. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1400592653.

Parry:2012:PLS

[PDL12] Matthew Parry, A. Philip Dawid, and Steffen Lauritzen. Proper local scoring rules. Annals of Statistics, 40 (1):561–592, February 2012. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1336396183.

Pensky:2013:SIL

[Pen13] Marianna Pensky. Spatially inhomogeneous linear inverse problems with possible singularities. Annals of Statistics, 41(5):2668–2697, October 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1384871349.

Pensky:2016:SLI

[Pen16] Marianna Pensky. Solution of linear ill-posed problems using overcomplete dictionaries. Annals of Statistics, 44 (4):1739–1764, August 2016. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1467894714.

Pensky:2017:MTE

[Pen17] Marianna Pensky. Minimax theory of estimation of linear functionals of the deconvolution density with or without sparsity. Annals of Statistics, 45(4):1516–1541, August 2017. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1498636865.

Pennec:2018:BSA

[Pen18] Xavier Pennec. Barycentric subspace analysis on manifolds. Annals of Statistics, 46(6A):2711–2746, December 2018. CODEN ASTSC7. ISSN 0090-5364. URL http://

projecteuclid.org/euclid.aos/1536307231.

Pensky:2019:DNM

[Pen19] Marianna Pensky. Dynamic network models and graphon estimation. Annals of Statistics, 47(4):2378–2403, August 2019. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/euclid.aos/
1558425649.

Phoa:2012:CAA

[Pho12] Frederick K. H. Phoa. A code arithmetic approach for quaternary code designs and its application to (1/64)thfractions. Annals of Statistics, 40(6):3161–3175, December 2012. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1361542078.

Pena:2011:PEM

[PHW11] Edsel A. Peña, Joshua D. Habiger, and Wensong Wu. Powerenhanced multiple decision functions controlling family-wise error and false discovery rates. Annals of Statistics, 39 (1):556–583, February 2011. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1297779856.

Park:2010:ADC

[PJS10] Byeong U. Park, Seok-Oh Jeong, and Léopold Simar. Asymptotic distribution of conical-hull estimators of directional edges. Annals of Statistics, 38(3):1320–1340, June 2010. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1268056618.

Papavasiliou:2011:PER

[PL11] Anastasia Papavasiliou and Christophe Ladroue. Parameter estimation for rough differential equations. Annals of Statistics, 39(4):2047–2073, August 2011. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1319595458.

Petersen:2016:FDA

[PM16a] Alexander Petersen and Hans-Georg Müller. Functional data analysis for density functions by transformation to a Hilbert space. Annals of Statistics, 44(1):183–218, February 2016. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1449755961.

Pillai:2016:UEC

[PM16b] Natesh S. Pillai and Xiao-Li Meng. An unexpected encounter with Cauchy and Lévy. Annals of Statistics, 44(5):2089–2097, October 2016. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1473685269.

Petersen:2019:FRR

[PM19] Alexander Petersen and Hans-Georg Müller. Fréchet regression for random objects with Euclidean predictors. Annals of Statistics, 47(2):691–719, April 2019. CODEN ASTSC7.

ISSN 0090-5364. URL http://projecteuclid.org/euclid.
aos/1547197235.

Palma:2010:EEL

[PO10] Wilfredo Palma and Ricardo Olea. An efficient estimator for locally stationary Gaussian long-memory processes. Annals of Statistics, 38(5):2958–2997, October 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1282315405.

Portnoy:2012:NRA

[Por12] Stephen Portnoy. Nearly root-n approximation for regression quantile processes. Annals of Statistics, 40(3):1714– 1736, June 2012. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1349196389.

Pazman:2014:ODA

[PP14] Andrej Pázman and Luc Pronzato. Optimum design accounting for the global nonlinear behavior of the model. Annals of Statistics, 42(4):1426–1451, August 2014. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1403715206.

Paul:2016:NED

[PPB16] Debashis Paul, Jie Peng, and Prabir Burman. Nonparametric estimation of dynamics of monotone trajectories. Annals of Statistics, 44(6):2401–2432, December 2016. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1479891623.

Plamadeala:2012:SMC

[PR12] Victoria Plamadeala and William F. Rosenberger. Sequential monitoring with conditional randomization tests. Annals of Statistics, 40(1):30–44, February 2012. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1331830773.

Perchet:2013:MAB

[PR13] Vianney Perchet and Philippe Rigollet. The multi-armed bandit problem with covariates. Annals of Statistics, 41(2):693–721, April 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1366980562.

Patschkowski:2016:ALD

[PR16] Tim Patschkowski and Angelika Rohde. Adaptation to lowest density regions with application to support recovery. Annals of Statistics, 44(1):255–287, February 2016. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1449755963.

Patschkowski:2019:LAC

[PR19] Tim Patschkowski and Angelika Rohde. Locally adaptive confidence bands. Annals of Statistics, 47(1):349–381, February 2019. CODEN ASTSC7. ISSN 0090-5364. URL http://

projecteuclid.org/euclid.aos/1543568591.

Perchet:2016:BBP

[PRCS16] Vianney Perchet, Philippe Rigollet, Sylvain Chassang, and Erik Snowberg. Batched bandit problems. Annals of Statistics, 44(2):660–681, April 2016. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1458245731.

Pensky:2010:CRE

[PS10] Marianna Pensky and Theofanis Sapatinas. On convergence rates equivalency and sampling strategies in functional deconvolution models. Annals of Statistics, 38(3):1793–1844, June 2010. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1269452655.

Pretorius:2018:ATN

[PS18] Charl Pretorius and Jan W. H. Swanepoel. On the asymptotic theory of new bootstrap confidence bounds. Annals of Statistics, 46(1):438–456, February 2018. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

Polunchenko:2010:OSR

[PT10] Aleksey S. Polunchenko and Alexander G. Tartakovsky. On optimality of the Shiryaev–Roberts procedure for detecting a change in distribution. Annals of Statistics, 38(6):3445–3457, December 2010. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1291126963.

Panaretos:2013:FAS

[PT13] Victor M. Panaretos and Shahin Tavakoli. Fourier analysis of stationary time series in function space. Annals of Statistics, 41(2):568–603, April 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1366980558.

Pan:2018:BDN

[PTWZ18] Wenliang Pan, Yuan Tian, Xueqin Wang, and Heping Zhang. Ball divergence: Nonparametric two sample test. Annals of Statistics, 46(3):1109–1137, June 2018. CODEN ASTSC7.

ISSN 0090-5364. URL http://projecteuclid.org/euclid.
aos/1525313077.

Paindaveine:2017:IMW

[PV17] Davy Paindaveine and Thomas Verdebout. Inference on the mode of weak directional signals: A Le Cam perspective on hypothesis testing near singularities. Annals of Statistics, 45(2):800–832, April 2017. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1494921958.

Paindaveine:2018:HDS

[PV18] Davy Paindaveine and Germain Van Bever. Halfspace depths for scatter, concentration and shape matrices. Annals of Statistics, 46(6B):3276–3307, December 2018. CODEN ASTSC7. ISSN 0090-5364. URL http://projecteuclid.

org/euclid.aos/1536631274.

Perry:2018:OSO

[PWBM18] Amelia Perry, Alexander S. Wein, Afonso S. Bandeira, and Ankur Moitra. Optimality and sub-optimality of PCA I: Spiked random matrix models. Annals of Statistics, 46 (5):2416–2451, October 2018. CODEN ASTSC7. ISSN

0090-5364. URL http://projecteuclid.org/euclid.aos/
1534492840.

Proksch:2018:MSI

[PWM18] Katharina Proksch, Frank Werner, and Axel Munk. Multiscale scanning in inverse problems. Annals of Statistics, 46 (6B):3569–3602, December 2018. CODEN ASTSC7. ISSN

0090-5364. URL http://projecteuclid.org/euclid.aos/
1536631284.

Pillai:2012:EUC

[PY12] Natesh S. Pillai and Jun Yin. Edge universality of correlation matrices. Annals of Statistics, 40(3):1737–1763, June 2012. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1349196390.

Panaretos:2016:APV

[PZ16] Victor M. Panaretos and Yoav Zemel. Amplitude and phase variation of point processes. Annals of Statistics, 44(2):771– 812, April 2016. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1458245735.

Qiu:2012:TBH

[QC12] Yumou Qiu and Song Xi Chen. Test for bandedness of high-dimensional covariance matrices and bandwidth estimation. Annals of Statistics, 40(3):1285–1314, June 2012. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1344610584.

Qiu:2018:DRF

[QCN18] Yumou Qiu, Song Xi Chen, and Dan Nettleton. Detecting rare and faint signals via thresholding maximum likelihood estimators. Annals of Statistics, 46(2):895–923, April 2018. CODEN ASTSC7. ISSN 0090-5364. URL http://

projecteuclid.org/euclid.aos/1522742440.

Qin:2019:CCA

[QH19] Qian Qin and James P. Hobert. Convergence complexity analysis of Albert and Chib’s algorithm for Bayesian probit regression. Annals of Statistics, 47(4):2320–2347, August 2019. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/euclid.aos/
1558425647.

Qian:2011:PGI

[QM11] Min Qian and Susan A. Murphy. Performance guarantees for individualized treatment rules. Annals of Statistics, 39(2):1180–1210, April 2011. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1304947047.

Qiao:2016:TAN

[QP16] Wanli Qiao and Wolfgang Polonik. Theoretical analysis of nonparametric filament estimation. Annals of Statistics, 44(3):1269–1297, June 2016. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1460381693.

Qin:2019:NIL

[QT19] Likuan Qin and Viktor Todorov. Nonparametric implied Lévy densities. Annals of Statistics, 47(2):1025–1060, April 2019. CODEN ASTSC7. ISSN 0090-5364. URL http://

projecteuclid.org/euclid.aos/1547197247.

Qu:2016:OEF

[QWW16] Simeng Qu, Jane-Ling Wang, and Xiao Wang. Optimal estimation for the functional Cox model. Annals of Statistics, 44(4):1708–1738, August 2016. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1467894713.

Qi:2010:AEF

[QZ10] Xin Qi and Hongyu Zhao. Asymptotic efficiency and finitesample properties of the generalized profiling estimation of parameters in ordinary differential equations. Annals of Statistics, 38(1):435–481, February 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1262271620.

Rao:2018:SIS

[Rao18] Suhasini Subba Rao. Statistical inference for spatial statistics defined in the Fourier domain. Annals of Statistics, 46(2):469–499, April 2018. CODEN ASTSC7. ISSN

0090-5364. URL http://projecteuclid.org/euclid.aos/
1522742426.

Ray:2017:ABM

[Ray17] Kolyan Ray. Adaptive Bernstein–von Mises theorems in Gaussian white noise. Annals of Statistics, 45(6):2511–2536, December 2017. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

Rothenhausler:2019:CDF

[RBM19] Dominik Rothenhäusler, Peter Bühlmann, and Nicolai Meinshausen. Causal Dantzig: Fast inference in linear structural equation models with hidden variables under additive interventions. Annals of Statistics, 47(3):1688–1722, June 2019. CODEN ASTSC7. ISSN 0090-5364. URL http://

projecteuclid.org/euclid.aos/1550026854.

Reynaud-Bouret:2010:AEH

[RBS10] Patricia Reynaud-Bouret and Sophie Schbath. Adaptive estimation for Hawkes processes; application to genome analysis. Annals of Statistics, 38(5):2781–2822, October 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1279638540.

Ramdas:2019:UTM

[RBWJ19] Aaditya K. Ramdas, Rina F. Barber, Martin J. Wainwright, and Michael I. Jordan. A unified treatment of multiple testing with prior knowledge using the p-filter. Annals of Statistics, 47(5):2790–2821, October 2019. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://

projecteuclid.org/euclid.aos/1564797864.

Rousseau:2012:BNE

[RCL12] Judith Rousseau, Nicolas Chopin, and Brunero Liseo. Bayesian nonparametric estimation of the spectral density of a long or intermediate memory Gaussian process. Annals of Statistics, 40(2):964–995, April 2012. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1342625458.

Rohe:2011:SCH

[RCY11] Karl Rohe, Sourav Chatterjee, and Bin Yu. Spectral clustering and the high-dimensional stochastic blockmodel. Annals of Statistics, 39(4):1878–1915, August 2011. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1314190618.

Richter:2019:CVL

[RD19] Stefan Richter and Rainer Dahlhaus. Cross validation for locally stationary processes. Annals of Statistics, 47 (4):2145–2173, August 2019. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://

projecteuclid.org/euclid.aos/1558512018.

Rothenhausler:2018:CIP

[REB18] Dominik Rothenhäusler, Jan Ernest, and Peter Bühlmann. Causal inference in partially linear structural equation models. Annals of Statistics, 46(6A):2904–2938, December 2018. CODEN ASTSC7. ISSN 0090-5364. URL http://

projecteuclid.org/euclid.aos/1536307237.

Reiss:2011:AEI

[Rei11] Markus Reiß. Asymptotic equivalence for inference on the volatility from noisy observations. Annals of Statistics, 39(2):772–802, April 2011. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1299680954.

Roman:2012:CAG

[RH12] Jorge Carlos Román and James P. Hobert. Convergence analysis of the Gibbs sampler for Bayesian general linear mixed models with improper priors. Annals of Statistics, 40(6):2823–2849, December 2012. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1360332185.

Rigollet:2012:KLA

[Rig12] Philippe Rigollet. Kullback–Leibler aggregation and misspecified generalized linear models. Annals of Statistics, 40(2):639–665, April 2012. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1337268207.

Robins:2017:MEF

[RLM^{+}17] James M. Robins, Lingling Li, Rajarshi Mukherjee, Eric Tchetgen Tchetgen, and Aad van der Vaart. Minimax estimation of a functional on a structured high-dimensional model. Annals of Statistics, 45(5):1951–1987, October 2017. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

Rockova:2018:BES

[Roč18] Veronika Ročková. Bayesian estimation of sparse signals with a continuous spike-and-slab prior. Annals of Statistics, 46(1): 401–437, February 2018. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

Rohe:2019:CTD

[Roh19] Karl Rohe. A critical threshold for design effects in network sampling. Annals of Statistics, 47(1):556–582, February 2019. CODEN ASTSC7. ISSN 0090-5364. URL http://

projecteuclid.org/euclid.aos/1543568598.

Rojo:2011:ELL

[Roj11] Javier Rojo. Erich Leo Lehmann — a glimpse into his life and work. Annals of Statistics, 39(5):2244–2265, October 2011. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1321020523.

Rousseau:2010:RCP

[Rou10] Judith Rousseau. Rates of convergence for the posterior distributions of mixtures of betas and adaptive nonparametric estimation of the density. Annals of Statistics, 38(1):146–180, February 2010. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1262271612.

Rousseau:2015:DFC

[Rou15] Judith Rousseau. Discussion of “Frequentist coverage of adaptive nonparametric Bayesian credible sets”. Annals of Statistics, 43(4):1444–1447, August 2015. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1434546208. See

[SvdVvZ15a].

Roysland:2012:CAG

[Røy12] Kjetil Røysland. Counterfactual analyses with graphical models based on local independence. Annals of Statistics, 40(4):2162–2194, August 2012. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1358951379.

Rinaldo:2013:MLE

[RPF13] Alessandro Rinaldo, Sonja Petrović, and Stephen E. Fienberg. Maximum lilkelihood estimation in the β-model. Annals of Statistics, 41(3):1085–1110, June 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1371150894.

Rivoirard:2012:BMT

[RR12] Vincent Rivoirard and Judith Rousseau. Bernstein–von Mises theorem for linear functionals of the density. Annals of Statistics, 40(3):1489–1523, June 2012. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1346850063.

Romano:2012:UAV

[RS12] Joseph P. Romano and Azeem M. Shaikh. On the uniform asymptotic validity of subsampling and the bootstrap. Annals of Statistics, 40(6):2798–2822, December 2012. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1360332184.

Rousseau:2017:ABE

[RS17] Judith Rousseau and Botond Szabo. Asymptotic behaviour of the empirical Bayes posteriors associated to maximum marginal likelihood estimator. Annals of Statistics, 45(2):833–865, April 2017. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1494921959.

Ren:2015:ANO

[RSZZ15] Zhao Ren, Tingni Sun, Cun-Hui Zhang, and Harrison H. Zhou. Asymptotic normality and optimalities in estimation of large Gaussian graphical models. Annals of Statistics, 43(3):991–1026, June 2015. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1431695636.

Rosenbaum:2010:SRU

[RT10] Mathieu Rosenbaum and Alexandre B. Tsybakov. Sparse recovery under matrix uncertainty. Annals of Statistics, 38(5):2620–2651, October 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1278861455.

Rigollet:2011:ESO

[RT11a] Philippe Rigollet and Alexandre Tsybakov. Exponential screening and optimal rates of sparse estimation. Annals of Statistics, 39(2):731–771, April 2011. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1299680953.

Rohde:2011:EHD

[RT11b] Angelika Rohde and Alexandre B. Tsybakov. Estimation of high-dimensional low-rank matrices. Annals of Statistics, 39(2):887–930, April 2011. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1299680958.

Roquain:2011:ECF

[RV11] Etienne Roquain and Fanny Villers. Exact calculations for false discovery proportion with application to least favorable configurations. Annals of Statistics, 39(1):584–612, February 2011. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1297779857.

Rinaldo:2010:GDC

[RW10a] Alessandro Rinaldo and Larry Wasserman. Generalized density clustering. Annals of Statistics, 38(5):2678–2722, October 2010. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1278861457.

Romano:2010:BCG

[RW10b] Joseph P. Romano and Michael Wolf. Balanced control of generalized error rates. Annals of Statistics, 38(1):598–633, February 2010. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1262271625.

Rinaldo:2019:BSS

[RWG19] Alessandro Rinaldo, Larry Wasserman, and Max G’Sell. Bootstrapping and sample splitting for high-dimensional, assumption-lean inference. Annals of Statistics, 47(6): 3438–3469, December 2019. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://

projecteuclid.org/euclid.aos/1572487399.

Ravikumar:2010:HDI

[RWL10] Pradeep Ravikumar, Martin J. Wainwright, and John D. Lafferty. High-dimensional Ising model selection using ℓ_{1}- regularized logistic regression. Annals of Statistics, 38 (3):1287–1319, June 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1268056617.

Raskutti:2019:CRH

[RYC19] Garvesh Raskutti, Ming Yuan, and Han Chen. Convex regularization for high-dimensional multiresponse tensor regression. Annals of Statistics, 47(3):1554–1584, June 2019. CODEN ASTSC7. ISSN 0090-5364. URL http://

projecteuclid.org/euclid.aos/1550026849.

Ren:2012:DLV

[RZ12] Zhao Ren and Harrison H. Zhou. Discussion: Latent variable graphical model selection via convex optimization. Annals of Statistics, 40(4):1989–1996, August 2012. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1351602532.

Sadeghi:2016:MCL

[Sad16] Kayvan Sadeghi. Marginalization and conditioning for LWF chain graphs. Annals of Statistics, 44(4):1792–1816, August 2016. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1467894716.

Saegusa:2019:LST

[Sae19] Takumi Saegusa. Large sample theory for merged data from multiple sources. Annals of Statistics, 47(3):1585–1615, June 2019. CODEN ASTSC7. ISSN 0090-5364. URL http://

projecteuclid.org/euclid.aos/1550026850.

Samworth:2012:OWN

[Sam12] Richard J. Samworth. Optimal weighted nearest neighbour classifiers. Annals of Statistics, 40(5):2733–2763, October 2012. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1359987536.

Samworth:2016:PHW

[Sam16] Richard J. Samworth. Peter Hall’s work on high-dimensional data and classification. Annals of Statistics, 44(5):1888–1895, October 2016. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1473685262.

Scott:2010:BEB

[SB10] James G. Scott and James O. Berger. Bayes and empirical-Bayes multiplicity adjustment in the variable-selection problem. Annals of Statistics, 38(5):2587–2619, October 2010.

CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1278861454.

Sarkar:2015:RNS

[SB15] Purnamrita Sarkar and Peter J. Bickel. Role of normalization in spectral clustering for stochastic blockmodels. Annals of Statistics, 43(3):962–990, June 2015. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1431695635.

Su:2017:FDO

[SBC17] Weijie Su, Małgorzata Bogdan, and Emmanuel Candès. False discoveries occur early on the Lasso path. Annals of Statistics, 45(5):2133–2150, October 2017. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

Song:2016:ACP

[SBK16] Rui Song, Moulinath Banerjee, and Michael R. Kosorok. Asymptotics for change-point models under varying degrees of mis-specification. Annals of Statistics, 44(1):153–182, February 2016. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1449755960.

Scornet:2015:CRF

[SBV15] Erwan Scornet, Gérard Biau, and Jean-Philippe Vert. Consistency of random forests. Annals of Statistics, 43 (4):1716–1741, August 2015. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1434546220.

Sen:2010:IBG

[SBW10] Bodhisattva Sen, Moulinath Banerjee, and Michael Woodroofe. Inconsistency of bootstrap: The Grenander estimator. Annals of Statistics, 38(4):1953–1977, August 2010. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1278861239.

Soltanolkotabi:2012:GAS

[SC12] Mahdi Soltanolkotabi and Emmanuel J. Candés. A geometric analysis of subspace clustering with outliers. Annals of Statistics, 40(4):2195–2238, August 2012. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1358951380.

Shang:2013:LGA

[SC13] Zuofeng Shang and Guang Cheng. Local and global asymptotic inference in smoothing spline models. Annals of Statistics, 41(5):2608–2638, October 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1384871347.

Shang:2015:NIG

[SC15] Zuofeng Shang and Guang Cheng. Nonparametric inference in generalized functional linear models. Annals of Statistics, 43(4):1742–1773, August 2015. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1434546221.

Su:2016:SAU

[SC16] Weijie Su and Emmanuel Candès. SLOPE is adaptive to unknown sparsity and asymptotically minimax. Annals of Statistics, 44(3):1038–1068, June 2016. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1460381686.

Schennach:2013:RBE

[Sch13] Susanne M. Schennach. Regressions with Berkson errors in covariates — a nonparametric approach. Annals of Statistics, 41(3):1642–1668, June 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1375362562.

Shao:2012:EHD

[SD12] Jun Shao and Xinwei Deng. Estimation in high-dimensional linear models with deterministic design matrices. Annals of Statistics, 40(2):812–831, April 2012. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1337268213.

Soltanolkotabi:2014:RSC

[SEC14] Mahdi Soltanolkotabi, Ehsan Elhamifar, and Emmanuel J. Candès. Robust subspace clustering. Annals of Statistics, 42(2):669–699, April 2014. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1400592174.

Schoen:2019:CCD

[SEG19] Eric D. Schoen, Pieter T. Eendebak, and Peter Goos. A classification criterion for definitive screening designs. Annals of Statistics, 47(2):1179–1202, April 2019. CODEN ASTSC7.

ISSN 0090-5364. URL http://projecteuclid.org/euclid.
aos/1547197252.

Song:2019:SMT

[SF19] Yanglei Song and Georgios Fellouris. Sequential multiple testing with generalized error control: An asymptotic optimality theory. Annals of Statistics, 47(3):1776–1803, June 2019. CODEN ASTSC7. ISSN 0090-5364. URL http://

projecteuclid.org/euclid.aos/1550026857.

Shi:2018:HDL

[SFSL18] Chengchun Shi, Ailin Fan, Rui Song, and Wenbin Lu. High-dimensional A-learning for optimal dynamic treatment regimes. Annals of Statistics, 46(3):925–957, June 2018. CODEN ASTSC7. ISSN 0090-5364. URL http://

projecteuclid.org/euclid.aos/1525313071.

Schwartzman:2011:MTL

[SGA11] Armin Schwartzman, Yulia Gavrilov, and Robert J. Adler. Multiple testing of local maxima for detection of peaks in 1d. Annals of Statistics, 39(6):3290–3319, December 2011. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1330958680.

Schmidt-Hieber:2014:AER

[SH14] Johannes Schmidt-Hieber. Asymptotic equivalence for regression under fractional noise. Annals of Statistics, 42 (6):2557–2585, December 2014. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1415801783.

Sagnol:2015:CEO

[SH15] Guillaume Sagnol and Radoslav Harman. Computing exact D-optimal designs by mixed integer second-order cone programming. Annals of Statistics, 43(5):2198–2224, October 2015. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1442364150.

Schmidt-Hieber:2013:MMS

[SHMD13] Johannes Schmidt-Hieber, Axel Munk, and Lutz Dümbgen. Multiscale methods for shape constraints in deconvolution: Confidence statements for qualitative features. Annals of Statistics, 41(3):1299–1328, June 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1372979639.

Steinberger:2019:PWF

[SL19] Lukas Steinberger and Hannes Leeb. Prediction when fitting simple models to high-dimensional data. Annals of Statistics, 47(3):1408–1442, June 2019. CODEN ASTSC7. ISSN

0090-5364. URL http://projecteuclid.org/euclid.aos/
1550026843.

Sun:2011:COM

[SLL11] Fasheng Sun, Dennis K. J. Lin, and Min-Qian Liu. On construction of optimal mixed-level supersaturated designs. Annals of Statistics, 39(2):1310–1333, April 2011. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1304947052.

Sun:2014:CNS

[SLQ14] Fasheng Sun, Min-Qian Liu, and Peter Z. G. Qian. On the construction of nested space-filling designs. Annals of Statistics, 42(4):1394–1425, August 2014. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1403715205.

Shu:2019:ELC

[SN19] Hai Shu and Bin Nan. Estimation of large covariance and precision matrices from temporally dependent observations. Annals of Statistics, 47(3):1321–1350, June 2019. CODEN ASTSC7. ISSN 0090-5364. URL http://projecteuclid.

org/euclid.aos/1550026839.

Seo:2018:LED

[SO18] Myung Hwan Seo and Taisuke Otsu. Local M-estimation with discontinuous criterion for dependent and limited observations. Annals of Statistics, 46(1):344–369, February 2018. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

Spokoiny:2012:PEF

[Spo12] Vladimir Spokoiny. Parametric estimation. Finite sample theory. Annals of Statistics, 40(6):2877–2909, December 2012. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1360332187.

Shi:2017:WSI

[SQ17] Peibei Shi and Annie Qu. Weak signal identification and inference in penalized model selection. Annals of Statistics, 45(3):1214–1253, June 2017. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1497319693.

Shalizi:2013:CUS

[SR13] Cosma Rohilla Shalizi and Alessandro Rinaldo. Consistency under sampling of exponential random graph models. Annals of Statistics, 41(2):508–535, April 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1366980556.

Szekely:2014:PDC

[SR14] Gábor J. Székely and Maria L. Rizzo. Partial distance correlation with methods for dissimilarities. Annals of Statistics, 42(6):2382–2412, December 2014. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1413810731.

Samonenko:2015:NPT

[SR15] Inga Samonenko and John Robinson. A new permutation test statistic for complete block designs. Annals of Statistics, 43(1):90–101, February 2015. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1416322037.

Seijo:2011:CPS

[SS11a] Emilio Seijo and Bodhisattva Sen. Change-point in stochastic design regression and the bootstrap. Annals of Statistics, 39(3):1580–1607, June 2011. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1307452129.

Seijo:2011:NLS

[SS11b] Emilio Seijo and Bodhisattva Sen. Nonparametric least squares estimation of a multivariate convex regression function. Annals of Statistics, 39(3):1633–1657, June 2011. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1311600278.

Suzuki:2013:FLR

[SS13] Taiji Suzuki and Masashi Sugiyama. Fast learning rate of multiple kernel learning: Trade-off between sparsity and smoothness. Annals of Statistics, 41(3):1381–1405, June 2013. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1375362553.

Shi:2019:LHT

[SSCL19] Chengchun Shi, Rui Song, Zhao Chen, and Runze Li. Linear hypothesis testing for high dimensional generalized linear models. Annals of Statistics, 47(5):2671–2703, October 2019. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/euclid.aos/
1564797860.

Sejdinovic:2013:EDB

[SSGF13] Dino Sejdinovic, Bharath Sriperumbudur, Arthur Gretton, and Kenji Fukumizu. Equivalence of distance-based and RKHS-based statistics in hypothesis testing. Annals of Statistics, 41(5):2263–2291, October 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1383661264.

Schervish:2014:DCM

[SSK14] M. J. Schervish, Teddy Seidenfeld, and J. B. Kadane. Dominating countably many forecasts. Annals of Statistics, 42(2):728–756, April 2014. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1400592176.

Shi:2019:TCQ

[SSL19] Chengchun Shi, Rui Song, and Wenbin Lu. On testing conditional qualitative treatment effects. Annals of Statistics, 47(4):2348–2377, August 2019. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://

projecteuclid.org/euclid.aos/1558425648.

Shpitser:2016:CIG

[ST16a] Ilya Shpitser and Eric Tchetgen Tchetgen. Causal inference with a graphical hierarchy of interventions. Annals of Statistics, 44(6):2433–2466, December 2016. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1479891624.

Stepanova:2016:DIF

[ST16b] Natalia A. Stepanova and Alexandre B. Tsybakov. Discussion of “Influential features PCA for high dimensional clustering”. Annals of Statistics, 44(6):2382–2386, December 2016. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1479891621. See [JW16a, JW16b].

Sadhanala:2019:AMT

[ST19] Veeranjaneyulu Sadhanala and Ryan J. Tibshirani. Additive models with trend filtering. Annals of Statistics, 47 (6):3032–3068, December 2019. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://

projecteuclid.org/euclid.aos/1572487382.

Sullivant:2010:TSG

[STD10] Seth Sullivant, Kelli Talaska, and Jan Draisma. Trek separation for Gaussian graphical models. Annals of Statistics, 38(3):1665–1685, June 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1269452651.

Stein:2011:RLW

[Ste11] Michael L. Stein. 2010 Rietz Lecture: When does the screening effect hold? Annals of Statistics, 39(6):2795–2819, December 2011. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1327413769.

Steinwart:2015:FAD

[Ste15] Ingo Steinwart. Fully adaptive density-based clustering. Annals of Statistics, 43(5):2132–2167, October 2015. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1442364148.

Strauch:2018:AID

[Str18] Claudia Strauch. Adaptive invariant density estimation for ergodic diffusions over anisotropic classes. Annals of Statistics, 46(6B):3451–3480, December 2018. CODEN ASTSC7.

ISSN 0090-5364. URL http://projecteuclid.org/euclid.
aos/1536631280.

Sherlock:2015:EPM

[STRR15] Chris Sherlock, Alexandre H. Thiery, Gareth O. Roberts, and Jeffrey S. Rosenthal. On the efficiency of pseudo-marginal random walk Metropolis algorithms. Annals of Statistics, 43(1):238–275, February 2015. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1418135621.

Shi:2014:OML

[STY14] Ce Shi, Yu Tang, and Jianxing Yin. Optimum mixed level detecting arrays. Annals of Statistics, 42(4):1546–1563, August 2014. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1407420008.

Segers:2014:SGC

[SvdAW14] Johan Segers, Ramon van den Akker, and Bas J. M. Werker. Semiparametric Gaussian copula models: Geometry and efficient rank-based estimation. Annals of Statistics, 42 (5):1911–1940, October 2014. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1410440629.

Szabo:2015:FCA

[SvdVvZ15a] Botond Szabó, A. W. van der Vaart, and J. H. van Zanten. Frequentist coverage of adaptive nonparametric Bayesian credible sets. Annals of Statistics, 43(4):1391–1428, August 2015. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1434546205. See discussion [Nic15, Cas15, Rou15, LM15, Gho15] and rejoinder [SvdVvZ15b].

Szabo:2015:RDF

[SvdVvZ15b] Botond Szabó, A. W. van der Vaart, and J. H. van Zanten. Rejoinder to discussions of “Frequentist coverage of adaptive nonparametric Bayesian credible sets”. Annals of Statistics, 43(4):1463–1470, August 2015. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/

1434546211. See [SvdVvZ15a].

Samworth:2010:AOB

[SW10a] R. J. Samworth and M. P. Wand. Asymptotics and optimal bandwidth selection for highest density region estimation. Annals of Statistics, 38(3):1767–1792, June 2010. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1269452654.

Seregin:2010:NEM

[SW10b] Arseni Seregin and Jon A. Wellner. Nonparametric estimation of multivariate convex-transformed densities. Annals of Statistics, 38(6):3751–3781, December 2010. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1291126972.

Su:2012:MLT

[SW12] Yu-Ru Su and Jane-Ling Wang. Modeling left-truncated and right-censored survival data with longitudinal covariates. Annals of Statistics, 40(3):1465–1488, June 2012. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1346850062.

Saegusa:2013:WLE

[SW13] Takumi Saegusa and Jon A. Wellner. Weighted likelihood estimation under two-phase sampling. Annals of Statistics, 41(1):269–295, February 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1364302743.

Sherwood:2016:PLA

[SW16a] Ben Sherwood and Lan Wang. Partially linear additive quantile regression in ultra-high dimension. Annals of Statistics, 44(1):288–317, February 2016. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1449755964.

Su:2016:SEE

[SW16b] Yu-Ru Su and Jane-Ling Wang. Semiparametric efficient estimation for shared-frailty models with doubly-censored clustered data. Annals of Statistics, 44(3):1298–1331, June 2016. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1460381694.

Sienkiewicz:2018:PQU

[SW18] Ela Sienkiewicz and Haonan Wang. Pareto quantiles of unlabeled tree objects. Annals of Statistics, 46(4):1513–1540, August 2018. CODEN ASTSC7. ISSN 0090-5364. URL

http://projecteuclid.org/euclid.aos/1530086424.

Spokoiny:2019:BTG

[SW19] Vladimir Spokoiny and Niklas Willrich. Bootstrap tuning in Gaussian ordered model selection. Annals of Statistics, 47(3):1351–1380, June 2019. CODEN ASTSC7. ISSN

0090-5364. URL http://projecteuclid.org/euclid.aos/
1550026841.

Shao:2011:SLD

[SWDW11] Jun Shao, Yazhen Wang, Xinwei Deng, and Sijian Wang. Sparse linear discriminant analysis by thresholding for high dimensional data. Annals of Statistics, 39(2):1241–1265, April 2011. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1304947049.

Sun:2019:UPD

[SWX19] Fasheng Sun, Yaping Wang, and Hongquan Xu. Uniform projection designs. Annals of Statistics, 47(1):641–661, February 2019. CODEN ASTSC7. ISSN 0090-5364. URL http://

projecteuclid.org/euclid.aos/1543568601.

Schiebinger:2015:GKS

[SWY15] Geoffrey Schiebinger, Martin J. Wainwright, and Bin Yu. The geometry of kernelized spectral clustering. Annals of Statistics, 43(2):819–846, April 2015. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1427115288.

Samworth:2012:ICA

[SY12] Richard J. Samworth and Ming Yuan. Independent component analysis via nonparametric maximum likelihood estimation. Annals of Statistics, 40(6):2973–3002, December 2012. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1360332190.

Sun:2014:SSD

[SYZL14] Yan Sun, Hongjia Yan, Wenyang Zhang, and Zudi Lu. A semiparametric spatial dynamic model. Annals of Statistics, 42(2):700–727, April 2014. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1400592175.

Serfling:2010:D

[SZ10] Robert Serfling and Yijun Zuo. Discussion. Annals of Statistics, 38(2):676–684, April 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1266586609.

Spokoiny:2015:BCS

[SZ15] Vladimir Spokoiny and Mayya Zhilova. Bootstrap confidence sets under model misspecification. Annals of Statistics, 43(6):2653–2675, December 2015. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1444222088.

Toulis:2017:AFS

[TA17] Panos Toulis and Edoardo M. Airoldi. Asymptotic and finitesample properties of estimators based on stochastic gradients. Annals of Statistics, 45(4):1694–1727, August 2017. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1498636871.

Tang:2012:LBI

[TBK12] Runlong Tang, Moulinath Banerjee, and Michael R. Kosorok. Likelihood based inference for current status data on a grid: A boundary phenomenon and an adaptive inference procedure. Annals of Statistics, 40(1):45–72, February 2012. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1331830774.

Tang:2011:TSH

[TBM11] Runlong Tang, Moulinath Banerjee, and George Michailidis. A two-stage hybrid procedure for estimating an inverse regression function. Annals of Statistics, 39(2):956– 989, April 2011. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1302268083.

Tong:2019:MLE

[TGC^{+}19] Xingwei Tong, Fuqing Gao, Kani Chen, Dingjiao Cai, and Jianguo Sun. Maximum likelihood estimation in transformed linear regression with nonnormal errors. Annals of Statistics, 47(4):1864–1892, August 2019. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://

projecteuclid.org/euclid.aos/1558425633.

Tibshirani:2014:APP

[Tib14] Ryan J. Tibshirani. Adaptive piecewise polynomial estimation via trend filtering. Annals of Statistics, 42(1):285–323, February 2014. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1395234979.

Taraldsen:2013:FTO

[TL13] Gunnar Taraldsen and Bo Henry Lindqvist. Fiducial theory and optimal inference. Annals of Statistics, 41(1):323–341, February 2013. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1364302745.

Taylor:2016:IAR

[TLT16] Jonathan E. Taylor, Joshua R. Loftus, and Ryan J. Tibshirani. Inference in adaptive regression via the Kac–Rice formula. Annals of Statistics, 44(2):743–770, April 2016. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1458245734.

Todorov:2015:JAE

[Tod15] Viktor Todorov. Jump activity estimation for pure-jump semimartingales via self-normalized statistics. Annals of Statistics, 43(4):1831–1864, August 2015. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1434546224.

Todorov:2017:TTV

[Tod17] Viktor Todorov. Testing for time-varying jump activity for pure jump semimartingales. Annals of Statistics, 45(3):1284– 1311, June 2017. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1497319695.

Tang:2018:LTE

[TP18] Minh Tang and Carey E. Priebe. Limit theorems for eigenvectors of the normalized Laplacian for random graphs. Annals of Statistics, 46(5):2360–2415, October 2018. CODEN ASTSC7. ISSN 0090-5364. URL http://projecteuclid.

org/euclid.aos/1534492839.

Tewes:2019:CSE

[TPN19] Johannes Tewes, Dimitris N. Politis, and Daniel J. Nordman. Convolved subsampling estimation with applications to block bootstrap. Annals of Statistics, 47(1):468–496, February 2019. CODEN ASTSC7. ISSN 0090-5364. URL http:/

/projecteuclid.org/euclid.aos/1543568595.

Tibshirani:2018:UAI

[TRTW18] Ryan J. Tibshirani, Alessandro Rinaldo, Rob Tibshirani, and Larry Wasserman. Uniform asymptotic inference and the bootstrap after model selection. Annals of Statistics, 46(3):1255–1287, June 2018. CODEN ASTSC7. ISSN

0090-5364. URL http://projecteuclid.org/euclid.aos/
1525313082.

Truquet:2019:LST

[Tru19] Lionel Truquet. Local stationarity and time-inhomogeneous Markov chains. Annals of Statistics, 47(4):2023–2050, August 2019. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL https://projecteuclid.org/
euclid.aos/1558425638.

Tchetgen:2012:STC

[TS12] Eric J. Tchetgen and Ilya Shpitser. Semiparametric theory for causal mediation analysis: Efficiency bounds, multiple robustness and sensitivity analysis. Annals of Statistics, 40(3):1816–1845, June 2012. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1350394518.

Tang:2013:UCV

[TSP13] Minh Tang, Daniel L. Sussman, and Carey E. Priebe. Universally consistent vertex classification for latent positions graphs. Annals of Statistics, 41(3):1406–1430, June 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1375362554.

Tibshirani:2011:SPG

[TT11] Ryan J. Tibshirani and Jonathan Taylor. The solution path of the generalized lasso. Annals of Statistics, 39(3):1335– 1371, June 2011. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1304514656.

Tibshirani:2012:DFL

[TT12a] Ryan J. Tibshirani and Jonathan Taylor. Degrees of freedom in lasso problems. Annals of Statistics, 40(2):1198– 1232, April 2012. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1342625466.

Todorov:2012:RLT

[TT12b] Viktor Todorov and George Tauchen. Realized Laplace transforms for pure-jump semimartingales. Annals of Statistics, 40(2):1233–1262, April 2012. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1342625467.

Tian:2018:SIR

[TT18] Xiaoying Tian and Jonathan Taylor. Selective inference with a randomized response. Annals of Statistics, 46(2):679–710, April 2018. CODEN ASTSC7. ISSN 0090-5364. URL http:

//projecteuclid.org/euclid.aos/1522742433.

Tsao:2013:ELF

[TW13] Min Tsao and Fan Wu. Empirical likelihood on the full parameter space. Annals of Statistics, 41(4):2176–2196, August 2013. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1382547517.

Tuo:2015:ECI

[TW15] Rui Tuo and C. F. Jeff Wu. Efficient calibration for imperfect computer models. Annals of Statistics, 43(6):2331–2352, December 2015. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1444222077.

Tang:2017:NGF

[TWT17] Chuan-Fa Tang, Dewei Wang, and Joshua M. Tebbs. Nonparametric goodness-of-fit tests for uniform stochastic ordering. Annals of Statistics, 45(6):2565–2589, December 2017. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

Tao:2013:OSV

[TWZ13] Minjing Tao, Yazhen Wang, and Harrison H. Zhou. Optimal sparse volatility matrix estimation for high-dimensional Itô processes with measurement errors. Annals of Statistics, 41(4):1816–1864, August 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1378386240.

Tang:2012:UFF

[TXL12] Yu Tang, Hongquan Xu, and Dennis K. J. Lin. Uniform fractional factorial designs. Annals of Statistics, 40(2):891– 907, April 2012. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1338515141.

Tan:2019:AMC

[TZ19a] Falong Tan and Lixing Zhu. Adaptive-to-model checking for regressions with diverging number of predictors. Annals of Statistics, 47(4):1960–1994, August 2019. CODEN ASTSC7.

ISSN 0090-5364 (print), 2168-8966 (electronic). URL https:

//projecteuclid.org/euclid.aos/1558425636.

Tan:2019:DPE

[TZ19b] Zhiqiang Tan and Cun-Hui Zhang. Doubly penalized estimation in additive regression with high-dimensional data. Annals of Statistics, 47(5):2567–2600, October 2019. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/euclid.aos/
1564797857.

Uhler:2012:GML

[Uhl12] Caroline Uhler. Geometry of maximum likelihood estimation in Gaussian graphical models. Annals of Statistics, 40(1):238–261, February 2012. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1333029964.

Uhler:2018:EFN

[ULR18] Caroline Uhler, Alex Lenkoski, and Donald Richards. Exact formulas for the normalizing constants of Wishart distributions for graphical models. Annals of Statistics, 46(1):90–118, February 2018. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

Uhler:2013:GFA

[URBY13] Caroline Uhler, Garvesh Raskutti, Peter Bühlmann, and Bin Yu. Geometry of the faithfulness assumption in causal inference. Annals of Statistics, 41(2):436–463, April 2013. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1366138197.

Verzelen:2017:DFS

[VAC17] Nicolas Verzelen and Ery Arias-Castro. Detection and feature selection in sparse mixture models. Annals of Statistics, 45 (5):1920–1950, October 2017. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

Volgushev:2019:DIQ

[VCC19] Stanislav Volgushev, Shih-Kang Chao, and Guang Cheng. Distributed inference for quantile regression processes. Annals of Statistics, 47(3):1634–1662, June 2019. CODEN ASTSC7. ISSN 0090-5364. URL http://projecteuclid.

org/euclid.aos/1550026852.

Vogt:2015:DGC

[VD15] Michael Vogt and Holger Dette. Detecting gradual changes in locally stationary processes. Annals of Statistics, 43(2):713– 740, April 2015. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1425398506.

vandeGeer:2013:PML

[vdGB13] Sara van de Geer and Peter Bühlmann. ℓ_{0}-penalized maximum likelihood for sparse directed acyclic graphs. Annals of Statistics, 41(2):536–567, April 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1366980557.

Verzelen:2010:AES

[Ver10] Nicolas Verzelen. Adaptive estimation of stationary Gaussian fields. Annals of Statistics, 38(3):1363–1402, June 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1268056620.

Vimond:2010:EES

[Vim10] Myriam Vimond. Efficient estimation for a subclass of shape invariant models. Annals of Statistics, 38(3):1885– 1912, June 2010. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1271271281.

Vu:2013:MSP

[VL13] Vincent Q. Vu and Jing Lei. Minimax sparse principal subspace estimation in high dimensions. Annals of Statistics, 41(6):2905–2947, December 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1388545673.

Velasco:2018:FDM

[VL18] Carlos Velasco and Ignacio N. Lobato. Frequency domain minimum distance inference for possibly noninvertible and noncausal ARMA models. Annals of Statistics, 46(2):555– 579, April 2018. CODEN ASTSC7. ISSN 0090-5364. URL

http://projecteuclid.org/euclid.aos/1522742429.

Vogt:2012:NRL

[Vog12] Michael Vogt. Nonparametric regression for locally stationary time series. Annals of Statistics, 40(5):2601–2633, October 2012. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1359987532.

VanderWeele:2012:GTI

[VR12] Tyler J. VanderWeele and Thomas S. Richardson. General theory for interactions in sufficient cause models with dichotomous exposures. Annals of Statistics, 40(4):2128–2161, August 2012. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1358951378.

Veitch:2019:SES

[VR19] Victor Veitch and Daniel M. Roy. Sampling and estimation for (sparse) exchangeable graphs. Annals of Statistics, 47 (6):3274–3299, December 2019. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://

projecteuclid.org/euclid.aos/1572487393.

VanderWeele:2013:DC

[VS13] Tyler J. VanderWeele and Ilya Shpitser. On the definition of a confounder. Annals of Statistics, 41(1):196–220, February 2013. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1364302740.

Vandermeulen:2019:OTA

[VS19] Robert A. Vandermeulen and Clayton D. Scott. An operator theoretic approach to nonparametric mixture models. Annals of Statistics, 47(5):2704–2733, October 2019. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/euclid.aos/
1564797861.

Verzelen:2010:GFT

[VV10] Nicolas Verzelen and Fanny Villers. Goodness-of-fit tests for high-dimensional Gaussian linear models. Annals of Statistics, 38(2):704–752, April 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1266586612.

vanZwet:2011:REL

[vZ11] Willem R. van Zwet. Remembering Erich Lehmann. Annals of Statistics, 39(5):2266–2279, October 2011. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1321020524.

Wainwright:2012:DLV

[Wai12] Martin J. Wainwright. Discussion: Latent variable graphical model selection via convex optimization. Annals of Statistics, 40(4):1978–1983, August 2012. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1351602530.

Walther:2010:OFD

[Wal10] Guenther Walther. Optimal and fast detection of spatial clusters with scan statistics. Annals of Statistics, 38(2):1010– 1033, April 2010. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1266586621.

Wang:2010:CSO

[Wan10] Weizhen Wang. On construction of the smallest one-sided confidence interval for the difference of two proportions. Annals of Statistics, 38(2):1227–1243, April 2010. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1266586628.

Wang:2011:GAC

[Wan11] Lan Wang. GEE analysis of clustered binary data with diverging number of covariates. Annals of Statistics, 39 (1):389–417, February 2011. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1291388380.

Wang:2013:AEQ

[Wan13] Yazhen Wang. Asymptotic equivalence of quantum state tomography and noisy matrix completion. Annals of Statistics, 41(5):2462–2504, October 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1383661270.

Wasserman:2014:DST

[Was14] Larry Wasserman. Discussion: “A significance test for the lasso”. Annals of Statistics, 42(2):501–508, April 2014. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/

1400592166. See [LTTT14c].

Wang:2017:LBM

[WB17] Y. X. Rachel Wang and Peter J. Bickel. Likelihood-based model selection for stochastic block models. Annals of Statistics, 45(2):500–528, April 2017. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1494921948.

Wang:2016:SCT

[WBS16] Tengyao Wang, Quentin Berthet, and Richard J. Samworth. Statistical and computational trade-offs in estimation of sparse principal components. Annals of Statistics, 44(5):1896–1930, October 2016. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1473685263.

Chang:2019:NTM

[wCM19] Hsin wen Chang and Ian W. McKeague. Nonparametric testing for multiple survival functions with noninferiority margins. Annals of Statistics, 47(1):205–232, February 2019. CODEN ASTSC7. ISSN 0090-5364. URL http://

projecteuclid.org/euclid.aos/1543568587.

Wolpert:2011:SEU

[WCT11] Robert L. Wolpert, Merlise A. Clyde, and Chong Tu. Stochastic expansions using continuous dictionaries: Lévy adaptive regression kernels. Annals of Statistics, 39(4):1916–1962, August 2011. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1314190619.

Wei:2010:D

[Wei10] Ying Wei. Discussion. Annals of Statistics, 38(2):670–675, April 2010. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1266586608.

Wang:2017:AEE

[WF17] Weichen Wang and Jianqing Fan. Asymptotics of empirical eigenstructure for high dimensional spiked covariance. Annals of Statistics, 45(3):1342–1374, June 2017. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1497319697.

Williams:2019:NVS

[WH19] Jonathan P. Williams and Jan Hannig. Nonpenalized variable selection in high-dimensional linear model settings via generalized fiducial inference. Annals of Statistics, 47(3):1723– 1753, June 2019. CODEN ASTSC7. ISSN 0090-5364. URL

http://projecteuclid.org/euclid.aos/1550026855.

Wang:2013:CNP

[WKL13] Lan Wang, Yongdai Kim, and Runze Li. Calibrating nonconvex penalized regression in ultra-high dimension. Annals of Statistics, 41(5):2505–2536, October 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1383661271.

Whiteley:2014:TPF

[WL14] Nick Whiteley and Anthony Lee. Twisted particle filters. Annals of Statistics, 42(1):115–141, February 2014. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1392733182.

Wang:2014:OEE

[WLCY14] Jiangyan Wang, Rong Liu, Fuxia Cheng, and Lijian Yang. Oracally efficient estimation of autoregressive error distribution with simultaneous confidence band. Annals of Statistics, 42(2):654–668, April 2014. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1400592173.

Wang:2011:EVS

[WLLC11] Li Wang, Xiang Liu, Hua Liang, and Raymond J. Carroll. Estimation and variable selection for generalized additive partial linear models. Annals of Statistics, 39(4):1827–1851, August 2011. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1311688537.

Wang:2014:OCS

[WLZ14] Zhaoran Wang, Han Liu, and Tong Zhang. Optimal computational and statistical rates of convergence for sparse nonconvex learning problems. Annals of Statistics, 42(6):2164–2201, December 2014. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1413810725.

Wong:2010:OPT

[WM10] Wing H. Wong and Li Ma. Optional Pólya tree and Bayesian inference. Annals of Statistics, 38(3):1433–1459, June 2010. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1268056622.

Weng:2018:OLP

[WMZ18] Haolei Weng, Arian Maleki, and Le Zheng. Overcoming the limitations of phase transition by higher order analysis of regularization techniques. Annals of Statistics, 46 (6A):3099–3129, December 2018. CODEN ASTSC7. ISSN

0090-5364. URL http://projecteuclid.org/euclid.aos/
1536307244.

Wang:2012:STN

[WP12] Qiying Wang and Peter C. B. Phillips. A specification test for nonlinear nonstationary models. Annals of Statistics, 40(2):727–758, April 2012. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1337268210.

Woodard:2013:CRM

[WR13] Dawn B. Woodard and Jeffrey S. Rosenthal. Convergence rate of Markov chain methods for genomic motif discovery. Annals of Statistics, 41(1):91–124, February 2013. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1362493041.

Wang:2019:AFM

[WRM19] Nanwei Wang, Johannes Rauh, and Hélène Massam. Approximating faces of marginal polytopes in discrete hierarchical models. Annals of Statistics, 47(3):1203–1233, June 2019. CODEN ASTSC7. ISSN 0090-5364. URL http://

projecteuclid.org/euclid.aos/1550026834.

Wu:2018:TGF

[WW18] Hau-Tieng Wu and Nan Wu. Think globally, fit locally under the manifold setup: Asymptotic analysis of locally linear embedding. Annals of Statistics, 46(6B):3805–3837, December 2018. CODEN ASTSC7. ISSN 0090-5364. URL http://

projecteuclid.org/euclid.aos/1536631291.

Wei:2019:GHT

[WWG19] Yuting Wei, Martin J. Wainwright, and Adityanand Guntuboyina. The geometry of hypothesis testing over convex cones: Generalized likelihood ratio tests and minimax radii. Annals of Statistics, 47(2):994–1024, April 2019. CODEN ASTSC7. ISSN 0090-5364. URL http://projecteuclid.

org/euclid.aos/1547197246.

Wang:2014:EMS

[WXQL14] Li Wang, Lan Xue, Annie Qu, and Hua Liang. Estimation and model selection in generalized additive partial linear models for correlated data with diverging number of covariates. Annals of Statistics, 42(2):592–624, April 2014. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1400592171.

Wang:2018:OMD

[WXX18] Lin Wang, Qian Xiao, and Hongquan Xu. Optimal maximin L_{1}-distance latin hypercube designs based on good lattice point designs. Annals of Statistics, 46(6B):3741–3766, December 2018. CODEN ASTSC7. ISSN 0090-5364. URL

http://projecteuclid.org/euclid.aos/1536631289.

Wang:2010:EPL

[WXZC10] Jane-Ling Wang, Liugen Xue, Lixing Zhu, and Yun Sam Chong. Estimation for a partial-linear single-index model. Annals of Statistics, 38(1):246–274, February 2010. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1262271615.

Wang:2017:EEL

[WY17] Qinwen Wang and Jianfeng Yao. Extreme eigenvalues of large-dimensional spiked Fisher matrices with application. Annals of Statistics, 45(1):415–460, February 2017. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1487667628.

Wu:2019:CPM

[WY19] Yihong Wu and Pengkun Yang. Chebyshev polynomials, moment matching, and optimal estimation of the unseen. Annals of Statistics, 47(2):857–883, April 2019. CODEN ASTSC7.

ISSN 0090-5364. URL http://projecteuclid.org/euclid.
aos/1547197241.

Wang:2010:VVM

[WZ10] Yazhen Wang and Jian Zou. Vast volatility matrix estimation for high-frequency financial data. Annals of Statistics, 38(2):943–978, April 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1266586619.

Wu:2012:PMT

[WZ12] Yuan Wu and Ying Zhang. Partially monotone tensor spline estimation of the joint distribution function with bivariate current status data. Annals of Statistics, 40(3):1609–1636, June 2012. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1346850067.

Wu:2018:GBS

[WZ18] Weichi Wu and Zhou Zhou. Gradient-based structural change detection for nonstationary time series M-estimation. Annals of Statistics, 46(3):1197–1224, June 2018. CODEN ASTSC7.

ISSN 0090-5364. URL http://projecteuclid.org/euclid.
aos/1525313080.

Wang:2017:CAM

[WZHO17] Jingshu Wang, Qingyuan Zhao, Trevor Hastie, and Art B. Owen. Confounder adjustment in multiple hypothesis testing. Annals of Statistics, 45(5):1863–1894, October 2017. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

Wang:2018:CPD

[WZY18] Guanghui Wang, Changliang Zou, and Guosheng Yin. Change-point detection in multinomial data with a large number of categories. Annals of Statistics, 46(5):2020–2044, October 2018. CODEN ASTSC7. ISSN 0090-5364. URL

http://projecteuclid.org/euclid.aos/1534492827.

Xu:2016:FVS

[XCL16] Min Xu, Minhua Chen, and John Lafferty. Faithful variable screening for high-dimensional convex regression. Annals of Statistics, 44(6):2624–2660, December 2016. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1479891630.

Xu:2012:AOE

[XH12] Ganggang Xu and Jianhua Z. Huang. Asymptotic optimality and efficient computation of the leave-subject-out cross-validation. Annals of Statistics, 40(6):3003–3030, December 2012. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1360332191.

Xie:2016:OSE

[XKB16] Xianchao Xie, S. C. Kou, and Lawrence Brown. Optimal shrinkage estimation of mean parameters in family of distributions with quadratic variance. Annals of Statistics, 44(2):564–597, April 2016. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1458245728.

Xue:2010:SEC

[XMW10] Hongqi Xue, Hongyu Miao, and Hulin Wu. Sieve estimation of constant and time-varying coefficients in nonlinear ordinary differential equation models by considering both numerical error and measurement error. Annals of Statistics, 38(4):2351–2387, August 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1278861251.

Xia:2013:CRE

[XQB13] Ningning Xia, Yingli Qin, and Zhidong Bai. Convergence rates of eigenvector empirical spectral distribution of large dimensional sample covariance matrix. Annals of Statistics, 41(5):2572–2607, October 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1384871346.

Xie:2013:SMS

[XS13] Yao Xie and David Siegmund. Sequential multi-sensor change-point detection. Annals of Statistics, 41(2):670–692, April 2013. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1366980561.

Xu:2017:IRL

[Xu17] Gongjun Xu. Identifiability of restricted latent class models with binary responses. Annals of Statistics, 45(2):675– 707, April 2017. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1494921954.

Xiao:2012:CME

[XW12] Han Xiao and Wei Biao Wu. Covariance matrix estimation for stationary time series. Annals of Statistics, 40(1):466–493, February 2012. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1334581750.

Xue:2012:RRB

[XZ12] Lingzhou Xue and Hui Zou. Regularized rank-based estimation of high-dimensional nonparanormal graphical models. Annals of Statistics, 40(5):2541–2571, October 2012. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1359987530.

Xia:2018:IAS

[XZ18] Ningning Xia and Xinghua Zheng. On the inference about the spectral distribution of high-dimensional covariance matrix based on high-frequency noisy observations. Annals of Statistics, 46(2):500–525, April 2018. CODEN ASTSC7. ISSN

0090-5364. URL http://projecteuclid.org/euclid.aos/
1522742427.

Xue:2012:NPC

[XZC12] Lingzhou Xue, Hui Zou, and Tianxi Cai. Nonconcave penalized composite conditional likelihood estimation of sparse Ising models. Annals of Statistics, 40(3):1403–1429, June 2012. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.
aos/1344610588.

Yang:2010:GP

[Yan10] Min Yang. On the de la Garza phenomenon. Annals of Statistics, 38(4):2499–2524, August 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1278861255.

Yuan:2010:RKH

[YC10] Ming Yuan and T. Tony Cai. A reproducing kernel Hilbert space approach to functional linear regression. Annals of Statistics, 38(6):3412–3444, December 2010. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1291126962.

Yang:2016:BMR

[YD16] Yun Yang and David B. Dunson. Bayesian manifold regression. Annals of Statistics, 44(2):876–905, April 2016. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1458245738.

Yu:2016:MSI

[YDS16] Zhou Yu, Yuexiao Dong, and Jun Shao. On marginal sliced inverse regression for ultrahigh dimensional model-free feature selection. Annals of Statistics, 44(6):2594–2623, December 2016. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1479891629.

Yen:2011:MMA

[Yen11] Tso-Jung Yen. A majorization–minimization approach to variable selection using spike and slab priors. Annals of Statistics, 39(3):1748–1775, June 2011. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1311600282.

Yamagata:2013:QLA

[YFG13] Koichi Yamagata, Akio Fujiwara, and Richard D. Gill. Quantum local asymptotic normality based on a new quantum likelihood ratio. Annals of Statistics, 41(4):2197–2217, August 2013. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1382547518.

Yoo:2016:SNP

[YG16] William Weimin Yoo and Subhashis Ghosal. Supremum norm posterior contraction and credible sets for nonparametric multivariate regression. Annals of Statistics, 44(3):1069– 1102, June 2016. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1460381687.

Yang:2012:BEL

[YH12] Yunwen Yang and Xuming He. Bayesian empirical likelihood for quantile regression. Annals of Statistics, 40(2):1102– 1131, April 2012. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1342625463.

Yin:2011:SDR

[YL11] Xiangrong Yin and Bing Li. Sufficient dimension reduction based on an ensemble of minimum average variance estimators. Annals of Statistics, 39(6):3392–3416, December 2011. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1330958684.

Yoshimori:2014:SOE

[YL14] Masayo Yoshimori and Partha Lahiri. A second-order efficient empirical Bayes confidence interval. Annals of Statistics, 42(4):1233–1261, August 2014. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1403715200.

Yan:2016:ADE

[YLZ16] Ting Yan, Chenlei Leng, and Ji Zhu. Asymptotics in directed exponential random graph models with an increasing bi-degree sequence. Annals of Statistics, 44(1):31–57, February 2016. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1449755956.

Yang:2015:ITH

[YP15] Yanrong Yang and Guangming Pan. Independence test for high dimensional data based on regularized canonical correlation coefficients. Annals of Statistics, 43(2):467–500, April 2015. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1424787425.

Yang:2017:RSK

[YPW17] Yun Yang, Mert Pilanci, and Martin J. Wainwright. Randomized sketches for kernels: Fast and optimal nonparametric regression. Annals of Statistics, 45(3):991–1023, June 2017. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1497319686.

Yang:2012:ILO

[YS12] Min Yang and John Stufken. Identifying locally optimal designs for nonlinear models: A simple extension with profound consequences. Annals of Statistics, 40(3):1665–1681, June 2012. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1349196387.

Yang:2015:MON

[YT15] Yun Yang and Surya T. Tokdar. Minimax-optimal nonparametric regression in high dimensions. Annals of Statistics, 43(2):652–674, April 2015. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1425398504.

Yu:2010:MCG

[Yu10] Yaming Yu. Monotonic convergence of a general algorithm for computing optimal designs. Annals of Statistics, 38(3):1593– 1606, June 2010. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1269452648.

Yuan:2012:DLV

[Yua12] Ming Yuan. Discussion: Latent variable graphical model selection via convex optimization. Annals of Statistics, 40 (4):1968–1972, August 2012. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1351602528.

Yang:2016:CCH

[YWJ16] Yun Yang, Martin J. Wainwright, and Michael I. Jordan. On the computational complexity of high-dimensional Bayesian variable selection. Annals of Statistics, 44(6):2497–2532, December 2016. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1479891626.

Yuan:2016:MOR

[YZ16] Ming Yuan and Ding-Xuan Zhou. Minimax optimal rates of estimation in high dimensional additive models. Annals of Statistics, 44(6):2564–2593, December 2016. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1479891628.

Zheng:2017:IUO

[ZAL17] Wei Zheng, Mingyao Ai, and Kang Li. Identification of universally optimal circular designs for the interference model. Annals of Statistics, 45(4):1462–1487, August 2017. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1498636863.

Zhang:2019:SSI

[ZBC19] Anru Zhang, Lawrence D. Brown, and T. Tony Cai. Semisupervised inference: General theory and estimation of means. Annals of Statistics, 47(5):2538–2566, October 2019. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/euclid.aos/
1564797856.

Zhou:2018:NPR

[ZBFL18] Wen-Xin Zhou, Koushiki Bose, Jianqing Fan, and Han Liu. A new perspective on robust M-estimation: Finite sample theory and applications to dependence-adjusted multiple testing.

Annals of Statistics, 46(5):1904–1931, October 2018. CODEN ASTSC7. ISSN 0090-5364. URL http://projecteuclid.

org/euclid.aos/1534492823.

Zheng:2015:SPC

[ZBY15] Shurong Zheng, Zhidong Bai, and Jianfeng Yao. Substitution principle for CLT of linear spectral statistics of highdimensional sample covariance matrices with applications to hypothesis testing. Annals of Statistics, 43(2):546–591, April 2015. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1424787428.

Zheng:2019:HTL

[ZCCL19] Shurong Zheng, Zhao Chen, Hengjian Cui, and Runze Li. Hypothesis testing on linear structures of high-dimensional covariance matrix. Annals of Statistics, 47(6):3300–3334, December 2019. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL https://projecteuclid.org/
euclid.aos/1572487394.

Zheng:2019:THD

[ZCGZ19] Shurong Zheng, Guanghui Cheng, Jianhua Guo, and Hongtu Zhu. Test for high-dimensional correlation matrices. Annals of Statistics, 47(5):2887–2921, October 2019. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/euclid.aos/
1564797867.

Zhao:2016:PLF

[ZCL16] Tianqi Zhao, Guang Cheng, and Han Liu. A partially linear framework for massive heterogeneous data. Annals of Statistics, 44(4):1400–1437, August 2016. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1467894703.

Zhong:2013:TAH

[ZCX13] Ping-Shou Zhong, Song Xi Chen, and Minya Xu. Tests alternative to higher criticism for high-dimensional means under sparsity and column-wise dependence. Annals of Statistics, 41(6):2820–2851, December 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1387313391.

Zhao:2018:RBC

[ZDMD18] Anqi Zhao, Peng Ding, Rahul Mukerjee, and Tirthankar Dasgupta. Randomization-based causal inference from splitplot designs. Annals of Statistics, 46(5):1876–1903, October 2018. CODEN ASTSC7. ISSN 0090-5364. URL http://

projecteuclid.org/euclid.aos/1534492822.

Zhang:2011:MTF

[ZFY11] Chunming Zhang, Jianqing Fan, and Tao Yu. Multiple testing via FDR _{L} for large-scale imaging data. Annals of Statistics, 39(1):613–642, February 2011. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1297779858.

Zhu:2010:SMR

[ZH10] Hongjian Zhu and Feifang Hu. Sequential monitoring of response-adaptive randomized clinical trials. Annals of Statistics, 38(4):2218–2241, August 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1278861247.

Zhang:2010:NUV

[Zha10] Cun-Hui Zhang. Nearly unbiased variable selection under minimax concave penalty. Annals of Statistics, 38(2):894– 942, April 2010. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1266586618.

Zhang:2013:NOM

[Zha13] Li Zhang. Nearly optimal minimax estimator for highdimensional sparse linear regression. Annals of Statistics, 41(4):2149–2175, August 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1382547516.

Zhang:2018:EQT

[Zha18] Yichong Zhang. Extremal quantile treatment effects. Annals of Statistics, 46(6B):3707–3740, December 2018. CODEN ASTSC7. ISSN 0090-5364. URL http://projecteuclid.

org/euclid.aos/1536631288.

Zhang:2019:CEL

[Zha19a] Anru Zhang. Cross: Efficient low-rank tensor completion. Annals of Statistics, 47(2):936–964, April 2019. CODEN ASTSC7. ISSN 0090-5364. URL http://projecteuclid.

org/euclid.aos/1547197244.

Zhao:2019:CBP

[Zha19b] Qingyuan Zhao. Covariate balancing propensity score by tailored loss functions. Annals of Statistics, 47(2):965–993, April 2019. CODEN ASTSC7. ISSN 0090-5364. URL http://

projecteuclid.org/euclid.aos/1547197245.

Zhang:2011:IUM

[ZHCC11] Li-Xin Zhang, Feifang Hu, Siu Hung Cheung, and Wai Sum Chan. Immigrated urn models — theoretical properties and applications. Annals of Statistics, 39(1):643–671, February 2011. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1297779859.

Zheng:2013:OCD

[Zhe13a] Wei Zheng. Optimal crossover designs for the proportional model. Annals of Statistics, 41(4):2218–2235, August 2013. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1382547519.

Zheng:2013:UOC

[Zhe13b] Wei Zheng. Universally optimal crossover designs under subject dropout. Annals of Statistics, 41(1):63–90, February 2013. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1362493040.

Zheng:2015:UOD

[Zhe15] Wei Zheng. Universally optimal designs for two interference models. Annals of Statistics, 43(2):501–518, April 2015. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1424787426.

Zhou:2010:NIQ

[Zho10] Zhou Zhou. Nonparametric inference of quantile curves for nonstationary time series. Annals of Statistics, 38 (4):2187–2217, August 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1278861246.

Zhou:2014:GGE

[Zho14a] Shuheng Zhou. Gemini: Graph estimation with matrix variate normal instances. Annals of Statistics, 42(2):532–562, April 2014. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1400592169.

Zhou:2014:IWS

[Zho14b] Zhou Zhou. Inference of weighted V -statistics for nonstationary time series and its applications. Annals of Statistics, 42(1):87–114, February 2014. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1389795746.

Zhu:2017:NAS

[Zhu17] Ying Zhu. Nonasymptotic analysis of semiparametric regression models with high-dimensional parametric coefficients. Annals of Statistics, 45(5):2274–2298, October 2017. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

Zhu:2019:SIA

[Zhu19] Ke Zhu. Statistical inference for autoregressive models under heteroscedasticity of unknown form. Annals of Statistics, 47(6):3185–3215, December 2019. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https:

//projecteuclid.org/euclid.aos/1572487390.

Zhu:2012:PSC

[ZIC12] Hongtu Zhu, Joseph G. Ibrahim, and Hyunsoon Cho. Perturbation and scaled Cook’s distance. Annals of Statistics, 40(2):785–811, April 2012. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1337268212.

Zhang:2011:CIC

[ZJS11] Mingyuan Zhang, Marshall M. Joffe, and Dylan S. Small. Causal inference for continuous-time processes when covariates are observed only at discrete times. Annals of Statistics, 39(1):131–173, February 2011. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1291388371.

Zhang:2011:FIC

[ZL11a] Xinyu Zhang and Hua Liang. Focused information criterion and model averaging for generalized additive partial linear models. Annals of Statistics, 39(1):174–200, February 2011. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1291388372.

Zheng:2011:EIC

[ZL11b] Xinghua Zheng and Yingying Li. On the estimation of integrated covariance matrices of high dimensional diffusion processes. Annals of Statistics, 39(6):3121–3151, December 2011. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1327672848.

Zhu:2011:GSW

[ZL11c] Ke Zhu and Shiqing Ling. Global self-weighted and local quasi-maximum exponential likelihood estimators for ARMA–GARCH/IGARCH models. Annals of Statistics, 39(4):2131–2163, August 2011. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1319595461.

Zhu:2012:MVC

[ZLK12] Hongtu Zhu, Runze Li, and Linglong Kong. Multivariate varying coefficient model for functional responses. Annals of Statistics, 40(5):2634–2666, October 2012. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1359987533.

Zhao:2013:HDI

[ZLLW13] Junlong Zhao, Chenlei Leng, Lexin Li, and Hansheng Wang. High-dimensional influence measure. Annals of Statistics, 41(5):2639–2667, October 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1384871348.

Zhong:2017:TCS

[ZLST17] Ping-Shou Zhong, Wei Lan, Peter X. K. Song, and Chih-Ling Tsai. Tests for covariance structures with high-dimensional repeated measurements. Annals of Statistics, 45(3):1185– 1213, June 2017. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1497319692.

Zhao:2012:CCD

[ZLZ12] Yunpeng Zhao, Elizaveta Levina, and Ji Zhu. Consistency of community detection in networks under degreecorrected stochastic block models. Annals of Statistics, 40(4):2266–2292, August 2012. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1358951382.

Zhao:2018:PCO

[ZLZ18] Tuo Zhao, Han Liu, and Tong Zhang. Pathwise coordinate optimization for sparse learning: Algorithm and theory. Annals of Statistics, 46(1):180–218, February 2018. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

Zhang:2018:CLE

[ZPG18] Bo Zhang, Guangming Pan, and Jiti Gao. CLT for largest eigenvalues and unit root testing for high-dimensional nonstationary time series. Annals of Statistics, 46(5):2186–2215, October 2018. CODEN ASTSC7. ISSN 0090-5364. URL

http://projecteuclid.org/euclid.aos/1534492833.

Zheng:2015:GAQ

[ZPH15] Qi Zheng, Limin Peng, and Xuming He. Globally adaptive quantile regression with ultra-high dimensional data. Annals of Statistics, 43(5):2225–2258, October 2015. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1442364151.

Zheng:2018:HDC

[ZPH18] Qi Zheng, Limin Peng, and Xuming He. High dimensional censored quantile regression. Annals of Statistics, 46(1):308– 343, February 2018. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

Zhu:2017:NVA

[ZPL^{+}17] Xuening Zhu, Rui Pan, Guodong Li, Yuewen Liu, and Hansheng Wang. Network vector autoregression. Annals of Statistics, 45(3):1096–1123, June 2017. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1497319689.

Zhang:2011:TAQ

[ZPMX11] Runchu Zhang, Frederick K. H. Phoa, Rahul Mukerjee, and Hongquan Xu. A trigonometric approach to quaternary code designs with application to one-eighth and onesixteenth fractions. Annals of Statistics, 39(2):931–955, April 2011. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1302268082.

Zhang:2013:TCM

[ZPW13] Rongmao Zhang, Liang Peng, and Ruodu Wang. Tests for covariance matrix with fixed or divergent dimension. Annals of Statistics, 41(4):2075–2096, August 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1382547513.

Zhang:2013:FSA

[ZS13] Xianyang Zhang and Xiaofeng Shao. Fixed-smoothing asymptotics for time series. Annals of Statistics, 41(3):1329– 1349, June 2013. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1372979640.

Zhou:2019:SPO

[ZvdAW19] Bo Zhou, Ramon van den Akker, and Bas J. M. Werker. Semiparametrically point-optimal hybrid rank tests for unit roots. Annals of Statistics, 47(5):2601–2638, October 2019. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/euclid.aos/
1564797858.

Zhang:2012:ITV

[ZW12] Ting Zhang and Wei Biao Wu. Inference of time-varying regression models. Annals of Statistics, 40(3):1376–1402, June 2012. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1344610587.

Zhang:2015:TVN

[ZW15] Ting Zhang and Wei Biao Wu. Time-varying nonlinear regression models: Nonparametric estimation and model selection. Annals of Statistics, 43(2):741–768, April 2015. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://projecteuclid.org/euclid.aos/
1425398507.

Zhang:2016:SDF

[ZW16] Xiaoke Zhang and Jane-Ling Wang. From sparse to dense functional data and beyond. Annals of Statistics, 44 (5):2281–2321, October 2016. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1473685276.

Zhang:2017:GAH

[ZW17] Danna Zhang and Wei Biao Wu. Gaussian approximation for high dimensional time series. Annals of Statistics, 45(5): 1895–1919, October 2017. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

Zou:2014:NML

[ZYFW14] Changliang Zou, Guosheng Yin, Long Feng, and Zhaojun Wang. Nonparametric maximum likelihood approach to multiple change-point problems. Annals of Statistics, 42(3):970– 1002, June 2014. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1400592649.

Zhao:2018:ARC

[ZYL18] Junlong Zhao, Guan Yu, and Yufeng Liu. Assessing robustness of classification using an angular breakdown point. Annals of Statistics, 46(6B):3362–3389, December 2018. CODEN ASTSC7. ISSN 0090-5364. URL http://

projecteuclid.org/euclid.aos/1536631277.

Zhang:2018:CMQ

[ZYS18] Xianyang Zhang, Shun Yao, and Xiaofeng Shao. Conditional mean and quantile dependence testing in high dimension. Annals of Statistics, 46(1):219–246, February 2018. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

Zhang:2016:MRC

[ZZ16] Anderson Y. Zhang and Harrison H. Zhou. Minimax rates of community detection in stochastic block models. Annals of Statistics, 44(5):2252–2280, October 2016. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1473685275.

Zhu:2018:MTI

[ZZX18] Liping Zhu, Yaowu Zhang, and Kai Xu. Measuring and testing for interval quantile dependence. Annals of Statistics, 46 (6A):2683–2710, December 2018. CODEN ASTSC7. ISSN

0090-5364. URL http://projecteuclid.org/euclid.aos/
1536307230.